### v1

# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

Abstract.
We study compact hyperbolic surfaces and multiplicative observables, establishing a large-scale analogue of Zelditch’s quantum mixing theorem with hypotheses that hold for both arithmetic and Weil-Petersson random surfaces of large genus. This complements the large-scale quantum ergodicity theorems of Le Masson and Sahlsten, which themselves serve as large-scale analogues of the quantum ergodicity theorem of Shnirelman, Zelditch, and de Vèrdiere, thereby providing a more complete picture of the asymptotic behaviour of observables in the large-scale limit. Our approach does not rely on ball averaging operators or Nevo’s ergodic theorem. Instead, we introduce a new method based on the hyperbolic wave equation and the quantitative exponential mixing of the geodesic flow established by Ratner and Matheus.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University
1. Introduction
1.1. Background
In quantum systems, the state of the system can be described in terms of the eigenfunctions of its energy operator. Their evolution is governed by the Schrödinger equation, the quantum analogue of Newton’s equations of motion. Observables correspond to measurable quantities of the system. Understanding the interplay between the energy operator, its eigenfunctions, and the observables leads to a deeper understanding of the dynamics of the system. A central topic in quantum chaos is the study of localization and delocalization phenomena for the eigenfunctions of the energy operator. In this paper, the Hamiltonian system is assumed to be the geodesic flow $\phi_{t}$ of a compact connected Riemannian manifold $(M,g)$ and the quantum Hamiltonian is taken to be $\sqrt{-\Delta}$, where $\Delta$ is the Laplace-Beltrami operator on the manifold. For more background and a better overview, we refer to the following surveys: [Mar06, Zel05, Sar11, Zel19, Ana22].
Consider a compact Riemannian manifold $(M,g)$ and the quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudodifferential operator representing an observable. We are interested in the diagonal elements $\langle\psi_{j},A\psi_{j}\rangle$, which correspond to the expected value of $A$ in the state $\psi_{j}$, and in the off-diagonal elements $\langle\psi_{j},A\psi_{k}\rangle$, which represent transition amplitudes from $\psi_{j}$ to $\psi_{k}$ induced by the observation of $A$.
Quantum ergodicity theorem (Shnirelman [Shn74], Zelditch [Zel87], and Colin de Verdière [Col85]) is a cornerstone result of quantum chaos and serves as a quantum analogue of the classical ergodic theorem. For an ergodic compact Riemannian manifold, it states that almost all subsequences of eigenfunctions corresponding to increasing eigenvalues satisfy
$\langle\psi_{j_{l}},A\psi_{j_{l}}\rangle\xrightarrow{l\to\infty}\omega(A),$
where $\omega(A)$ denotes the average of the principal symbol $\sigma_{A}$ of $A$ (the classical counterpart of observable $A$) over the unit cotangent bundle of $M$. Roughly speaking, the theorem asserts that the probability distributions induced by the eigenfunctions converge to the uniform distribution in the weak sense of testing the distributions against observables.
Shnirelman’s theorem concerns the diagonal elements of observables. Zelditch [Zel90, Zel96b] extended it to the off-diagonal elements, given that the compact Riemannian surface is now weak mixing, proving that for any $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
$\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}-\tau|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j},A\psi_{k}\right\rangle\right|^{2}<\epsilon.$
Informally, this means that the summed transition amplitudes between states separated approximately by energy $\tau$ decay to zero for almost all subsequences of eigenfunctions as the eigenvalue increases.
A unified formulation encompassing both of these results can be given following the works of multiple authors: [Shn74, Col85, Zel87, Laz93, Zel90, GL93, Zel96a, Zel96b, ZZ96, Sun97]. The version below adapts Theorems 4.1 and 4.4 from [Zel05].
Theorem 1.1 (Semi-classical quantum mixing, [Sun97, Zel05]).
Let $(M,g)$ be a compact Riemannian manifold (possibly with a boundary) with quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenvalues $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$. Let $A$ be a zeroth-order pseudodifferential operator. Consider the following properties:
  • (i)
    $\lim_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\big|\langle\psi_{j},A\psi_{j}\rangle-\omega(A)\big|^{2}=0,$
  • (ii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{\lambda_{j}<\lambda}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}|<\delta(\epsilon)\end{subarray}}\left|\langle\psi_{j},A\psi_{k}\rangle\right|^{2}<\epsilon,$
  • (iii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j},A\psi_{k}\rangle|^{2}<\epsilon.$
Then the geodesic flow $\varphi_{t}$ is ergodic on $(S^{*}M,d\mu)$ if and only if properties (i) and (ii) hold. It is weakly mixing on $(S^{*}M,d\mu)$ if and only if properties (i), (ii), and (iii) hold. Here $S^{*}M$ denotes the unit tangent bundle of $M$ and $d\mu$ is the Liouville measure on $S^{*}M$.
This theorem exhibits a delicate connection between the classical and quantum settings. If properties (i) and (ii) hold in the quantum setting, the corresponding classical geodesic flow is ergodic, and conversely. Likewise, if Properties (i)(iii) hold, the classical flow is weak mixing, and vice versa. We therefore say that a quantum system satisfying Properties (i) and (ii) is quantum ergodic, while one satisfying Properties (i)(iii) is quantum mixing.
Our focus will be on the compact connected hyperbolic surfaces, which are inherently chaotic due to their (constant) negative curvature. In this context, the following theorem of Ratner [Rat87], with quantitative bounds due to Matheus [Mat13], will be central. The theorem is stated here in a slightly modified form using only the physical space observables.
Theorem 1.2 (Position space exponential mixing).
Let $X$ be a compact connected hyperbolic surface with geodesic flow $\phi_{t}$ and with $\lambda_{1}(X)$ being the smallest eigenvalue of the Laplace-Beltrami operator $\Delta$. Let $f,g\in L^{2}(X)$. Then for all $t\geq 0$,
$\bigg|\big\langle f\circ\varphi_{t},g\big\rangle_{L^{2}(T^{1}X)}-\int_{T^{1}X}f\,dx\int_{T^{1}X}g\,dx\bigg|\leq 11\,e^{\beta(\lambda_{1}(X))}\,(1+t)\,e^{-\beta(\lambda_{1}(X))t}\,\|f\|_{2}\,\|g\|_{2},$
where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4},\end{cases}$
and where $f(x)=f(x,\theta)$ and $g(x)=g(x,\theta)$ are viewed as functions on $T^{1}X$, the unit tangent bundle.
The more general phase space version (see [Rat87, Mat13]) implies that the geodesic flow on a compact connected hyperbolic surface is weak mixing with exponential decay which further implies ergodicity. By Theorem 1.1, a compact connected hyperbolic surface satisfies properties (i)(iii), and hence is both quantum ergodic and quantum weak mixing.
Remark 1.3.
From Theorem 1.1, we see that neither the $2$-torus nor the $2$-sphere is quantum ergodic or quantum weakly mixing. However, if we restrict the observables to be taken from the collection of the physical-space observables, i.e. the multiplicative operators corresponding to functions, then a $2$-torus with irrational aspect ratio is quantum ergodic (though still not quantum weak mixing). The $2$-sphere remains neither quantum ergodic nor quantum weak mixing. These examples can be analyzed explicitly due to the availability of closed-form expressions for their eigenfunctions and eigenvalues. An illustarting example showing this for a discrete circle with $N$ elements can be acquired using observable $a(x)=e^{\frac{2\pi i\,x}{N}}$; this was pointed out by Mostafa Sabri. This idea should translate to the $2$-dimensional surfaces as well. We see that the choice of observables matters.
These phenomena on Riemannian surfaces naturally motivate the search for analogous results in other contexts, such as discrete settings. A pseudo-analogue of the above corollary to Theorem 1.1 was obtained for certain finite regular graphs by Anantharaman and Le Masson [AL15]. Their result, however, is not a strict analogue, since the high-energy limit is not meaningful for finite graphs, which possess only finitely many eigenvalues. Instead, they considered a sequence $(G_{n})$ of graphs with increasing size $n$ and a fixed energy window $I$, as opposed to a fixed structure with increasing energy as in Shnirelman’s theorem. A special case of their result states that, under suitable assumptions, if $(a_{n})$ is a sequence of diagonal observables acting on functions on the vertices and if $(\Delta_{n})$ is a sequence of graph Laplacians each with eigenvalues $\{\lambda_{j}^{(n)}\}$ and eigenfunctions $\{\psi_{j}^{(n)}\}$, then
$\frac{1}{\#\{j:\lambda_{j}^{(n)}\in I\}}\sum_{\lambda_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{\operatorname{Tr}(a_{n})}{n}\right|^{2}\xrightarrow{n\to\infty}0.$
As in Shnirelman’s theorem, the averages of the expectation values converge to the uniform average. We refer to such results, in which the structure grows while energy remains bounded, as large-scale results, in contrast to the original results in which the energy frows. We refer to the latter results as large-energy results. The large-scale quantum ergodicity of graphs and related theorems is an active field with a vigorous interplay with the continuous settings. For more realted results in the discrete graph setting, see [DP12, BL13, EKYY13, TVW13, AL15, Gei15, BLL16, Ana17, AS17, BHKY17, BKY17, AS19, BS19, BHY19, McK22, MS23]. An especially impressive result establsihed in [BHKY17, BHY19, BKY17] shows a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] for random regular graphs; the randomization effectively suppresses pathological cases helping to establish stronger results compared to the deterministic setting.
Since the large-scale graph quantum ergodicity theorem is not a true analogue of Shnirelman’s theorem, it is natural to seek a corresponding surface result highlighting the distinctions from the original large-energy setting. Such an analogue for hyperbolic surfaces was established by Le Masson and Sahlsten [LS17] (see Subsection 1.2 for details). Their result resembles closely that of Anantharaman and Le Masson, with minor modifications reflecting the transition from discrete graphs to continuous surfaces.
Despite the distinct nature of the two limits, there exists a heuristic correspondence between the large-energy and large-scale settings. In the large-energy regime, as the energy grows, the eigenfunctions exhibit shorter wavelengths, resolving increasingly fine structures. In the large-scale regime, as the structure grows, finer details become accessible within a fixed energy window. Hence, these two asymptotic regimes may be regarded as analogous in spirit.
To our knowledge, the only existing large-scale off-diagonal results establishing analogues of properties (i), (ii), and (iii) of Theorem 1.1 (see Subsection 1.2) are those of Cipolloni, Erdős, and Schröder [CES21] for Wigner matrices. Their setting resembles that of Anantharaman and Le Masson, but with graphs replaced by random matrices and observables by general matrices. Their result is remarkably strong as they require no averaging: they prove that
$\left|\left\langle\psi_{j}^{(n)},A_{n}\psi_{k}^{(n)}\right\rangle-\delta_{jk}\tfrac{\operatorname{Tr}(A_{n})}{n}\right|$
decays to zero at the optimal and uniform rate of $n^{-1/2}$ for all $j,k$, thereby establishing a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94], and in fact extending it to the off-diagonal elements. This result is relevant as the random matrix theory is closely related to the quantum ergodicity on graphs and surfaces, especially if Berry’s conjecture (see [Ber77]) holds. For more related results for random matrices, see [ESY09a, ESY09b, BY17].
A new upcoming result by Bordenave, Letroit, and Sabri [BLS] establishes a quantum ergodicy and quantum mixing results on growing Schreier graphs where even the quantum ergodicy was previously unknown. To our knowledge, this is first mixing result in the graph setting. This is especially interesting as graphs are intuitively closer to the original setting on manifolds when compared to random matrices.
1.2. Main theorems
We prove a large-scale analogue of the corollary of Theorem 1.1 for physical-space observables, that is, for functions viewed as multiplicative operators, thereby complementing the result of Le Masson and Sahlsten [LS17]. We also prove a probabilistic analogue of this, complementing the probabilistic quantum ergodicity result of Le Masson and Sahlsten [LS24]. The small difference is that we use the quantum Hamiltonian $\sqrt{-\Delta-1/4}$ instead of $\sqrt{-\Delta}$ but causes only a spectral shift. We already refer to Section 3 for more technical overview of the proof ideas as well as the novel ideas of the proofs used.
The proof follows the ideas of Le Masson and Sahlsten in that we establish a quantitative version which yields limiting forms of the theorem. We again divide the argument into a spectral and a geometric component. However, we employ a different propagator, namely the hyperbolic wave propagator
$P_{t}=h_{t}\!\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right),\qquad h_{t}(x)=x^{-1}\sin(tx).$
The intuition behind $P_{t}$ is that $u_{tt}=-\Delta\,u$, the wave equation in the hyperbolic plane, has solution $u=2^{-1/2}\,P_{t}\,u_{0}$ with initial data
$u(z,0)=0,\qquad u_{t}(z,0)=2\pi f,$
for some function $f$. For details, see [LP82]. A convenient form of the integral kernel of $P_{t}$ is established in Section 5.
Unlike the ball-averaging operator used by Le Masson and Sahlsten, which has no immediate connection to the underlying classical dynamics, the hyperbolic wave propagator is obtained from the quantization of the geodesic flow, thus linking the quantum and classical settings in a natural manner. This choice of propagator renders the spectral side particularly clean, though it makes the geometric side more involved compared to the approach of Le Masson and Sahlsten.
To access the off-diagonal elements $\langle\psi_{j},a\psi_{k}\rangle$, where the energy levels of $\psi_{j}$ and $\psi_{k}$ may differ substantially, we also introduce a modified propagator: $P_{t,\tau}=\cos(t\tau)\,P_{t}$. Using $P_{t}$ and $P_{t,\tau}$ simultaneously allows us to capture the behaviour of the off-diagonal elements. This, however, complicates both the spectral and geometric analyses compared to the quantum ergodicity case, which concerns only the diagonal elements. In particular, the Heisenberg evolution of the observable $a$, $P_{t}^{*}aP_{t}$, appearing in the proof of quantum ergodicity by Le Masson and Sahlsten, is replaced here by $P_{t,\tau}^{*}aP_{t}$. The physical interpretation of this operator is not immediately transparent, but it can be verified that
$P_{t,\tau}^{*}aP_{t}=P_{t,\tau/2}^{*}aP_{t,\tau/2}+P_{-t,\tau/2}^{*}aP_{-t,\tau/2},$
meaning it equals the sum of two Heisenberg evolutions of $a$ with $P_{t,\tau/2}$, one forward and one backward in time.
Finally, we replace rather abstract Nevo’s theorem with the exponential mixing theorem (see Theorem 1.2), which elucidates the connection between classical and quantum dynamics. The idea of using the exponential mixing originates in the upcoming article of the author with Lequen, Mikkelsen, Sahlsten, and Ueberschär [HLM+]. The use of exponential mixing is both simple and transparent: the exponential decay of correlations of $a$ and $a\circ\varphi_{t}$, even for small $t$, is essential in obtaining our result. Small propagation times suffice because we work with observables in a macroscopic sense, unlike the large-energy quantum ergodicity and quantum weak mixing setting, which employs microlocal analysis and requires long propagation times for decorrelation to occur at the microlocal scale. The more technical approach taken in the proofs is presented in Section 3.
We now state the main theorems.
Theorem 1.4 (Large-scale quantum ergodicity and quantum mixing).
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
$0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
Assume that the sequence has the following properties:
  • (BSC)
    Benjamini–Schramm convergence: for every $R>0$,
    $\lim_{n\to\infty}\frac{\big|\{x\in X_{n}:\operatorname{InjRad}_{X_{n}}(x)<R\}\big|}{|X_{n}|}=0,$
  • (EXP)
    Expander property: the sequence admits a uniform lower bound on the spectral gap,
  • (UND)
    Uniform discreteness: the sequence admits a uniform lower bound on the injectivity radius.
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator. Then the following properties hold:
  • (i)
    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
  • (ii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon,$
  • (iii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon.$
Properties (i)(iii) are the large-scale analogues of Properties (i)(iii) in Theorem 1.1. Accordingly, we refer to systems satisfying Properties (i) and (ii) as large-scale quantum ergodic, and those satisfying (i)(iii) as large-scale quantum mixing. As in Theorem 1.1, the choice of observables plays a role, recall the example on the unit circle where quantum mixing fails discussed in the earlier remark, Remark 1.3. We note that this idea can be modified to prove that the $2$-torii of increasing area are not quantum mixing for physical-space observables, though there are sequences of $2$-torii that are, interestingly, quantum ergodic using the explicit formula of the eigenfunctions. We refer to the earlier illustrating example discussed This observation also illustrates that large-scale quantum ergodicity and large-scale quantum weak mixing are properties specific to certain sequences of hyperbolic surfaces rather than to general sequences of Riemannian manifolds.
Le Masson and Sahlsten [LS17] have already proven that property (i) holds in this setting. Hence to establish Theorem 1.4, it remains to prove properties (ii) and (iii). To prove the previous theorem, we need the following quantitative version of it.
Theorem 1.5 (Quantitative version).
Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
$0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
$0<\delta<\tfrac{2}{9}\min(I),\qquad T\delta=\tfrac{\pi}{2},$
and fix $\tau\in\mathbb{R}$. Then
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The coefficient appearing is
$C_{8}=50000\pi.$
Remark 1.6.
We observe from the statement of Theorem 1.5 that in the proof of Theorem 1.4 in Section 8, we could relax many of the conditions. E.g. we could let the spectral gap and the injectivity radius to shrink slowly as we proceed in the sequence. Thereby Theorem 1.4 could be extended to encompass a larger family of sequences.
An example of a sequence of surfaces that is both quantum ergodic and quantum mixing in the above sense is a sequence of compact arithmetic surfaces. This follows from Katz et al. [KSV07], who showed that the injectivity radius grows (ensuring properties (BSC) and (UND)), and from Kim et al. [KSR03], who established a uniform lower bound on the spectral gap (ensuring property (EXP)). The result of Kim et al. is the best deterministic bound currently available. The first, weaker, bound is Selberg’s $3/16$ bound, see [Sel65].
Moreover, the class of admissible sequences is quite broad which is seen when using probabilistic approach using the Weil-Petersson random model; for background, see [Mir13, Wri20]. By Le Masson and Sahlsten [LS24], for a sequence $(X_{g})$ of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is the Weil-Petersson probability measure on the moduli space $\mathcal{M}_{g}$ of hyperbolic surfaces of genus $g$, the condition (i) of Theorem 1.4 holds with high probability as $g$ grows to infinity. We state now a probabilistic analogue of Theorem 1.4, thus showing that a growing Weil-Petersson random surface sequence is quantum ergodic and quantum mixing with high probability and further complementing the results of Le Masson and Sahlsten [LS24].
Theorem 1.7 (Probabilistic version).
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus. Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
$0=\lambda_{0}^{(g)}<\lambda_{1}^{(g)}\leq\dots\leq\lambda_{j}^{(g)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator. Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
As mentioned, Property (i) was proved by Le Masson and Sahlsten [LS24]. Hence it remains to prove Properties (ii) and (iii) where the appraoch of the proof follows closely that of Theorem 1.4. However, as the assumptions are lighter in the probabilistic setting we do not have Properties (BSC), (EXP), and (UND) available, nor do we have the large-scale Weyl law, Theorem 8.1. We need analogues of these in the proof of Theorem 1.7. Theorem 8.2 consisting of recent results in the random surface theory offers us the required tools.
The use of Weil-Petersson random surfaces in quantum ergodicity is currently very active. Randomization effectively suppresses pathological cases helping to establish stong results that the analogous deterministic settings are far away from acquiring. For results in Benjamini-Schramm convergence and injectivity radius in Weyl-Peterson random model, see [Mir13, Mon21]. For recent results on the spectral gap, see [AM24, LW24, Mag24, HMT25, HMN25, WX25]. There are other random models and we especially refer to late groundbreaking results using the random covering model, see [MN20, MNP22, HM23]. See an interesting application of Property (EXP) to the Bose-Einstein condensation of random hyperbolic surfaces [LS22].
1.3. Further discussions
Like Le Masson and Sahlsten [LS17], we restrict the energy window to a compact interval in $\left(1/4,\infty\right)$. This assumption plays a crucial role in our proofs. However, it is plausible that with a different treatment of the low-lying modes—namely, those eigenvalues and corresponding eigenfunctions with eigenvalues in [0,1/4)[0,1/4)—one could extend the results to an energy window covering the full range $0,\infty)$.
In contrast to Theorem 1.1, Theorem 1.4 is not a characterization-type result linking the classical and quantum dynamics. A natural question arises: can Theorem 1.4 be complemented to yield such a characterization? As noted in Remark 1.6, the sequences of surfaces may in fact possess slowly decaying injectivity radii and spectral gaps. Given the connection between weak mixing and the spectral gap, it is plausible to conjecture that a sequence of hyperbolic manifolds converging to a hyperbolic space in the Benjamini–Schramm sense is large-scale quantum weak mixing if and only if the spectral gap tends to zero sufficiently slowly. Establishing such characterizations for both large-scale quantum ergodicity and large-scale quantum mixing would be remarkable.
While full characterization results may still be out of reach, the present results appear extendable—albeit with varying levels of difficulty. Several directions for possible extensions present themselves. For instance, can analogous results be established in other structures such as graphs? We conjecture that the graph result of Brooks, Lindenstrauss, and Le Masson [BLL16] could be extended to include off-diagonal elements by adapting the techniques developed here, thereby enabling a proof of large-scale weak mixing in the graph setting of Anantharaman and Le Masson. Doing quantum mixing on the graphs seems highly timely as Bordenave, Letroit, and Sabri have an upcoming result on mixing on growing Schreier graphs [BLS].
Another avenue of extensions concerns the choice of observables: in Theorem 1.1, the observable class was quite broad—zeroth-order pseudodifferential operators—so it seems likely that with minor modifications, a larger class of observables could be accommodated in the large-scale setting as well. This has already been achieved in the large-scale quantum ergodicity contexts for graphs [AL15] and surfaces [ABL23], where the kernels of the observables are concentrated near the diagonal but not restricted to it, unlike in the present work.
Yet another natural question is whether the theorem can be adapted to the case of variable curvature, or to accommodate more general energy operators than the Laplacian. For graphs, analogous results are known: Anantharaman and Sabri [AS17, AS19] established large-scale quantum ergodicity for non-regular graphs with general Schrödinger operators, under suitable restrictions. Non-regular graphs can be viewed as discrete analogues of variable-curvature surfaces. It would also be of great interest to construct counterexamples—sequences of surfaces demonstrating that the conditions imposed on the sequences are indeed necessary. In the graph setting, the results of Anantharaman and Le Masson were complemented in this manner by MacKenzie [McK22]; however, constructing such examples in the surface setting remains substantially more difficult.
We pose perhaps an obvious question: is it possible to remove the averaging present in our theorems? In the large-energy regime, this corresponds to the quantum unique ergodicity conjecture of Rudnick and Sarnak, see [RS94]. The conjecture has been resolved in certain special cases (notably by Lindenstrauss for arithmetic surfaces [Lin06]), but remains open in general. Analogous large-scale results are known for graphs; for instance, large-scale quantum unique ergodicity for random regular graphs was proved by Bauerschmidt et al. [BHY19]. We conjecture that the conclusions of Theorem 1.4 remain valid even without averaging. We refer to such statements as large-scale quantum unique ergodic and large-scale quantum unique mixing. These would parallel the averaging-free results of Cipolloni et al. [CES21]. A possibility is to try to establish this first using random models, such as Weil-Petersson model, or by using arithmetic surfaces and thus complement Lindenstrauss [Lin06]. In particular, for arithmetic surfaces, averaging-free off-diagonal bounds are of significant interest: strong estimates on $|\langle\psi_{j},\psi_{k}\psi_{l}\rangle|$ could contribute to a proof of the Lindelöf hypothesis, to which quantum chaos is elegantly connected via the Watson $L$-function [Wat02].
In any case, a deeper understanding of averaging-free results may provide a pathway toward results in many-body quantum chaos—such as the Eigenstate Thermalization Hypothesis proposed by Deutsch in the physics literature [Deu18]. Deutsch predicts that $|\langle\psi_{j},a\psi_{k}\rangle|$ goes to zero as the distance of the eigenvalues of $\psi_{j}$ and $\psi_{k}$ goes to infinity. Looking at the large-scale results discussed here, Theorems 1.4 and 1.7, we see that parameter $\tau$ controls the distance of the eigenvalues. However, the eigenvalues are taken from the fixed energy windows which means that for large enough $\tau$, the averages of the off-diagonal elements will be zero as the large distance is not possible in the energy window. However, if we let the energy window grow, then $\tau$ can also grow without giving trivial conclusions. It is natural to assume that in this setting of growing energy window and growing $\tau$, the average will close to zero, but verifying this requires careful study of the sequence of growinf surfaces, $\tau$, and the energy window.
2. Preliminaries
In this section, we recall some basic concepts and results that will be used throughout this article, without proofs. For background on hyperbolic geometry concerning the hyperbolic plane and hyperbolic surfaces, we refer to Katok [Kat92], and for details on the Laplace–Beltrami operator, to Iwaniec [Iwa02].
2.1. The hyperbolic plane and hyperbolic surfaces
The hyperbolic plane $\mathbb{H}$ is identified with the upper half-plane
$\{x+iy\in\mathbb{C}:y>0\},$
equipped with the Riemannian metric
$ds^{2}=y^{-2}(dx^{2}+dy^{2}).$
For points $z,z^{\prime}\in\mathbb{H}$, we denote by $d(z,z^{\prime})$ the hyperbolic distance between them. The associated volume measure is
$d\mu(x+iy)=y^{-2}\,dx\,dy,$
and we write $|A|$ for the hyperbolic volume of a set $A\subseteq\mathbb{H}$.
The group $PSL_{2}(\mathbb{R})$, consisting of $2\times 2$ real matrices with determinant one, modulo $\pm I$, acts on $\mathbb{H}$ by Möbius transformations:
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot z=\frac{az+b}{cz+d}.$
This action identifies $PSL_{2}(\mathbb{R})$ with the group of orientation-preserving isometries of $\mathbb{H}$.
The unit tangent bundle of $\mathbb{H}$ is
$T^{1}\mathbb{H}=\{(z,v)\in\mathbb{H}\times T_{z}\mathbb{H}:\|v\|_{z}=1\}.$
The group $PSL_{2}(\mathbb{R})$ acts on $T^{1}\mathbb{H}$ via
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot(z,v)=\left(\frac{az+b}{cz+d},\frac{v}{(cz+d)^{2}}\right).$
Evaluating this action at the point $(i,i)$ shows that $PSL_{2}(\mathbb{R})$ is homeomorphic to $T^{1}\mathbb{H}$; hence, these two spaces can be identified.
The geodesic flow is the one-parameter family of maps
$\varphi_{t}:T^{1}\mathbb{H}\to T^{1}\mathbb{H},\qquad t\in\mathbb{R},$
such that for any $(z,v)\in T^{1}\mathbb{H}$, the curve $(\eta(t),\eta^{\prime}(t))=\varphi_{t}(z,v)$ is the unique geodesic parametrized by arc length satisfying $\eta(0)=z$ and $\eta^{\prime}(0)=v$. The geodesic
$(i,i)\mapsto\varphi_{t}(i,i)=\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}\cdot(i,i)$
corresponds to the imaginary axis. For any $(z,v)\in T^{1}\mathbb{H}$, there exists $g\in PSL_{2}(\mathbb{R})$ such that $g\cdot(i,i)=(z,v)$, and thus
$\varphi_{t}(z,v)=g\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}\cdot(i,i).$
Hence, identifying $T^{1}\mathbb{H}$ with $PSL_{2}(\mathbb{R})$, the geodesic flow corresponds to the right action of the one-parameter subgroup
$A=\left\{a_{t}=\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}:t\in\mathbb{R}\right\}.$
We denote by $z,z^{\prime}]$ the geodesic segment connecting two distinct points $z,z^{\prime}\in\mathbb{H}$, by $z,z^{\prime}](w)$ the tangent direction at a point $w\in[z,z^{\prime}]$, and by $|[z,z^{\prime}]|$ its length.
For any distinct points $z,z^{\prime}\in\mathbb{H}$, there exists a unique geodesic of length $r$ from $z$ to $z^{\prime}$. Using the geodesic flow, there exist unique $\theta\in\mathbb{S}^{1}$ and $r\in(0,\infty)$ such that $z^{\prime}$ is the projection of $\varphi_{r}(z,\theta)$ onto the first coordinate. The change of variables $z^{\prime}\mapsto(r,\theta)$ defines the polar coordinates around $z$, in which the metric and volume element are
$ds^{2}=dr^{2}+\sinh^{2}(r)\,d\theta^{2},\qquad d\mu(r,\theta)=\sinh(r)\,dr\,d\theta.$
Then a point $x\in\mathbb{H}$ can be written as
$x=\pi_{1}(\phi_{r}(z,\theta)),$
where the polar coordinates are around $z$; $r$ and $\theta$ depend on $x$, and $\pi_{1}$ is the projection onto the first element of pair in $T^{1}\mathbb{H}$.
The Laplacian on $\mathbb{H}$ is the differential operator
$\Delta=-y^{2}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right),$
which commutes with all isometries of $\mathbb{H}$.
2.2. Hyperbolic surfaces
A compact connected hyperbolic surface can be realized as a quotient
$X=\Gamma\setminus\mathbb{H},$
where $\Gamma\subset PSL_{2}(\mathbb{R})$ is a discrete subgroup consisting solely of hyperbolic elements, that is, matrices whose trace satisfies $|\mathrm{tr}(\gamma)|>2$. In this way, $X$ can be represented by a fundamental domain in $\mathbb{H}$.
A function on $X$ can equivalently be regarded as a $\Gamma$-periodic function on $\mathbb{H}$, and vice versa. Moreover, the integral of a function $f$ over $X$ equals the integral of the corresponding $\Gamma$-invariant function over a fundamental domain.
For a hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, the unit tangent bundle can be identified with $\Gamma\setminus PSL_{2}(\mathbb{R})$, and under this identification the geodesic flow takes the form
$\varphi_{t}(\Gamma g)=\Gamma g\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}$.
The injectivity radius of $X$ at a point $z\in X$ is defined by
$\operatorname{InjRad}_{X}(z):=\tfrac{1}{2}\min_{\gamma\in\Gamma\setminus\{\mathrm{id}\}}d(z,\gamma z),$
and the (global) injectivity radius of $X$ is
$\operatorname{InjRad}_{X}:=\sup_{z\in X}\operatorname{InjRad}_{X}(z).$
Since the Laplacian on $\mathbb{H}$ commutes with isometries, it descends naturally to a Laplacian on the quotient $X=\Gamma\setminus\mathbb{H}$. To distinguish the two, we denote by $\Delta_{\mathbb{H}}$ the Laplacian acting on all smooth functions on $\mathbb{H}$, and by $\Delta_{X}$ the Laplacian acting on $\Gamma$-invariant functions. On a compact connected hyperbolic surface, the Laplacian $\Delta_{X}$ has a discrete spectrum
$0=\lambda_{0}<\lambda_{1}\leq\lambda_{2}\leq\dotsb\xrightarrow{j\to\infty}\infty.$
3. Strategy of proofs
Three main theorems, Theorems 1.4, 1.5, and 1.7, are proven in Section 8. In this section we give a rough, overarching idea of the approach taken in the proofs. Most of the paper is dedicated for trying to establish Theorem 1.5. Proving this theorem is technical so the proof is split into multiple sections. These sections, Sections 4, 5, 6, and 7, are presented before the Section 8 so that the auxiliary results have been covered when the main theorems are under focus.
In Theorem 1.5 we want to bound the double sum appearing in Expression (3.3) for a fixed surface $X=\Gamma\setminus\mathbb{H}$ under suitable conditions. Hence we want to bound the following:
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}$ (3.1)
We can split the inner product as follows:
$|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,\psi_{k}\,dt\right\rangle\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|},$ (3.2)
when
$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right),\qquad h_{t}(x)=\frac{\sin(tx)}{x},$
and
$P_{t,\tau}=h_{t,\tau}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right),\qquad h_{t,\tau}=\cos(t\tau)\,h_{t}.$
This split could be referred to as geometric-spectral split, since the denumerator contains the spectral data of the problem and the numerator the geometric data of the problem. $P_{t}$ is the hyperbolic wave propagator (see [LP82]). It is the quantization of the geodesic flow, hence a very natural object. Using this propagator makes the spectralside easy but complicates the analysis of the geometric side when comparing the setting to that of Le Masson and Sahlsten (see [LS17]) who used a ball averaging propagator which is not as natural as the wave propagator. $P_{t,\tau}$, a modified version of $P_{t}$, is used simultaneously with $P_{t}$ so that we can access the off-diagonal elements. This adds another layer of complexity to the proof when comparing to the setting in which only the diagonal elements are considered.
The denumerator of Expression (3.2) contains the spectral data of the problem. We acquire bounds for the denominator in Section 4. Using the bounds acquired there gives an upper bound for Expression (3.1):
$64\pi^{2}\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2},$
where the Hilbert-Schmidt norm upper bound was acquired by enlargening the sums to include all the basis elements after the spectral bound was pulled out of the sum. Now the Hilbert-Schmidt norm contains the geometric data of the problem.
Bouding the Hilbert-Schmidt norm is very technical. In Section 5 we establish convenient forms for the integral kernels of $P_{t}$ and $P_{t,\tau}$. Using the acquired formulas, we bound the Hilbert-Schmidt norm in Section 6. This is based on carefully bounding the Hilbert-Schmidt norm expression and finally splitting it into two cases. In one we consider an integral over points with small injectivity radius and in the other an integral over points with large injectivity radius. To bound these two cases, we especially need the exponential mixing theorem, Theorem 1.2, and bounds for certain geometric weight $F_{t,t^{\prime},\rho}$ which appears in both the small and large injectivity radius cases, containing the core geometric information of the problem. Bounding $F_{t,t^{\prime},\rho}$ is done in Section 7. We remark that the exponential mixing theorem replaces the Nevo’s theorem appearing in the same role in [LS17] as the exponential mixing here.
With the Hilbert-Schmidt norm bounded, we acquire the final bounds for Expression (3.1) to conclude Theorem 1.5:
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq E\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where we refer to Theorem 1.5 for the quantities appearing in the upper bound.
The limit version of the main theorems, Theorem 1.4, is proven using Theorem 1.5 together with the large-scale Weyl law of Le Masson and Sahlsten, Theorem 8.1. In the proof we want to prove that for any $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for any $\tau\in\mathbb{R}$ we have that
$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle|^{2}\right)<\epsilon.$ (3.3)
Proving this proves only conditions (ii) and (iii); Le Masson and Sahlsten’s proof of quantum ergodicity covers case (i), see [LS17]. We remark here that the techniques used here can also be used to give a new proof for the quantum ergodicity theorem. Using the bounds acquired for the double sum in Theorem 1.5 gives the following upper bound for the previous expression:
$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,E\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}(X_{n}))^{3}}\,\left(\|a_{n}\|_{2}^{2}\,+\,8\pi^{2}\|a_{n}\|_{\infty}^{2}\,|D_{n}\setminus D_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right)\right),$
where we refer to Theorem 1.4 for the quantities appearing in the upper bound. Using property (BSC) helps us to get a universal lower bound for $\beta(\lambda_{1}(X_{n}))$ and using property (UND) helps us to get a universal lower bound for $\operatorname{InjRad}_{X_{n}}$. Weyl law together with property (BSC) shows that
$\frac{|D_{n}\setminus D_{n}(4T)|}{|N(X_{n},I)|}\xrightarrow{n\to\infty}0,$
as the Weyl law relates the growth of the number of eigenvalues in the energy window $I$ to that of the volume of the surface, $|X_{n}|$, and the ratio of the points with small injectivity radius and the volume of the surface goes to zero due to property (BSC). The core idea is to use properties (BSC), (EXP), and (UND) together with Weyl law to remove the limit supremum from the upper bound. After this, we have an upper bound in which $\delta$ appears. We can choose a suitable $\delta$ to get $\epsilon$ as an upper bound to conclude the proof.
To finally prove the probabilistic main theorem, Theorem 1.7, we take a very similar approach as with the proof of Theorem 1.4 as these theorems are very similar. We only prove Properties (ii) and (iii) of Theorem 1.4 as Le Masson and Sahlsten have already proved Property (i), see [LS24]. We remark that in probabilistic setting we do not impose Properties (BSC), (EXP), and (UND) and hence we do not have the large-scale Weyl law at out disposal. To remedy this, we get similar properties from the theory of random surfaces. These are recorded in Theorem 8.2. We can use these properties in the place of Properties (BSC), (EXP), and (UND) in a very similar way to get the conclusion to hold with high probability as we proceed in the sequence.
To conclude the section, we remark that the main novelty of this paper method-wise is the use of the exponential mixing theorem, Theorem 1.2, instead of using Nevos’s theorem to get the needed extra decay to establish Theorem 1.5. Another novelty is to separate a weight arising from the choice of propagator, denoted here by $F_{t,t^{\prime},\rho}$, which gives much required nimbleness in the proof; it can also be approximated separately without the observable function $a$ interfering.
4. Spectral data
In this section we prove a bound for spectral quantity
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}$
under certains assumptions on $a,b,t,\tau,$ and $T$, given that
$h_{t}(x)=x^{-1}\,\sin(tx),\qquad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x).$
The spectral quantity appears in the proof of Theorem 1.5 and there it contains the spectral data of the problem.
We state the proposition.
Proposition 4.1 (Spectral data).
Let $\tau\in\mathbb{R}$. Assume that $0<m<a,b<\infty$ and that
$a-b-\tau\in(-\delta,\delta),$
where $0<\delta<\tfrac{2}{9}\,m$ and where $T\,\delta=\pi/2$. Then
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}\,<\,8\pi ab.$
The proof consists of relatively straightforward integration and careful application of the relationships of the variables and coefficients. The simplicity follows from a convenient choice of $h_{t}$ and $h_{t,\tau}$.
Proof of Proposition 4.1.
We set out to bound
$G_{\tau,T}=\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|$
from below to prove the bound. $G_{t,\tau}$ is illustrated in Figure 1.
Figure 1. Contour plot of $G_{\tau,T}(a,b)$. Two pronounced “ridges” are visible along the lines $b=a+\tau$ and $b=a-\tau$. The ridges get pronounced as $T$ increases. For a fixed $\delta$, $G_{\tau,t}(a,b)$ has a positive lower bound when $a-b-\tau\in(-\delta,\delta)$. Notice also how the ridges decay as $a$ and $b$ grow.
By using the definitions of $h_{t}$ and $h_{t,\tau}$, we write $G_{\tau,T}(a,b)$ as
$\frac{1}{Tab}\left|\int_{0}^{T}\cos(t\tau)\sin(ta)\sin(tb)dt\right|.$ (4.1)
By using trigonometric sum-to-product identities, we have that $\cos(t\tau)\sin(ta)$ equals
$\frac{1}{2}\left(\sin\left(t(\tau+a)\right)-\sin\left(t(\tau-a)\right)\right),$
and further that $\cos(t\tau)\sin(ta)\sin(tb)$ equals
$\frac{1}{4}\left(\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\right).$
Using this equality, we have that Expression (4.1) equals
$\frac{1}{4Tab}\left|\int_{0}^{T}\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\,dt\right|.$
The sine terms of the integral can all be integrated separately and thus the integral yields the following equal expression:
$\frac{1}{4Tab}\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}-\frac{\sin(T(a+b-\tau))}{a+b-\tau}-\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|$
Applying the reverse triangle inequality yields a following lower bound for the previous expression:
$\frac{1}{4Tab}\bigg(\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}\right|-\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|-\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|\bigg).$ (4.2)
We proceed to bound the absolute value terms separately. Consider the first absolute value term of the previous expression. Consider function
$x\mapsto\frac{\sin(T\,x)}{x}.$
It is positive when $x\in\left(-\pi/T,\pi/T\right)$, and it also decreases as $|x|$ grows in this interval. Because $T\delta=\pi/2$ and $a-b-\tau\in(-\delta,\delta)$, we can apply the observations regarding the previously studied function to get that
$\frac{\sin\!\big(T(a-b-\tau)\big)}{a-b-\tau}\ \geq\ \frac{\sin(T\delta)}{\delta}=\frac{1}{\delta}.$
We also notice that
$\frac{\sin(Tx)}{x}>0$
when $x\in\left(-\pi/T,\pi/T\right)$, which means that for every $x\in\mathbb{R}$ we have that
$\frac{\sin(Tx)}{x}\geq-\frac{1}{\pi/T}=-\frac{1}{2\delta}.$
The previous two observations mean that
$\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}$
is necessarily positive and hence
$\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}\right|=\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}>\frac{1}{2\delta}.$
Next, consider the second and third absolute value terms of Expression (4.2). We observe that
$\left|\frac{\sin(Tx)}{x}\right|\leq\frac{1}{|x|},$
implying that
$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|\leq\frac{1}{|a+b-\tau|}.$
We also observe that
$|a+b-\tau|\geq 2b-|a-b-\tau|>2m-\delta>8\delta,$
where we used the reverse triangle inequality together with the fact that $|a-b-\tau|$ is smaller than $\delta$ and that $2m$ is larger than $9\delta$. Using these observations gives that
$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|\leq\frac{1}{|a+b-\tau|}<\frac{1}{8\delta}.$
A similar approach yields that
$\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|<\frac{1}{8\delta}.$
These two observations mean that
$-\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|-\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|>-\frac{1}{4\delta}.$
With all the absolute value terms bounded, Expression (4.2) has the following lower bound:
$\frac{1}{4Tab}\,\frac{1}{4\delta}=\frac{1}{16\frac{\pi}{2}ab}=\frac{1}{8\pi ab}.$
In the previous we used that $T\delta=\pi/2$.
Combining these bounds gives that
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|\ \geq\ \frac{1}{4Tab}\cdot\frac{1}{4\delta}\ =\ \frac{1}{16\,\frac{\pi}{2}\,ab}\ =\ \frac{1}{8\pi\,ab}.$
Therefore,
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}=G_{\tau,T}^{-1}<8\pi ab,$
as claimed. ∎
5. Propagators
In order to estimate the Hilbert–Schmidt norm
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2},$
the geometric data of the problem that appears in the proof of Theorem 1.5, it is necessary to analyze the kernels of the hyperbolic wave propagator $P_{t}$ and a slightly mofified version of it, $P_{t,\tau}$, for these kernels are then used to open the Hilbert-Schmidt norms. Although the hyperbolic wave propagator is a natural choice, it exhibits greater thechnical complexity than, for instance, the ball averaging operator (cf. [LS17]). As mentioned earlier, this choice makes the specral side easier.
We state a proposition giving a convenient form for the integral kernels of $P_{t}$ and $P_{t,\tau}$.
Proposition 5.1 (Propagators).
Let
$P_{t}=h_{t}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),\quad P_{t,\tau}=h_{t,\tau}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),$
be defined on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$ with $\Delta_{X}$ being the Laplace-Beltrami operator and where
$h_{t}(x)=\frac{\sin(tx)}{x},\quad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x).$
Then $P_{t}$ admits integral kernel
$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\quad K_{t}(x,y)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
and $P_{t,\tau}$ admits integral kernel $\cos(t\tau)\,K_{t}^{\Gamma}$.
Given that
$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right)$
and that $\tilde{P}_{t}$ is the operator with the integral kernel $K_{t}^{\Gamma}$ which was presented in the statement of the proposition, the proposition can be proved by showing that inner product $\langle P_{t}f,g\rangle$ equals inner product $\langle\tilde{P}_{t}f,g\rangle$ where $f,g\in C^{\infty}(X)$ as $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ with $X$ being compact. The core idea is representing $P_{t}$ as a limit of propagators $P_{t}^{(\varepsilon)}$ that have a convenient explicit formulae for their integral kernels.
The proof relies on multiple auxiliary results presented after the proof of Proposition 5.1 for clarity’s sake
Proof of Proposition 5.1.
The assertion for $P_{t,\tau}$ follows directly from the assertion for $P_{t}$ since $P_{t,\tau}=\cos(t\tau)P_{t}$. Hence it suffices to consider $P_{t}$ and $K_{t}$.
Due to way $P_{t}$ is defined, it is bounded and has property
$P_{t}\psi_{j}=h_{t}(\rho_{j})\,\psi_{j},$
where $\psi_{j}$ is an eigenfunction of the Laplacian on $X$ with eigenvalue $\lambda_{j}$ such that $\rho_{j}=\sqrt{\lambda_{j}-1/4}$. Define regularized operators for $\varepsilon>0$:
$P_{t}^{(\varepsilon)}=h_{t}^{(\varepsilon)}\Big(\sqrt{-\Delta-\tfrac{1}{4}}\Big),\quad h_{t}^{(\varepsilon)}(x)=\frac{\sin(tx)}{x}\,e^{-\frac{\varepsilon^{2}x^{2}}{2}}.$
Due to way $P_{t}^{(\epsilon)}$ is defined, it is bounded and has property
$P_{t}^{(\varepsilon)}\psi_{j}=h_{t}^{(\varepsilon)}(\rho_{j})\,\psi_{j}.$
Since $h_{t}^{(\varepsilon)}$ converges to $h_{t}$ pointwise as $\varepsilon$ shrinks, we have that $P_{t}^{(\varepsilon)}f$ converges to $P_{t}f$ pointwise as $\varepsilon$ shrinks for every $f\in C^{\infty}(X)$.
Proposition 5.2 ensures the existence of integral operator $Q_{t}^{(\varepsilon)}$ on $X$ with integral kernel
$K_{t}^{(\varepsilon)}(x,y)=\sum_{\gamma\in\Gamma}k_{t}^{(\varepsilon)}(x,\gamma y),$
$k_{t}^{(\varepsilon)}(x,y)=-\frac{1}{\pi i}\frac{1}{2\sqrt{2}\pi}\int_{-\infty}^{\infty}\Bigg(\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\Bigg)\rho h^{(\varepsilon)}(\rho)\,d\rho,$
such that
$Q_{t}^{(\varepsilon)}\psi_{j}=h_{t}^{(\varepsilon)}(\rho_{j})\psi_{j}.$
Lemma 5.3 allows us to rewrite $k_{t}^{(\varepsilon)}$ in a more tractable form so that
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where
$\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}$
denotes the standard Gaussian mollifier. Since an operator on $X$ is uniquely determined by its action on the orthonormal basis, it follows that $Q_{t}^{(\varepsilon)}=P_{t}^{(\varepsilon)}$.
Let $\tilde{P}_{t}$ denote an integral operator on $X$ with kernel $K_{t}^{\Gamma}$. To show that $\tilde{P}_{t}$ is well-defined and bounded, we use the Schur test (see [HS12]), so it suffices to verify that for all $x\in X$
$\int_{X}K_{t}(x,y)\,dy$
is finite as $K_{t}^{\Gamma}(x,y)$ is symmetric in $x$ and $y$. Open the previous expression using the definition of $K_{t}$:
$\frac{1}{2\sqrt{2}\pi}\,\int_{X}\sum_{\gamma\in\Gamma}\frac{\bm{1}_{t>d(x,\gamma y)}}{\sqrt{\cosh(t)-\cosh(d(x,\gamma y))}}dy.$
Uniting the integral and the sum into an integral over the universal cover gives the following equal expression:
$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}dy.$
Writing this using polar coordinates around $x$ yields
$\frac{2\pi}{2\sqrt{2}\pi}\int_{0}^{t}\frac{\sinh(\rho)}{\sqrt{\cosh(t)-\cosh(\rho)}}d\rho=\frac{4\pi}{2\sqrt{2}\pi}\sqrt{\cosh(t)-1},$
where coefficient $2\pi$ appeared first from the polar coordinates since there was no dependency on the direction, and the equality followed from computing the integral. The result is finite and $\tilde{P}$ is well-defined and bounded.
To show that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, as claimed in the statement of the proposition, it suffices to verify that for every $f,g\in C^{\infty}(X)$
$\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle=0\iff\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ where $X$ is compact. To show this, start by considering the absolute value term of the previous expression. It can be written as
$\left|\int_{X}g(x)^{*}\big(P_{t}f(x)-\tilde{P}_{t}f(x)\big)dx\right|,$
where we opened the inner product and united the two integrals. By pulling the absolute values inside the integral and then bounding $|g|$ with the $\sup$ norm, we get the following upper bound for the previous expression:
$\|g\|_{\infty}\int_{X}\left|P_{t}f(x)-\tilde{P}_{t}f(x)\right|dx.$
Focus on the absolute value integrand appearing in the previous expression. We can write it in the following form by writing $P_{t}f$ using the limit and then by writing the operator action using the kernels:
$\left|\lim_{\varepsilon\to 0}\int_{X}K_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}^{\Gamma}(x,y)f(y)dy\right|.$
Writing $K_{t}^{\Gamma}$ and $K_{t}^{(\varepsilon)}$ using their sum representations over $\Gamma$ and then uniting the integral and the sum into an integral over the universal cover yields the following equal expression, where we also view $f$ as $\Gamma$ periodic function:
$\left|\lim_{\varepsilon\to 0}\int_{\mathbb{H}}k_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}(x,y)f(y)dy\right|.$
By pulling the absolute values inside the integral and bouding $|f|$ with the $\sup$ norm, we get the following upper bound:
$\|f\|_{\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|k_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}(x,y)\right|dy.$
Focus on the integral in the previous expression. By opening $k_{t}$ and $K_{t}$, the integral can be written as
$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\left|\left(\int_{\mathbb{R}}\frac{\left(\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)\right)\,\bm{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}dv\right)-\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\right|dy.$
By defining
$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
we may write the preceding expression as
$\frac{1}{2\sqrt{2}\pi}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(t)+\chi_{d(x,y)}(-t)\right|\,dy,$
where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$. The additional term $\chi_{d(x,y)}(-t)$ equals 0 so adding it had no effect on the equality. Using the triangle inequality, the previous expression is bounded by
$\frac{1}{2\sqrt{\pi}}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|+\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(-t)\right|\,dy.$
Finally, Lemma 5.4 guarantees that $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. The line of reasoning leading to this cocnlusion means that
$\langle P_{t}f,g\rangle=\langle\tilde{P}_{t}f,g\rangle,$
thus concluding the proof. ∎
The remaining section contains the lemmas and propositions used in the proof of Proposition 5.1. First, we record a slightly modified formulation of Proposition 5 in [Mar04]. It gives a family of operators with explicit formulae for their integral kernels on the compact connected hyperbolic surface. Especially, these operators share the eigenfunctions with the Laplacian of the surface and of which eigenvalues have a representation using the eigenvalues of the Laplacian.
Proposition 5.2.
Let $h:\{x+iy:|y|<C\}\to\mathbb{C}$ satisfy:
  1. i.
    $h$ is analytic,
  2. ii.
    $h$ is even,
  3. iii.
    $h(x+iy)$ decays superpolynomially as $|x|\to\infty$,
[…]
$K^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}K(x,\gamma y),$
where
$K(x,y)=-\frac{1}{\pi i}\frac{1}{2\sqrt{2}\pi}\int_{-\infty}^{\infty}\left(\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\right)\rho h(\rho)\,d\rho,$
satisfies property
$P\psi_{j}=h(\rho_{j})\psi_{j},$
where $\psi_{j}$ is a Laplacian eigenfunction on $X$ with eigenvalue $\lambda_{j}$ and $\rho_{j}=\sqrt{\lambda_{j}-1/4}$.
Next, we derive a convenient expression for $k_{t}^{(\varepsilon)}$, an integral kernel that is crucial in establishing a convenient formula for the integral operator $P_{t}$.
Lemma 5.3.
Let
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{\pi i}\int_{-\infty}^{\infty}\left(-\frac{1}{2\pi\sqrt{2}}\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\right)\rho h_{t}^{(\varepsilon)}(\rho)\,d\rho,$
where $x,y\in\mathbb{H}$ and
$h_{t}^{(\varepsilon)}(\rho)=\frac{\sin(t\rho)}{\rho}\,e^{-\frac{\varepsilon^{2}\rho^{2}}{2}}.$
Then
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where
$\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}$
is the standard Gaussian mollifier.
The proof is based on noting that there is a Fourier transform inside the expression used to define $k_{t}^{(\varepsilon)}$ in the statement. Then by using the inverse Fourier transform of $\phi_{\varepsilon}$ and the Fourier transform of $s_{t}(x)=x^{-1}\,\sin(t\,x)$ (in the distributional sense) together with the Fourier convolution identity, we acquire the wanted conclusion.
Proof of Lemma 5.3.
Applying Fubini’s theorem to interchange the order of integration, we obtain that $k_{t}^{(\varepsilon)}$ equals
$\frac{1}{\pi i}\frac{1}{2\sqrt{2}\,\pi}\int_{d(x,y)}^{\infty}\frac{1}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\left(\int_{-\infty}^{\infty}e^{-i\rho v}\rho h_{t}^{(\varepsilon)}(\rho)\,d\rho\right)dv.$ (5.1)
The inner integral is the Fourier transform of $\rho\mapsto\rho h_{t}^{(\varepsilon)}(\rho)$, where the Fourier transform is
$\mathcal{F}(f)(x)=\int_{\mathbb{R}}f(y)e^{-ixy}\,dy$
and hence the Inverse Fourier transform is
$F^{-1}f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}f(y)\,e^{ixy}dy.$
We observe that
$\rho\,h_{t}^{(\varepsilon)}(\rho)=\sin(t\rho)\,e^{-\frac{\varepsilon^{2}\rho^{2}}{2}}=2\pi\,\sin(t\rho)\,\mathcal{F}^{-1}(\phi_{\varepsilon})(\rho),$
since
$\mathcal{F}^{-1}(\phi_{\varepsilon})(x)=\frac{1}{2\pi}\,e^{-\frac{\varepsilon^{2}\,x^{2}}{2}}.$
By using the Fourier convolution identity and the Fourier transform of sine we observe that the inner integral of Expression (5.1) equals
$\mathcal{F}\left(\rho\mapsto 2\pi\,\sin(t\rho)\,\mathcal{F}^{-1}(\phi_{\varepsilon})\right)(v)=\frac{2\pi}{2\pi}\left(\mathcal{F}(s_{t})\ast\phi_{\varepsilon}\right)(v)=\pi i\left(\phi_{\varepsilon}(v+t)-\phi_{\varepsilon}(v-t)\right),$
where $s_{t}$ denotes function $x\mapsto\sin(tx)$. Substituting this into Expression (5.1) yields the following equal expression:
$\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(v-t)-\phi_{\varepsilon}(v+t)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(t+v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where the equality followed using the symmetry of the Gaussian mollifier.
Finally, the following lemma establishes convergence of $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. Interestingly, the convolution here is over the real line but the convergence is studied in the $\mathbb{H}$.
Lemma 5.4.
Let
$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}$
and let
$\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}.$
Then map $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$ for every $t\in\mathbb{R}$ as $\varepsilon\to 0$, where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$.
In the proof we show that
$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy\xrightarrow{\varepsilon\to 0}0.$
The idea is to present this integral as a convolution $\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)$, where $A$ is a continuous function with $A(t)=0$. Then the conclusion can be drawn using the properties of the convolution.
Proof of Lemma 5.4.
We want to show that the following expression goes to 0 as $\varepsilon$ goes to zero:
$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy=\int_{\mathbb{H}}\left|\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\big(\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\big)\,dv\right|dy.$
By the triangle inequality and Fubini’s theorem, this is bounded by
$\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\int_{\mathbb{H}}\left|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\right|\,dy\,dv=\phi_{\varepsilon}\ast_{\mathbb{R}}A(t),$
where
$A(v)=\int_{\mathbb{H}}|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|\,dy.$
By using the polar coordinates around $x$ together with the triangle inequality we have that
$A(v)\leq\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\,dy+\int_{\mathbb{H}}\frac{\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dy<\infty.$
Hence we observe that $v\mapsto|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|$ has a dominating function with which the dominated convergence theorem can be used to observe that $A(v)\to 0$ as $v\to t$. $A$ is continuous, and as it is continuous, we have that
$\lim_{\varepsilon\to 0}\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)=0,$
establishing the claim. ∎
6. Geometric data
In this section, we bound the Hilbert-Schmidt norm
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2},$
which is the geometric data appearing in the proof of Theorem1.5 in Section 8. $P_{t}$ and $P_{t,\tau}$ are the integral operators we extensively studied in the previous section, Section 5.
Proposition 6.1 (Geometric data).
Let $P_{t}$ and $P_{t,\tau}$ be integral operators on a compact connected hyperbolic surface $X$ with Laplace-Beltrami operator $\Delta_{X}$ such that $P_{t}$ has integral kernel
$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\quad K_{t}(x,y)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
and $P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)\,K_{t}^{\Gamma}$. Then
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D_{1}\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)$
where $\beta=\beta(\lambda_{1}(X))$, where $\lambda_{1}(X)$ is the spectral gap of $\Delta_{X}$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The constant appearing is
$C_{7}=\frac{7200}{\pi}.$
This is the most technical part of the whole paper. The core idea is first to write $K_{t}^{\Gamma}$ and $K_{t,\tau}^{\Gamma}$ using $K_{t}$ while simultaneously simplifying the expressions by unifying the sums over the Fuchsian group $\Gamma$ with the integrals over the fundamental domains of the surface $X$ to get integrals over the whole hyperbolic plane. This careful splitting allows us to separate a certain weight function $F_{t,t^{\prime},\rho}$ into a separate integral in which it is easier to bound. Section 7 is dedicated for the analysis of $F_{t,t^{\prime},\rho}$. Finally, we split the expression into two sides, one with the points with large injectivity radius and the other with the points with small injectivity radius. Both cases require careful analysis to acquire reasonable bounds, especially the side with the large injectivity radius for which we need the exponential mixing theorem, Theorem 1.2.
The proof relies on multiple auxiliary lemmas presented after the proof. This approach makes the proof clearer.
Proof of Proposition 6.1.
Writing the Hilbert-Schmidt norm as a double integral over the fundamental domain $D$ of $X$ yields the following expression:
$\int_{D}\int_{D}\left|\frac{1}{T}\int_{0}^{T}\int_{D}K_{t,\tau}^{\Gamma}a(z)K_{t}^{\Gamma}(z,y)dt\right|^{2}dydx.$
Using that $K^{\Gamma}_{t,\tau}=\cos(t\tau)K_{t}^{\Gamma}$ and by expanding the squares, we can write the previous expression as follows:
$\int_{D\times D}\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D\times D}K_{t}^{\Gamma}(x,z)a(z)K_{t}^{\Gamma}(z,y)\,K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dz^{\prime}dzdt^{\prime}dtdydx.$
Using Fubini’s theorem to reorder the integrals yields the following equals expression:
$\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D\times D\times D\times D}K_{t}^{\Gamma}(x,z)a(z)K_{t}^{\Gamma}(z,y)K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dxdydz^{\prime}dzdt^{\prime}dt.$
This expression is bounded by adding the absolute values around it. Then we pull the absolute values inside to get the following upper bound:
$\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}\int_{D}\int_{D}K_{t}^{\Gamma}(x,z)a(z)K^{\Gamma}_{t}(z,y)\,K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dydxdz^{\prime}dz\right|dt^{\prime}dt.$
This can be written in a more compact form when gathering the factors containing $x$ into one integral and the factors containing $y$ into another integral and noticing that these two integrals are the same thus resulting into an integral squared:
$\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{D}K_{t}^{\Gamma}(x,z)K_{t^{\prime}}^{\Gamma}(x,z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt.$
Lemma 6.2 allows us to write the previous expression using the hyperbolic plane kernels $K_{t}$ and $K_{t^{\prime}}$ instead of the surface kernels $K_{t}^{\Gamma}$ and $K_{t^{\prime}}^{\Gamma}$:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K_{t}(x,z)K_{t^{\prime}}(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt$
Coefficient $(8\,\pi^{2})^{-1}$ came from the definition of $K_{t}^{\Gamma}$:
$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y).$
Lemma 6.3 allows us to rewrite the integral inside the absolute value of the previous integral expression in a very convenient form, eliminating but one sum. Thus the previous expression equals:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{\mathbb{H}}a(z)a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K_{t}(x,z)K_{t^{\prime}}(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K_{t}(y,\gamma z)K_{t^{\prime}}(y,z^{\prime})dy\right)dz^{\prime}dz\right|dt^{\prime}dt$
which further equals
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{\mathbb{H}}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt,$ (6.1)
when $F_{t,t^{\prime},\rho}$ is defined to equal
$\int_{\mathbb{H}}K_{t}(x,w)K_{t^{\prime}}(x,w^{\prime})dx,$
where $w,w^{\prime}\in\mathbb{H}$ are any points seprated by distance $\rho$. In Section 7, where we study $F_{t,t^{\prime},\rho}$ carefully, in Lemma 7.1 we will see that the chosen points $w,w^{\prime}$ do not matter, only their distance.
The definition of $F_{t,t^{\prime},d(z,z^{\prime})}$ shows that $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, which follows from the definition of $K_{t}$ and $K_{t^{\prime}}$. Using this observation gives us the following equal expression for Expression (6.1):
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (6.2)
Focus on the absolute values in the previous expression. They are bounded from above by
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$ (6.3)
where the integration over $D$ is split into two: the points having injectivity radius bigger than $4T$, $D(4T)$, and the points having injecitivity radius smaller than $4T$, $D\setminus D(4T)$. The large injectivity radius case and the small injectivity radius case are bounded separately.
The small injectivity radius term is bounded in Lemma 6.4. Using its bound gives the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
In the previous, the second term is bounded by the third term as the second term has no sum and the third one does. Hence the previous expression is bounded by
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
+ $+\ 2\,\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$
where the second term is the small injectivity radius case, which is bounded in Lemma 6.5. Using this lemma gives the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ 8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$ (6.4)
The latter term is bounded by
$8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}11e^{\beta}\sinh(\rho)(1+\rho)e^{-\beta\rho}e^{2\beta T}F_{t,t^{\prime},\rho}^{2}d\rho,$
since
$11\,e^{\beta}(1+\rho)e^{-\beta\rho}e^{\beta\,2T}\geq 1.$
Hence Expression (6.4) is bounded by:
$11e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
We have now acquired good bounds for the absolute values appearing in Expression (6.1), and this expression is thus bounded by
$\frac{11}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
As $F_{t,t^{\prime},\rho}$ equals $F_{t^{\prime},t,\rho}$ and since we integrate from 0 to $T$ both $t$ and $t^{\prime}$, we can use symmetry to integrate the inner integral of $t^{\prime}$ from 0 to only $t$ twice. Hence we write the previous expression as
$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$ (6.5)
The innermost integral can be bounded using Proposition 6.6, which uses the bound acquired for $F_{t,t^{\prime},\rho}$ in Section 7. We get the following upper bound for the previous expression:
$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}dt^{\prime}dt,$
where
$C_{6}=480\pi.$
The remaining two integrals have
$\frac{2T\,C_{6}}{\beta^{3}}$
as an upper bound, which is seen after evaluating these integrals. Then the previous integral expression has the following upper bound:
$\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where
$\frac{22\cdot 2\cdot C_{6}\cdot e}{8\pi^{2}}\leq\frac{7200}{\pi}=C_{7},$
where we used that $e^{\beta}$ is less than $e$. This is also an upper bound for the Hilbert-Schmidt norm we set out to bound. It is of the wanted form which concludes the proof.
The remaining of this section contains the lemmas needed to prove Proposition 6.1 together with the contents of Section 7. The following lemma allows us to write the integration containing the kernels living on the hyperbolic surface using the kernels living on the hyperbolic plane.
Lemma 6.2.
Let
$K^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}K(x,\gamma y),\qquad L^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}L(x,\gamma y)$
be integral kernels on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, where $K$ and $L$ are integral kernels on $\mathbb{H}$ that depend only on distance $d(x,y)$. Then
$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma_{2}z^{\prime})dx,$
where $D$ is a fundamental domain of $X$.
The proof is on uniting the sum over the Fuchsian group and the integration over the fundamental domain into an integration over the hyperbolic plane.
Proof of Lemma 6.2.
We have that
$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(x,\gamma_{1}z)L(x,\gamma_{2}z^{\prime})\,dx$
by opening the kernels using their sum form. Doing a change of variables $\gamma_{2}\mapsto\gamma_{1}\gamma_{2}$ for the sum allows us to write the previous expression as
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(x,\gamma_{1}z)L(x,\gamma_{1}\gamma_{2}z^{\prime})dx.$
Using that $K(x,\gamma_{1}z)=K(\gamma_{1}^{-1}x,z)$ and that $L(x,\gamma_{1}\gamma_{2}z^{\prime})=L(\gamma_{1}^{-1}x,\gamma_{2}z^{\prime})$ allows us then to write the previous expression as:
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(\gamma_{1}^{-1}x,z)L(\gamma_{1}^{-1}x,\gamma_{2}z^{\prime})dx.$
Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover gives us the following equal expression:
$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})\,dx,$
where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. This concludes the proof.
The following lemma allows us to simplify the notation by removing one sum over the Fuchsian group.
Lemma 6.3.
Let $K$ and $L$ be integral kernels on compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$. Let $a$ be a function on $X$. Then:
$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}$
$=\qquad$ $\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime},$
where $D$ is a fundamental domain of $X$.
The proof is based on the same principle as in the proof of Lemma 6.2: we unite the sum over the Fuchsian group and the integral over the fundamental domain into an integration over the hyperbolic plane.
Proof of Lemma 6.3.
Focus on the following expression:
$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}.$
We expand the squares and reorder the integrals and the sums using Fubini’s theorem to get the following expression:
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}z^{\prime})dy\right)dz^{\prime}.$
Doing a change of variables $\gamma_{2}\mapsto\gamma_{2}^{-1}\gamma_{1}$ in the sum allows us to write the previous expression as
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}^{-1}\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
which further equals
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma_{2}y,\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
when using that $L(x,\gamma y)=L(\gamma^{-1}x,y)$. Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover is possible since $a^{*}$ can be seen as $\Gamma$-periodic function on $\mathbb{H}$. This gives the following equal expression:
$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy\right)dz^{\prime},$ (6.6)
where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. For the latter integral we have that
$\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy=\int_{\mathbb{H}}K(\gamma^{-1}y,z)L(y,z^{\prime})dy=\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy,$
where the first equality follows from the change of variables $y\mapsto\gamma^{-1}y$, and the second equality follows from the fact that $L(\gamma^{-1}x,y)=L(x,\gamma y)$. With this observation, we can rewrite Expression (6.6) as
$\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime}.$
The following lemma deals with the integration over the part of the fundamental domain with large injectivity radius, giving us the upper bound consisting of two terms: one is an integral with integrand with exponential decay and the other is an integration over the points with the small injectivity radius. The first term is desired and can be united with the bound acquired from the small injectivity radius case, but the second term needs to be further bounded. This is taken care of when we deal with all the parts dealing with the parts with small injectivity radius, which is done in Lemma 6.5.
Lemma 6.4.
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. $D(4T)$ denotes the points with injectivity radius bigger than $4T$. Let $a$ be a mean zero function on $X$. Then
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right|$
$=\ \|a\|_{2}^{2}\,\int_{0}^{2T}\,11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}\,(1+\rho)\,e^{-\beta\rho}\,d\rho$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|,$ (6.7)
where $\beta=\beta(\lambda_{1}(X))$, $\lambda_{1}(X)$ is the spectral gap of Laplace-Beltrami operator $\Delta_{X}$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The proof is based on modifying the expression so that the exponential mixing theorem, Theorem 1.2, can be applied.
Proof of Lemma 6.4.
As $z\in D(4T)$, we have that $d(z,\gamma z)$ is larger than $4T$ for any non-identity group element $\gamma$. We also have that $z^{\prime}\in B(z,2T)$. This means that $d(\gamma z,z^{\prime})$ is larger than $2T$ for any non-identity group element. As $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, we have that only the identity element can have a non-zero contribution. Expression (6.4) equals
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.8)
By adding and subtracting
$\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz$
in the absolute values of Expression (6.8) and then appyling the triangle inequality gives the following expression as an upper bound for Expression (6.8):
$\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.9)
Focus on the first term. We write $z^{\prime}$ using polar coordinates around $z$, so $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, and the first term equals
$\left|\int_{D}a(z)\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)\,a^{*}(\varphi_{\rho}(z,\theta))\,F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz\right|,$
where $a$ can be seen as a function on the unit tangent bundle by defining that $a(z,\theta)=a(z)$. The integrals of this expression can be reordered using Fubini’s theorem yielding the following expression:
$\left|\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\int_{D}\int_{\mathbb{S}^{1}}a(z)\,a^{*}(\varphi_{\rho}(z,\theta))\,d\theta\,dz\,d\rho\right|.$
By pulling the absolute values inside the first integral, we have the following upper bound, where the inner two integrals are written as an inner product:
$\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,\left|\left\langle a,a\circ\varphi_{\rho}\right\rangle_{L^{2}(D\times\mathbb{S}^{1})}\right|d\rho.$
We can apply the Exponential mixing theorem, Theorem 1.2, to the absolute value of the inner product which then yields the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}(1+\rho)e^{-\beta\rho}\,d\rho.$
Hence Expression (6.9) and hence the original expression we set out to bound are bound by
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$
The following lemma deals with the integration over the part of the fundamental domain with small injectivity radius giving an upper bound for it. This can then be united with the bound acquired for the large injectivity radius case in Lemma 6.4.
Lemma 6.5.
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. $D(4T)$ denotes the points with injectivity radius larger than $4T$. Then
$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
is bounded from above by
$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho$
The proof is based on careful use of Hölder’s inequality together with a careful management of the sum elements with non-zero contribtution.
Proof of Lemma 6.5.
Consider
$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
By pulling the absolute values inside this is bounded by
$\int_{D\setminus D(4T)}|a(z)|\int_{B(z,2T)}\left|a^{*}(z^{\prime})\right|F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz,$
which is further bounded by the following expression when using the supremum norm bounds for $|a|$ and $|a^{*}|$:
$\|a\|_{\infty}^{2}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$
In the previous expression, the sum elements with a possible non-zero contribution are those $\gamma\in\Gamma$ such that $d(\gamma z,z^{\prime})$ is smaller than $2T$ for some $z^{\prime}\in B(z,2T)$. This collection depends on $z$ and we denote it by $\Gamma(z)$. We have that
$\#\Gamma(z)\leq\frac{|B(z,4T)|}{\operatorname{InjRad}_{X}}\leq 2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}},$
which follows from that the minimum distance of $\gamma z$ and $\gamma^{\prime}z$, where $\gamma\neq\gamma^{\prime}$ is $\operatorname{InjRad}_{X}$, and that $\gamma z$ must lie in $B(z,4T)$ for it to be possible that $d(\gamma z,z^{\prime})$ is smaller than $2T$, since $z^{\prime}\in B(z,2T)$. Hence the previous integral expression has the following upper bound when using Hölder’s inequality and $\Gamma(z)$ instead of $\Gamma$:
$\|a\|_{\infty}^{2}\,\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}\right)^{\frac{1}{2}}\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}\left(\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right)^{2}\right)^{\frac{1}{2}}.$ (6.10)
Focus on the double integral in the second factor. Using Hölder’s inequality to sum and using the bound observed for $\Gamma(z)$ gives the following upper bound for it:
$2\pi\,\int_{D\setminus D(4T)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz.$
By using that $d(\gamma z,z^{\prime})$ equals $d(z,\gamma^{-1}z^{\prime})$ and interchanging the order of the sum and the integral using the linearity of integral, the previous expression equals
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,\gamma^{-1}z^{\prime})}^{2}dz^{\prime}dz,$
which further equals under a change of variables $z^{\prime}\mapsto\gamma z^{\prime}$ the following:
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
Due to the presence of $F_{t,t^{\prime},d(z,z^{\prime})}$, for non-zero contribution, we must have that $d(z,z^{\prime})$ is less than $2T$. This means that the previous expression equals
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)\cap B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz,$
which is then bounded by
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
As the $\gamma$-dependency is now eliminated, the sum contributes the number of elements it sums over as a coefficient, which leads to the following equal expression:
$4\pi^{2}\,\frac{e^{8T}}{\operatorname{InjRad}_{X}^{2}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
Hence we get the following bound for Expression (6.10), as now we have a bound for the second factor:
$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
By writing $z^{\prime}$ using polar coordinates around $z$, whence $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, we get the following equal expression:
$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz,$
which then equals
$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho,$
as there was no $\theta$ dependency nor $z$ dependency.
The following lemma bounds a complicated integral in which $F_{t,t^{\prime},\rho}$ appears. The bounds for $F_{t,t^{\prime},\rho}$ are established in Section 7.
Lemma 6.6.
Let $\beta$ be a positive constant that is smaller than $1$. Then:
$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}d\rho\leq\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
where
$F_{t,t^{\prime},\rho}\leq 32\pi\,\cdot\,\begin{cases}&\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad\text{ if }\rho<t-t^{\prime}\\ &\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},\qquad\text{ if }t-t^{\prime}<\rho<t\\ &\frac{1}{\sqrt{\sinh(\rho)}},\qquad\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
and where
$C_{6}=480\pi.$
The proof uses bounds for $F_{t,t^{\prime},\rho}$. Acquiring good enough bounds for the integral to converge requires very careful analysis which is done in Section 7. There we acquire three different bounds for $F_{t,t^{\prime},\rho}$ working in separate cases. Hence to bound the integral, we split it into three cases and use the bounds for $F_{t,t^{\prime},\rho}$. One of the cases is more technical as both the Beta function and the Gamma function appear in the analysis.
Proof.
We have that
$(1+\rho)\leq\frac{2}{\beta}\,e^{\frac{\beta}{2}-1}\,e^{-\frac{\beta}{2}\rho},$
which can be verified by studying the derivatives of the functions. Using this observation gives the following upper bound for the integral we are bounding:
$\frac{1}{\beta}\,\int_{0}^{2T}\sinh(\rho)\,e^{-\frac{\beta}{2}\rho}\,F_{t,t^{\prime},\rho}^{2}d\rho.$
By splitting the integral and using the bounds for $F_{t,t^{\prime},\rho}$, we have the following upper bound for the integral expression we set out to bound:
$\frac{32\pi}{\beta}\left(\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{\sqrt{\sinh(\rho)}\,e^{-\frac{\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\ +\ \int_{t-t^{\prime}}^{t}\frac{e^{-\frac{\beta}{2}\rho}}{(1-e^{t-t^{\prime}-\rho})^{\frac{1}{4}}}d\rho\ +\ \int_{t}^{t+t^{\prime}}e^{-\frac{\beta}{2}\rho}d\rho\right).$ (6.11)
Let us focus on the first term inside the parentheses. It is bounded by
$\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{e^{\frac{1-\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\leq\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\int_{0}^{t-t^{\prime}}\frac{1}{\sqrt{t-t^{\prime}-\rho}}d\rho,$
where the inequality followed from using the upper boundary of the integration domain. The previous then equals after evaluating the integral:
$2\,\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\sqrt{t-t^{\prime}}\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
The inequality followed from the observation that a positive $x$ divided by $\sinh(x)$ is bounded from above by $1$.
Next, let us bound the second term inside the parentheses of Expression (6.11). We begin by writing it in a different form using a change of variables $u=t^{\prime}+\rho-t$:
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{t^{\prime}}e^{-\frac{\beta}{2}u}\,\left(1-\frac{1}{\cosh(u)}\right)^{-\frac{1}{4}}\,du.$
We do another change of variables $z=e^{-u}$ to get
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,\frac{dz}{z},$
where especially
$1-\frac{1}{\cosh(u)}=\frac{(1-z)^{2}}{1+z^{2}}.$
The previous integral expression equals
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}-1}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,dz\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{1}z^{\frac{\beta}{2}-1}\,(1-z)^{-\frac{1}{2}}\,dz,$
where the inequality followed from that $(1+z^{2})^{\frac{1}{4}}$ is bounded by $2$ and the integration domain is enlargened. The previous expression then equals
$2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,B\left(\frac{\beta}{2},\frac{1}{2}\right),$ (6.12)
where $B$ is the Beta function which can be written using Gamma functions (the reference for the Beta and Gamma functions appearing here is [Olv10, Chp.6]):
$B\left(\frac{\beta}{2},\frac{1}{2}\right)=\frac{\Gamma\left(\frac{1}{2}\right)\,\Gamma\left(\frac{\beta}{2}\right)}{\Gamma\left(\frac{1}{2}+\frac{\beta}{2}\right)}.$
We have that $\Gamma\left(\frac{1}{2}\right)$ equals $\sqrt{\pi}$. We also have that
$\Gamma(x)=\frac{1}{x}\Gamma(x+1),$
and that
$\frac{\Gamma(x+1)}{\Gamma\left(x+\frac{1}{2}\right)}\leq(x+1)^{\frac{1}{2}},$
for positive $x$. Hence
$B\left(\frac{\beta}{2},\frac{1}{2}\right)\leq\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}.$
Hence Expression (6.12) is bounded by
$2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
Let us focus then on the third term inside the parentheses of Expression (6.11). The integral can be evaluated thus giving the following equal expression:
$\frac{2}{\beta}\left(e^{-\frac{\beta}{2}t}-e^{-\frac{\beta}{2}(t+t^{\prime})}\right),$
which has the following upper bound
$\frac{2\,e^{-\frac{\beta}{2}(t-t^{\prime})}}{\beta}.$
All three terms inside the parentheses of Expression (6.11) are bounded and they give the final upper bound for the original integral:
$\frac{32\,\pi}{\beta}\,e^{-\frac{\beta}{2}(t-t^{\prime})}\left(2\ +\ 2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\ +\ \frac{2}{\beta}\right)\leq\frac{480\pi}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
where we especially noted that $\beta<1$ for the inequality.
7. Weight function
Consider the following integral:
$F_{t,t^{\prime},\rho}=\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}},$
where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. The following lemma shows that the chosen points do not matter as long as the distance between the points is the same.
Lemma 7.1.
Let $z,z^{\prime},w,w^{\prime}\in\mathbb{H}$ such that $d(z,z^{\prime})$ equals $d(w,w^{\prime})$. Then
$\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}$
$=\ \int_{B(w,t)\cap B(w^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}}.$ (7.1)
The proof is based on using the properties of the isometries in the hyperbolic plane.
Proof of Lemma 7.1.
There exists an isometry $f$ such that $f(z)$ equals $w$ and that $f(z^{\prime})$ equals $w^{\prime}$. Using a change of variables $x\mapsto f^{-1}(x)$, we get the following equal expression for the left-hand side of Expression (7.1):
$\int_{f\left(B(z,t)\cap B(z^{\prime},t^{\prime})\right)}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(f^{-1}(x),z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(f^{-1}(x),z^{\prime}))}}$
$=\ \int_{f\left(B(z,t)\cap B(z^{\prime},t^{\prime})\right)}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}},$
where the equality followed from the fact that $d(f^{-1}(x),z)=d(x,f(z))$ which then equals $d(x,w)$, and similarly $d(f^{-1}(x),z^{\prime})=d(x,w^{\prime})$. As $f$ maps $z$ to $w$ and $z^{\prime}$ to $w^{\prime}$, we have that
$f\left(B(z,t)\cap B(z^{\prime},t^{\prime})\right)=B(w,t)\cap B(w^{\prime},t^{\prime}).$
Using the previous observation gives us that the previous integral expression equals
$\int_{B(w,t)\cap B(w^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}}.$
The main goal of this section is to give an upper bound for $F_{t,t^{\prime},\rho}$. Here, we additionally assume that $t>t^{\prime}$ for the case $t^{\prime}<t$ is symmetric. $F_{t,t^{\prime},\rho}$ is a crucial weight appearing in Section 6 when proving bounds for the geometric data appearing in the proof of Theorem 1.5 in Proposition 6.1. Bounding this weight is done in three lemmas, Lemmas 7.3, 7.4, and 7.5, of which settings are illustrated, respectively, in the following three figures:
Figure 2. The case of $t>t^{\prime}+\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is just the smaller ball $B(z^{\prime},t^{\prime})$ for it is enveloped by the bigger ball $B(z,t)$. The bound in this case is established using the polar coordinates around $z^{\prime}$.
Figure 3. The case of $\rho<t<t^{\prime}+\rho$, where $a+b=\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The middle point $z^{\prime}$ of the smaller ball is contained in the bigger ball. The integration domain is split into two parts, $D_{1}$ and $D_{2}$ by two geodesics of length $t^{\prime}$ going from $z^{\prime}$ to points $w$ and $w^{\prime}$ which are the points where the two balls intersect. The bound in this case is established separately for $D_{1}$ and $D_{2}$. For $D_{1}$, we use the polar coordinates around $z^{\prime}$ and for $D_{2}$ the polar coordinates around $z$.
Figure 4. The case of $t<\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The middle point $z^{\prime}$ of the smaller ball is not contained in the bigger ball. The bound in this case is established using the polar coordinates around $m$ which is the intersection of geodesics $z,z^{\prime}]$ and $w,w^{\prime}]$, where $w$ and $w^{\prime}$ are the intersections of the two balls.
From these three lemmas we get the following bound for $F_{t,t^{\prime},\rho}$.
Proposition 7.2.
$F_{t,t^{\prime},\rho}\leq C_{1}\,\cdot\,\begin{cases}\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad&\text{ if }\rho<t-t^{\prime}\\ \frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{\frac{1}{8}},\qquad&\text{ if }t-t^{\prime}<\rho<t\\ \frac{1}{\sqrt{\sinh(\rho)}},\qquad&\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
where
$C_{1}=\max\left(C_{2},C_{3},C_{4}\right)=\max(32,32\pi,99)=32\pi.$
The following lemma gives the bound for $F_{t,t^{\prime},\rho}$ in the first case, see Figure 2.
Lemma 7.3.
Assume that $t$ is larger than $t^{\prime}+\rho$. Then
$F_{t,t^{\prime},\rho}\leq\frac{C_{2}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
where
$C_{2}=32.$
The proof of lemma is based on using two different lower bounds for $\cosh(t)-\cosh(R)$, when working in the polar coordinates around $z^{\prime}$. Using these two bounds simultaneously allows us to deal with the problematic small angles, when the boundaries of two balls are close to each other.
Proof of Lemma 7.3.
We write the integral using the polar coordinates around $z^{\prime}$, taking $\theta=0$ in the direction $z,z^{\prime}](z^{\prime})$. Hence $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$. The situation is illustrated in Figure 2. Hence $F_{t,t^{\prime},\rho}$ is written as
$\int_{0}^{t^{\prime}}\int_{-\pi}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
$=\qquad$ $2\int_{0}^{t^{\prime}}\int_{0}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}},$ (7.2)
where half of the angles are dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ using the polar coordinates $r$ and $\theta$.
We want to find a lower bound for $\cosh(t)-\cosh(R)$. We remark that we have the following inequality since $t>t^{\prime}+\rho>R$:
$\cosh(t)-\cosh(R)\geq\cosh(t)-\cosh(t^{\prime}+\rho).$
This implies the following inequality by using the mean value theorem and the convexity of the hyperbolic sine for the positive input:
$\cosh(t)-\cosh(R)\geq\sinh(t^{\prime})\,\sinh(\rho)\,(t-t^{\prime}-\rho).$ (7.3)
We can also get another inequality. We write $\cosh(t)$ as follows:
$\cosh(t)=\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime}).$
Then, by using the hyperbolic law of cosines, we have that
$\cosh(R)=\cosh(\rho)\cosh(r)+\sinh(\rho)\sinh(r)\cos(\theta).$
Hence $\cosh(t)-\cosh(R)$ equals
$\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$
Since $r<t^{\prime}$ and $\rho<t-t^{\prime}$, we have the following inequality:
$\cosh(t)-\cosh(R(r,\theta))\geq(1-\cos(\theta))\,\sinh(t)\,\sinh(t-t^{\prime}).$ (7.4)
We write Expression (7.2) as
$2\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}d\theta dr,$
which has the following upper bound
$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(1-\cos(\theta))^{\frac{1}{4}}}d\theta dr,$
which was acquired by using the inequalities appearing in Expressions (7.3) and (7.4). The following equal expressions are acquired by evaluating the integrals:
$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,2^{\frac{5}{4}}\,K\left(\frac{1}{2}\right)dr$
$=\ \frac{2^{2+\frac{5}{4}}\,K(\frac{1}{2})\,\sqrt{\cosh(t^{\prime}-1)}}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
where $K$ is the complete elliptic integral of the first kind (see [Olv10, Chp.19]). The previous expression has a further upper bound, where we also bound the elliptic integral by $2$:
$\frac{2^{\frac{17}{4}}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\leq 32\,\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}.$
We proceed to a lemma bounding the second case, see Figure 3.
Lemma 7.4.
Assume that $\rho<t<\rho+t^{\prime}$. Then:
$F_{t,t^{\prime},\rho}\leq C_{3}\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},$
where
$C_{3}=32\pi.$
The idea of the proof is to split the integration domain into two separate domains. The bounding of the other domain resembles the approach of the previous lemma in that we establish two different lower bounds for $\cosh(t)-\cosh(R)$, which allow us to get a bound when used simultaneously. Bounding the other domain turns out to be simple. The following figure, Figure 5, further illustrates these separate domains:
((a)) The integration domain $D_{1}$ consists of those points that are outside the triangle formed in the ball $B(z^{\prime},t^{\prime})$. In this case, the polar coordinates around $z^{\prime}$ are used.
((b)) The integration domain $D_{2}$ consists of those points that are inside the triangle formed in the ball $B(z^{\prime},t^{\prime})$. In this case, polar coordinates around $z$ are used.
Figure 5. The case is split into two case which are bounded separately.
Proof of Lemma 7.4.
The integration domain of $F_{t,t^{\prime},\rho}$ can be split into a sum of integration over two disjoint domains, $D_{1}$ and $D_{2}$, see Figure 3. We bound these separately. $D_{1}$ is better illustrated in Figure 5(a) and $D_{2}$ in Figure 5(b).
Focus on $D_{1}$ first. We write it using the polar coordinates around $z^{\prime}$. Thus $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$, where $\theta=0$ corresponds to direction $z^{\prime},z](z^{\prime})$. In this way, the integral over $D_{1}$ equals
$\int_{-\Theta}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
= $=\ 2\int_{0}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
where half the angles were dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ in the polar coordinates $r$ and $\theta$. Notice that the maximum angle $\Theta$ is larger than $\pi/2$. We split the integral into two terms as follows:
$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
+ $+\ 2\int_{\frac{\pi}{2}}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$ (7.5)
By using the hyperbolic law of cosines we get that
$\cosh(t)=\cosh(\rho)\cosh(t^{\prime})-\sinh(\rho)\sinh(t^{\prime})\cos(\Theta),$ (7.6)
$\cosh(R)=\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$ (7.7)
Hence $\cosh(t)-\cosh(R)$ equals
$\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\cos(\theta)\sinh(r)-\cos(\Theta)\sinh(t^{\prime})).$ (7.8)
Hence we have the following inequalities:
$\cosh(t)-\cosh(R)\geq\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))$ (7.9)
and
$\cosh(t)-\cosh(R)\geq\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r)).$ (7.10)
Using the inequality of Expression (7.9), we get that the first term of Expression (7.5) is bounded by
$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
By evaluating the integral over $r$, we get the following equal expression:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)-\cos(\Theta)}}\,d\theta.$
This has the following upper bound, when we notice that both $\cos(\theta)$ and $\cos(\Theta)$ are positive. Additionally the ratio with the hyperbolic sine and cosine is bounded by $1$:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)}}\,d\theta.$
The following bound follows from the well known upper bound for the integral:
$\frac{12}{\sqrt{\sinh(\rho)}}.$
Let us focus on the second term of Expression (7.5). We write it as follows by interchanging the integrals using Fubini’s theorem:
$2\,\int_{0}^{t^{\prime}}\int_{\frac{\pi}{2}}^{\Theta}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}\,d\theta\,dr.$
We use the two inequalities appearing in Expressions (7.9) and (7.10) to get the following upper bound for the previous expression:
$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\cos(\theta)-\cos(\Theta))^{\frac{1}{4}}}\,d\theta\,dr.$
Using the mean value theorem, we get the following upper bound:
$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\Theta-\theta)^{\frac{1}{4}}}d\theta dr.$
Integrating the inner integral yields
$\frac{4}{3}\,\left(\Theta-\frac{\pi}{2}\right)^{\frac{3}{4}}.$
Hence we have the following bound for the previous integral expression:
$\frac{4\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}dr$
$=\ \frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}},$ (7.11)
where the equality followed from the evaluation of the remaining integral.
Focus momentarily on $\sin(\Theta)$. Expression (7.6) can be modified into the following form using the hyperbolic identities:
$\cos(\Theta)=1-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}.$
The subtracted division term must be larger than $1$ as the cosine is negative due to $\Theta$ being in $(\frac{\pi}{2},\pi)$. We have that
$\sin(\arccos(1-x))=\sqrt{x}\sqrt{2-x}\geq\sqrt{2-x},$
when $x>1$. Hence:
$\sin(\Theta)\geq\sqrt{2-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}}=\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\sinh(t^{\prime})\sinh(\rho)}},$
where the equality followed using hyperbolic identities. The previous implies the following inequality:
$\sin(\Theta)\geq\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\cosh(t^{\prime}+\rho)}}=\sqrt{1-\frac{\cosh(t)}{\cosh(t^{\prime}+\rho)}}.$
This then gives us the following inequality:
$\sin(\Theta)\geq\sqrt{1-\frac{1}{\cosh(t^{\prime}+\rho-t)}}.$ (7.12)
Applying the inequality of Expression (7.12) in Expression (7.11) gives the following upper bound:
$\frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
This has the following further upper bound:
$\frac{16\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
The integration over $D_{1}$ has the following upper bound, when both of the terms of Expression (7.5) are added together:
$\frac{1}{\sqrt{\sinh(\rho)}}\left(12+16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\right)$
$\leq\ 16\pi\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
Next, we will focus on bounding the integral over $D_{2}$. We write it using the polar coordinates around $z$. Thus $x=\pi_{1}(\phi_{R}(z,\theta))$, where $\theta=0$ corresponds to direction $z,z^{\prime}](z)$. Then the integral equals:
$\int_{-\Psi}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta$
= $=\ 2\int_{0}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta,$
where the equality followed from using the symmetry. $r=r(R,\theta)$ is $d(x,z^{\prime})$ in the polar coordinates $R$ and $\theta$. $R(\theta)$ denotes the minimum distance that needs to be propagated into direction $\theta$ so that we are in $D_{2}$. Notice also that $\Psi$ is less than $\pi/2$. By using the hyperbolic law of cosines we have that
$\cosh(t^{\prime})=\cosh(\rho)\cosh(t)-\sinh(\rho)\sinh(t)\cos(\Psi),$
$\cosh(r)=\cosh(\rho)\cosh(R)-\sinh(\rho)\sinh(R)\cos(\theta).$
As $\cosh(r)$ grows as $R$ grows, we have the following lower bound for $\cosh(t^{\prime})-\cosh(r)$:
$\sinh(\rho)\sinh(t)(\cos(\theta)-\cos(\Psi)).$
Hence the integral expression has the following upper bound:
$\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\frac{1}{\sqrt{\sinh(t)}}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}}dRd\theta.$
This has an upper bound when we integrate from 0 instead of $R(\theta)$; the following expression is acquired after we evaluate the integral:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{\sqrt{\cosh(t)-1}}{\sqrt{\sinh(t)}}d\theta.$
This is then bounded by
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}d\theta,$
which has the following upper bound using the mean value theorem:
$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\sin(\theta)}}.$
Using the well-known bound for sine, we get the following upper bound:
$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\frac{2}{\pi}\theta}}d\theta.$
This is actually the Beta integral (see [Olv10, Chp. 5]), so we get the following equal expression:
$\frac{4\sqrt{\pi}\pi}{\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)}}.$
The integration over both $D_{1}$ and $D_{2}$ is done. Uniting the bounds of these two cases gives us the final bound for $F_{t,t^{\prime},\rho}$:
$\frac{1}{\sqrt{\sinh(\rho)}}\left(16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\ +\ \frac{4\pi\sqrt{\pi}}{\sqrt{2}}\right)$
$\leq\qquad$ $\frac{32\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
The following lemma takes care of the last case, see Figure 4.
Lemma 7.5.
Assume that $t>\rho$. Then:
$F_{t,t^{\prime},\rho}\leq C_{4}\,\frac{1}{\sqrt{\sinh(\rho)}},$ (7.13)
where
$C_{4}=99.$
The proof is again based on splitting the integration domain into two. This time the splitting is done by $w,w^{\prime}]$ which splits the integration into two sides, one closer to $z$ and one closer to $z^{\prime}$. The technique used to bound both sides is the same, based on establising carefully bounds using hyperbolic geometry.
Proof of Lemma 7.5.
The integration domain of $F_{t,t^{\prime},\rho}$ is the intersection of $B(z,t)$ and $B(z^{\prime},t^{\prime})$. Since $t<\rho+t^{\prime}$, the intersection is not a ball. Let $w$ and $w^{\prime}$ be the intersections of the boundaries of the balls, and let $m$ be the intersection of $z,z^{\prime}]$ with $w,w^{\prime}]$. We parametrize the domain using the polar coordinates around $m$ with $\theta=0$ corresponding to direction $z,z^{\prime}](m)$. Hence $x=\pi_{1}(\phi_{r}(m,\theta))$. To easen the notation, we call $|[z,m]|=a$, $|[m,z^{\prime}]|=b$, and $|[m,w]|=h$. We will present $d(z,x)$ and $d(z^{\prime},x)$ using polar coordinates, denoted by $R=R(r,\theta)$ and $P=P(r,\theta)$, respectfully. For illustration, see Figure 4. Hence $F_{t,t^{\prime},\rho}$ is in polar coordinates
$\int_{0}^{2\pi}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}},$
where $r(\theta)$ is the maximum distance we can go from $m$ to direction $\theta$ to still stay in the integration domain. We can split the integration into two parts:
$\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}}$
$+\ \int_{\frac{\pi}{2}}^{\frac{3\pi}{2}}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}}.$ (7.14)
The first term integrates over the side closer to $z^{\prime}$ while the second term integrates the side closer to $z$, these sides are separated by $w,w^{\prime}]$.
Let us focus on the integrand of the first term. Using the hyperbolic law of cosines, we get the following equalities:
$\cosh(t)=\cosh(a)\cosh(r(\theta))+\sinh(a)\sinh(r(\theta))\cos(\theta),$
$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
where we note that $\cos(\theta)$ is positive. We also notice that by using the hyperbolic law of cosines we get that
$\cosh(t)=\cosh(a)\cosh(h),$
$\cosh(t^{\prime})=\cosh(b)\cosh(h),$
which implies that
$\cosh(t^{\prime})=\frac{\cosh(b)}{\cosh(a)}\cosh(t).$
Inserting the earlier equality of $\cosh(t)$ into the previous yields the following:
$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))+\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))\cos(\theta).$
We want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\sinh(a)\left(\sinh(r(\theta))-\sinh(r)\right),$
which then has a lower bound:
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right).$
Next, we want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right),$
and then has a lower bound
$\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
Thus $(\cosh(t)-\cosh(R))(\cosh(t^{\prime})-\cosh(P))$ is bounded from below by
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
This equals
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(b)\sinh(a)\sinh(r(\theta))+\cosh(a)\sinh(b)\sinh(r)\right),$
and is then bounded from below by
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(b)\sinh(a)+\cosh(a)\sinh(b)\right).$
As $a+b=\rho$, the previous equals
$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
Now, let us focus on the integrand of the second term of Expression (7.14). Using the hyperbolic law of cosines, we get the following equalities:
$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))-\sinh(b)\sinh(r(\theta))\cos(\theta),$
$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
where $\cos(\theta)$ is negative. We also notice that, using the hyperbolic law of cosines,
$\cosh(t)=\cosh(a)\cosh(h),$
$\cosh(t^{\prime})=\cosh(b)\cosh(h),$
which implies that
$\cosh(t)=\frac{\cosh(a)}{\cosh(b)}\cosh(t^{\prime}).$
The previous gives the following equality after inserting an earlier equality of $\cosh(t^{\prime})$ in:
$\cosh(t)=\cosh(a)\cosh(r(\theta))-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))\cos(\theta).$
We want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)-\cos(\theta)\sinh(b)\left(\sinh(r(\theta))+\sinh(r)\right).$
This has a lower bound
$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right).$
Next, we want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))-\sinh(a)\sinh(r)\right),$
and then has a lower bound
$(-1)\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
Thus $(\cosh(t)-\cosh(R(r,\theta)))(\cosh(t^{\prime})-\cosh(P(r,\theta)))$ is bounded from below by
$(-1)\,\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)\,\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
This equals
$(-1)\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(a)\sinh(b)\sinh(r(\theta))+\cosh(b)\sinh(a)\sinh(r)\right),$
and is then bounded from below by
$(-1)\,\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(a)\sinh(b)+\cosh(b)\sinh(a)\right).$
As $a+b=\rho$, the previous equals
$(-1)\,\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
These observations for the integrands of the terms of Expression (7.14) imply that we get the following upper bound for $F_{t,t^{\prime},\rho}$:
$\frac{1}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\int_{0}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\,d\theta$ (7.15)
Focus on the inner integral. Assume first that $r(\theta)<2$. The integral is bounded by the following by using the mean value theorem:
$\int_{0}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)}.$
This is bounded by $2\sqrt{2}$. Then, assume that $r(\theta)>2$. We split the integration as follows:
$\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr.$
$a$ is specified later. We have the following upper bound for the previous:
$\frac{1}{\sqrt{\sinh(a)}}\,\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr.$
The first term was bounded using the boundary of the integration domain and the second term was bounded using the mean value theorem. By evaluating the integrals, we acquire the following equal expression:
$\frac{2\,\sqrt{\cosh(r(\theta))-\cosh(a)}}{\sqrt{\sinh(a)}}\ +\ 2\,\sqrt{r(\theta)-a}\leq\frac{2\,\sqrt{r(\theta)-a}\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta))}}\ +\ 2\,\sqrt{r(\theta)-a}.$
The inequality followed from using the mean value theorem. We choose $a$ to equal $r(\theta)-1$. Then the previous equals
$\frac{2\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta)-1)}}\ +\ 2\leq\frac{2\,\sqrt{\sinh(2)}}{\sqrt{\sinh(1)}}\ +\ 2,$
where for the inequality we used that $r(\theta)>2$. The previous is bounded by $9$.
The previous observations mean that Expression (7.15) has the following upper bound:
$\frac{9}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\,d\theta.$
The integral is bounded from above by $11$, so we can conclude that
$F_{t,t\,\rho}\leq\frac{99}{\sqrt{\sinh(\rho)}}.$
8. Proof of main theorems
In this section we give proofs to the main theorems of this paper, Theorems 1.4, 1.5, and 1.7. We first prove the quantitative main theorem, Theorem 1.5. Most of the paper is dedicated to this end. Proving it is highly technical and thus for the clarity of the proof we have presented auxiliary results in separate sections, Sections 4, 5, 6, and 7.
After proving Theorem 1.5, we prove the limit form main theorem, Theorem 1.4. It follows straightforwardly from Theorems 1.5 and 8.1.
Proving Theorem 1.7 relies on using Theorems 1.5 and 8.2. The proof is very similar to that of Theorem 1.4 but due to the probabilistic setting imposing lighter conditions on the sequence, we need Theorem 8.2 to record some key results in random surface theory.
We proceed to prove Theorem 1.5. We restate the theorem for the convenience of the reader.
Theorem.
Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
$0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
$0<\delta<\tfrac{2}{9}\min(I),\qquad T\delta=\tfrac{\pi}{2},$
and fix $\tau\in\mathbb{R}$. Then
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The coefficient appearing is
$C_{8}=50000\pi.$
The proof presented here is clear as the technical details are dealt in different sections, see Sections 6, 5, 6, and 7. The core idea of the proof is to split the problem into a geometric side and a spectral side using a well-chosen propagator. The geometric side and the spectral side can then be dealt separately.
Proof of Theorem 1.5.
Define functions
$h_{t}(x)=x^{-1}\sin(tx),\qquad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x),$
and operators
$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right),\qquad P_{t,\tau}=h_{t,\tau}\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right).$
Then due to way the operators are defined, we have that
$P_{t}\psi_{j}=h_{t}(\rho_{j})\psi_{j},\qquad P_{t,\tau}\psi_{j}=h_{t,\tau}(\rho_{j})\psi_{j}.$
Using the previous, we can write that
$|\langle\psi_{j},a\psi_{k}\rangle|=\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|^{-1}\left|\left\langle\psi_{j},\,\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\,\psi_{k}\right\rangle\right|.$
The left factor containing spectral data is bounded using Proposition 4.1,which can be used since $I$ is compact in $(0,\infty)$, $T\delta=\pi/2$, and $\delta$ is smaller than $(2/9)\cdot\min(I)$:
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})h_{t}(\rho_{k})\,dt\right|^{-1}\leq 8\pi\rho_{j}\,\rho_{k}\leq 8\pi\max(I)^{2}.$
Thus,
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}|\langle\psi_{j},a\psi_{k}\rangle|^{2}$ $\leq 64\pi^{2}\,\max(I)^{4}\,\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\psi_{k}\right\rangle\right|^{2}$
$\leq 64\pi^{2}\,\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2},$
where the Hilbert-Schmidt norm bound is acquired when we consider all the basis elements. The Hilbert-Schmidt norm contains the geometric data of the problem.
Next we want to apply Proposition 6.1 to bound the Hilbert-Schmidt norm. This requires understanding the integral kernels of $P_{t}$ and $P_{t,\tau}$ for which a convenient form is given in Proposition 5.1. We apply Proposition 6.1 to acquire the following upper bound for the previous expression:
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where we opted to use $|X\setminus X(4T)|$ to denote the volume of the points of $X$ with injectivity radius smaller that $4T$. Also,
$C_{7}=\frac{7200}{\pi}.$
Hence we can conclude that
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where
$C_{7}\cdot 64\pi^{2}\leq 50000\pi=C_{8}.$
We now turn to Theorem 1.4, the limit form main theorem. We also restate the theorem for convenience.
Theorem.
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
$0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
Assume that the sequence has the following properties:
  • (BSC)
    Benjamini–Schramm convergence: for every $R>0$,
    $\lim_{n\to\infty}\frac{\big|\{x\in X_{n}:\operatorname{InjRad}_{X_{n}}(x)<R\}\big|}{|X_{n}|}=0,$
  • (EXP)
    Expander property: the sequence admits a uniform lower bound on the spectral gap,
  • (UND)
    Uniform discreteness: the sequence admits a uniform lower bound on the injectivity radius.
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator. Then the following properties hold:
  • (i)
    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
  • (ii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon,$
  • (iii)
    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
    $\limsup_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\right\rangle\right|^{2}<\epsilon.$
Properties (ii) and (iii) remain to be proven; Property (i) was established by Le Masson and Sahlsten (see Theorem 1.1 [LS17]). We remark that the proof presented here for properties (ii) and (iii) could be adapted with slight additions to also include property (i), thus giving a new proof for the large-scale quantum ergodicity theorem of Le Masson and Sahlsten.
Passing to the limit and thus acquiring a small upper bound $\epsilon$ is based on carefully using Properties (BSC), (EXP), and (UND) together with the large-scale Weyl law which is used to asymptotically relate the volume of the surface with the number of eigenvalues in an energy window, when we proceed in the sequence of surfaces. Before the proof, we state the large-scale Weyl law (see [LS17, Sec.9]).
Theorem 8.1 (Large-scale Weyl law).
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces satisfying property (BSC) stated in Theorem 1.4. $X_{n}$ has Laplace-Beltrami operator $\Delta_{X_{n}}$ with eigenvalues
$0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
Then
$\frac{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}{|X_{n}|}=\frac{\#\{j\ :\ \lambda_{j}^{(n)}\in J\}}{|X_{n}|}\xrightarrow{n\to\infty}C,$
where $C$ is a constant depending on the sequence of surfaces and the interval $I$.
We have included using both $\lambda_{j}^{(n)}$ and $\rho_{j}^{(n)}$ since Le Masson and Sahlsten proved their theorem using $\lambda_{j}^{(n)}$ approach. However, the theorem works just as well for $\rho_{j}^{(n)}$ as was seen in the statement of the previous theorem.
We proceed to prove Theorem 1.4.
Proof of Theorem 1.4.
We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)>0$ such that, for all $\tau\in\mathbb{R}$,
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\sum_{j:\rho_{j}^{(n)}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle\right|^{2}<\epsilon,$ (8.1)
where
$N(X_{n},I)=\#\{j\ :\ \rho_{j}^{(n)}\in I\}.$
Let us define a following function:
$\bar{a}_{n}=a_{n}-\frac{1}{|X_{n}|}\int_{X_{n}}a_{n}\,dx.$
$\bar{a}_{n}$ is a mean 0 normalization of $a_{n}$. Due to orthogonality of $\psi_{j}^{(n)}$ and $\psi_{k}^{(n)}$, we have that
$\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle=\langle\psi_{j}^{(n)},\bar{a}_{n}\psi_{k}^{(n)}\rangle.$
Hence we have the following equal expression for the left-hand side of Expression (8.1):
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\sum_{j:\rho_{j}^{(n)}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(n)},\bar{a}_{n}\psi_{k}^{(n)}\rangle\right|^{2}$
We can apply Theorem 1.5 to the double sum given that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(n)})^{3}}\,\left(\|\bar{a}_{n}\|_{2}^{2}\,+\,8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{n}\setminus X_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right),$ (8.2)
where $|X_{n}\setminus X_{n}(4T)|$ is the volume of the points of $X_{n}$ with injectivity radius being smaller than $4T$, $T\,\delta=\pi/2$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The constant appearing is
$C_{8}=50000\pi.$
Then we apply the properties of the sequence to tidy the previous expression. First we note that
$\|\bar{a}_{n}\|_{2}\leq\sqrt{|X_{n}|}\,\|\bar{a}_{n}\|_{\infty}.$
And then, since $(a_{n})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
$\|\bar{a}_{n}\|_{\infty}\leq 2\,a_{max}.$
Property (EXP) means that there is a universal lower bound for the spectral gaps which then implies that $\beta(\lambda_{1}^{(n)})$ has a universal lower bound. Denote this by $\beta_{min}$. Finally, Property (UND) means that there is a universal lower bound for the injectivity radii. Denote this by $\operatorname{InjRad}_{min}$. Using these observations gives the following upper bound for Expression (8.2):
$\limsup_{n\to\infty}\ 2\cdot C_{8}\,\frac{\max(I)^{4}\,a_{max}^{2}}{N(X_{n},I)\,T\,\beta_{min}^{3}}\left(|X_{n}|+\frac{8\pi^{2}\,|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{\operatorname{InjRad}_{min}}\right)$
$=\ \frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\left(\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right)+\frac{8\pi^{2}\,e^{6T}}{\operatorname{InjRad}_{min}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}\setminus X_{n}(4T)|}{N(X_{n},I)}\right)\right),$
where the equality followed from splitting the limit supremum into two terms.
Consider the second limit supremum inside the parentheses of the previous expression. Due to the large-scale Weyl law, Theorem 8.1, we have that $N(X_{n},I)$ grows asymptotically like the volume of surfaces $|X_{n}|$. This means that asymptotically the ratio inside the limit supremum closes to 0 as we proceed in the sequence since Property (BSC) forces the ratio of the volume of the points with small injectivity radius and the volume of the surface to go to zero. Hence the previous expression equals
$\frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right).$
Using the large-scale Weyl law again to the remaining limit supremum, we see that the ratio approaches some constant $C_{9}$. Hence the previous expression equals
$\frac{2C_{8}\,C_{9}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}$
$=\ \frac{4\,C_{8}\,C_{9}\,\delta\,\max(I)^{4}}{\beta_{min}^{3}\,\pi},$
where for the equality we used that $T\,\delta=\pi/2$. If we had chosen originally that
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{8}\,\max(I)^{4}}{\beta_{min}^{3}\,\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
In the end we prove the probabilistic theorem, Theorem 1.7. We restate the theorem for convenience.
Theorem.
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus. Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
$0=\lambda_{0}^{(g)}<\lambda_{1}^{(g)}\leq\dots\leq\lambda_{j}^{(g)}\leq\dots\xrightarrow{j\to\infty}\infty$
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator. Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
Properties (ii) and (iii) remain to be proven; Property (i) was already proven by Le Masson and Sahlsten [LS24]. The proof is very close to that of Therem 1.4. Now we do not assume Properties (BSC), (EXP), and (UND), so these need to be replaced. This is done by the following theorem acquired from the theory of the Weil-Petersson random surfaces.
Theorem 8.2.
Let $X$ be a $\mathbb{P}_{g}$-random surface, where $\mathbb{P}_{g}$ is the Weil-Petersson probability distribution on the moduli space $\mathcal{M}_{g}$ with genus $g$. It has eigenvalues
$0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
together with associated values
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}.$
Then with probability 1o(1)1-o(1) $X$ has the following properties:
  • (a)
    $\operatorname{InjRad}_{X}\geq g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}},$
  • (b)
    $\frac{|\{x\in X\ :\ \operatorname{InjRad}_{X}(x)<\frac{1}{6}\,\log(g)\}|}{|X|}=\mathcal{O}(g^{-\frac{1}{4}}),$
  • (c)
    $\lambda_{1}\geq\frac{1}{4}-\eta,$
where $\eta\in(0,\frac{1}{4})$. Additionally,
  • (d)
    $\{j\ :\ \rho_{j}\in I\}=\#\{j\ :\ \lambda_{j}\in J\}=|X|\cdot\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right),$
where $I\subset(0,\infty)$ is a compact interval and
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
The theorem is acquired from two articles: [Mon22, HMT25]. Properties (a), (b), and (d) together with the rates is acquired from [Mon22], and Property (c) together with rate is acquired from [HMT25]. When comparing the proof of Theorem 1.7 with the proof of Theorem 1.4, Property (a) replaces Property (UND), Property (b) replaces Property (BSC), and Property (c) replaces Property (EXP). Additionally, Property (d) is a suitably modified version of the large-scale Weyl law of Le Masson and Sahlsten working in this probabilistic setting.
Proof of Theorem 1.7.
We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)$ such that for every $\tau\in\mathbb{R}$, we have that
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle\right|^{2}<\epsilon,$ (8.3)
where
$N(X_{g},I)=\#\{j\ :\ \rho_{j}^{(g)}\in I\}.$
Let us define a following function:
$\bar{a}_{g}=a_{g}-\frac{1}{|X_{g}|}\int_{X_{g}}a_{g}\,dx.$
$\bar{a}_{g}$ is a mean 0 normalization of $a_{g}$. Due to orthogonality of $\psi_{j}^{(g)}$ and $\psi_{k}^{(g)}$, we have that
$\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle=\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle.$
Hence the left-hand side of Expression (8.3) equals
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle\right|^{2}.$
We can apply Theorem 1.5 to the double sum givem that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(g)})^{3}}\,\left(\|\bar{a}_{g}\|_{2}^{2}+8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right),$
where $|X_{g}\setminus X_{g}(4T)|$ is the volume of the points of $X_{g}$ with injectivity radius less than $4T$, $T\,\delta=\pi/2$, and where
$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
The constant appearing is
$C_{8}=50000\pi.$
First notice that
$\|a_{g}\|_{2}\leq\sqrt{|X_{g}|}\,\|a_{g}\|_{\infty}.$
Since the sequence $(a_{g})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
$\|\bar{a}_{g}\|\leq 2\,a_{max}.$
Hence the previous limit supremum expression is bounded from above by
$2\cdot C_{8}\,\frac{a_{max}^{2}}{T}\,\limsup_{g\to\infty}\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right).$ (8.4)
Using Theorem 8.2, we have that
$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
$\leq\ \frac{1}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(\frac{3}{16}-\eta\right)^{3}}\,\left(|X_{g}|+8\pi^{2}\,|X_{g}|\,\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
with propability $1-\mathcal{O}(g^{-c(\eta)})$, where $\eta\in(0,3/16)$, $c(\eta)$ is a positive constant depending on $\eta$, and where
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
The previous expression has a tidier upper bound:
$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+8\pi^{2}\,e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right).$
This means that with high propability Expression (8.4) equals
$2\,C_{8}\,C_{10}\,\frac{a_{max}^{2}}{T}=\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\,\delta$
where equality followed from using that $T\,\delta=\pi/2$. $C_{10}$ is a constant upper bound coming from
$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}.$
If we had chosen originally that
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
Acknowledgements.
The author would like to thank Prof. Mostafa Sabri, Prof. Tuomas Sahlsten, and Prof. Farrell Brumley for discussions that led to this article. The author also thanks Prof. Tuomas Sahlsten, Prof. Henrik Ueberschär, Dr. Félix Lequen, Dr. Søren Mikkelsen, and Dr. Joe Thomas for their guidance while writing this article.
Our use of the exponential is inspired by the upcoming article [HLM+] of the author with Félix Lequen, Søren Mikkelsen, Tuomas Sahlsten and Henrik Ueberschär on quantum ergodicity for eigenfunctions for Schrödinger operators on hyperbolic surfaces. We thank the other authors of that paper.
The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University.

### v2

@@
 Abstract.
-We study compact hyperbolic surfaces and multiplicative observables, establishing a large-scale analogue of Zelditch’s quantum mixing theorem with hypotheses that hold for both arithmetic and Weil-Petersson random surfaces of large genus.
-This complements the large-scale quantum ergodicity theorems of Le Masson and Sahlsten, which themselves serve as large-scale analogues of the quantum ergodicity theorem of Shnirelman, Zelditch, and de Vèrdiere, thereby providing a more complete picture of the asymptotic behaviour of observables in the large-scale limit.
-Our approach does not rely on ball averaging operators or Nevo’s ergodic theorem.
+We study compact hyperbolic surfaces and multiplication observables, establishing a large-scale analogue of Zelditch’s quantum mixing theorem with hypotheses that hold for both arithmetic and Weil–Petersson random surfaces of large genus.
+This complements the large-scale quantum ergodicity theorems of Le Masson and Sahlsten, which themselves are large-scale analogues of the quantum ergodicity theorem of Shnirelman, Zelditch, and Colin de Verdière, thereby providing a more complete picture of the asymptotic behavior of observables in the large-scale limit.
+Our approach does not rely on the ball averaging operator or Nevo’s ergodic theorem.
 Instead, we introduce a new method based on the hyperbolic wave equation and the quantitative exponential mixing of the geodesic flow established by Ratner and Matheus.
-The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University
+The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research Council of Finland project: Quantum chaos on large and many body systems, grant numbers 347365, 353738, and Aalto University.
 1. Introduction
@@
 Observables correspond to measurable quantities of the system.
-Understanding the interplay between the energy operator, its eigenfunctions, and the observables leads to a deeper understanding of the dynamics of the system.
-A central topic in quantum chaos is the study of localization and delocalization phenomena for the eigenfunctions of the energy operator.
-In this paper, the Hamiltonian system is assumed to be the geodesic flow $\phi_{t}$ of a compact connected Riemannian manifold $(M,g)$ and the quantum Hamiltonian is taken to be $\sqrt{-\Delta}$, where $\Delta$ is the Laplace-Beltrami operator on the manifold.
-For more background and a better overview, we refer to the following surveys: [Mar06, Zel05, Sar11, Zel19, Ana22].
-Consider a compact Riemannian manifold $(M,g)$ and the quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$.
-Let $A$ be a zeroth-order pseudodifferential operator representing an observable.
-We are interested in the diagonal elements $\langle\psi_{j},A\psi_{j}\rangle$, which correspond to the expected value of $A$ in the state $\psi_{j}$, and in the off-diagonal elements $\langle\psi_{j},A\psi_{k}\rangle$, which represent transition amplitudes from $\psi_{j}$ to $\psi_{k}$ induced by the observation of $A$.
-Quantum ergodicity theorem (Shnirelman [Shn74], Zelditch [Zel87], and Colin de Verdière [Col85]) is a cornerstone result of quantum chaos and serves as a quantum analogue of the classical ergodic theorem.
-For an ergodic compact Riemannian manifold, it states that almost all subsequences of eigenfunctions corresponding to increasing eigenvalues satisfy
+Understanding the interplay between the energy operator, its eigendata (eigenvalues and eigenfunctions), and the observables leads to a deeper understanding of the dynamics of the system.
+A central topic in quantum chaos is the study of localization and delocalization phenomena of the eigenfunctions of the energy operator.
+Here, the Hamiltonian system is assumed to be the geodesic flow $\varphi_{t}$ of a compact connected Riemannian manifold $(M,g)$ and the quantum Hamiltonian is taken to be $\sqrt{-\Delta}$, where $\Delta$ is the Laplace–Beltrami operator of the manifold.
+For more background and a better overview, we refer to the following surveys: [ZEL05, SAR11, ZEL19, ANA02, MAR02].
+Consider a compact Riemannian manifold $(M,g)$ and the quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace–Beltrami operator $\Delta$.
+Let $A$ be a zeroth-order pseudo-differential operator representing an observable.
+We are interested in the diagonal average $\langle\psi_{j},A\psi_{j}\rangle$, which corresponds to the expectation value of $A$ in the state $\psi_{j}$, and in the off-diagonal average $\langle\psi_{j},A\psi_{k}\rangle$, which represents the transition amplitude from $\psi_{j}$ to $\psi_{k}$.
+Quantum ergodicity theorem (Shnirelman [ŠNI74], Zelditch [ZEL87], and Colin de Verdière [COL85]) is a cornerstone result of quantum chaos and serves as a quantum analogue of the classical ergodic theorem.
+For an ergodic compact Riemannian manifold, it states that almost all subsequences of eigenfunctions corresponding to increasing eigenvalues satisfy that
 $\langle\psi_{j_{l}},A\psi_{j_{l}}\rangle\xrightarrow{l\to\infty}\omega(A),$
-where $\omega(A)$ denotes the average of the principal symbol $\sigma_{A}$ of $A$ (the classical counterpart of observable $A$) over the unit cotangent bundle of $M$.
+where $\omega(A)$ denotes the average of the principal symbol $\sigma_{A}$ of $A$ (the classical counterpart of observable $A$) over the unit co-tangent bundle of $M$.
 Roughly speaking, the theorem asserts that the probability distributions induced by the eigenfunctions converge to the uniform distribution in the weak sense of testing the distributions against observables.
-Shnirelman’s theorem concerns the diagonal elements of observables.
-Zelditch [Zel90, Zel96b] extended it to the off-diagonal elements, given that the compact Riemannian surface is now weak mixing, proving that for any $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
+The quantum ergodicity theorem focuses on the diagonal averages of the observables.
+Zelditch [ZEL90, ZEL96b] extended it to the off-diagonal averages, given that the geodesic flow on the compact Riemannian surface is now weak mixing, proving that for any $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$
 $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j:\mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}-\tau|<\delta(\epsilon)\end{subarray}}\left|\left\langle\psi_{j},A\psi_{k}\right\rangle\right|^{2}<\epsilon.$
-Informally, this means that the summed transition amplitudes between states separated approximately by energy $\tau$ decay to zero for almost all subsequences of eigenfunctions as the eigenvalue increases.
-A unified formulation encompassing both of these results can be given following the works of multiple authors: [Shn74, Col85, Zel87, Laz93, Zel90, GL93, Zel96a, Zel96b, ZZ96, Sun97].
+This means that almost all subsequences of eigenfunctions with increasing eigenvalues, $(\psi_{j_{l}})$, satisfy that the transition amplitudes into a small energy window of separation $\tau$ from the eigenvalue of $\psi_{j_{l}}$ can be made arbitrarily small as $l\to\infty$ by choosing a small enough energy window.
+A unified formulation encompassing both of these results can be given following the works of multiple authors: [ŠNI74, COL85, ZEL87, ZEL90, GL93, LAZ93, ZEL96a, ZEL96b, ZZ96, SUN97].
 The version below adapts Theorems 4.1 and 4.4 from [ZEL05].
 Theorem 1.1 (Semi-classical quantum mixing, [SUN97, ZEL05]).
-Let $(M,g)$ be a compact Riemannian manifold (possibly with a boundary) with quantum Hamiltonian $\sqrt{-\Delta}$ with eigenvalues $\{\mu_{j}=\sqrt{\lambda_{j}}\}$ and eigenvalues $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace-Beltrami operator $\Delta$.
-Let $A$ be a zeroth-order pseudodifferential operator.
+Let $(M,g)$ be a compact Riemannian manifold (possibly with a boundary) with quantum Hamiltonian $\sqrt{-\Delta}$ together with eigenvalues $\{\mu_{j}=~\sqrt{\lambda_{j}}\}$ and eigenfunctions $\{\psi_{j}\}$, where $\{\lambda_{j}\}$ and $\{\psi_{j}\}$ are the eigenvalues and eigenfunctions of the Laplace–Beltrami operator $\Delta$.
+Let $A$ be a zeroth-order pseudo-differential operator.
 Consider the following properties:
@@
   • (II)
-    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that
-    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j\ :\ \mu_{j}<\mu\}}\sum_{\lambda_{j}<\lambda}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}|<\delta(\epsilon)\end{subarray}}\left|\langle\psi_{j},A\psi_{k}\rangle\right|^{2}<\epsilon,$
+    For every $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that
+    $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j:\mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}|<\delta(\epsilon)\end{subarray}}\left|\langle\psi_{j},A\psi_{k}\rangle\right|^{2}<\epsilon,$
   • (III)
-    For every $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
+    For every $\epsilon>0$, there exists $\delta(\epsilon)>0$ such that for all $\tau\in\mathbb{R}$,
     $\limsup_{\mu\to\infty}\ \frac{1}{\#\{j:\mu_{j}<\mu\}}\sum_{j:\mu_{j}<\mu}\sum_{\begin{subarray}{c}k:\mu_{k}<\mu,\\ k\neq j,\\ |\mu_{j}-\mu_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j},A\psi_{k}\rangle|^{2}<\epsilon.$
 Then the geodesic flow $\varphi_{t}$ is ergodic on $(S^{*}M,d\mu)$ if and only if Properties (I) and (II) hold.
-It is weakly mixing on $(S^{*}M,d\mu)$ if and only if properties (i), (ii), and (iii) hold.
-Here $S^{*}M$ denotes the unit tangent bundle of $M$ and $d\mu$ is the Liouville measure on $S^{*}M$.
+It is weakly mixing on $(S^{*}M,d\mu)$ if and only if Properties (I)–(III) hold.
+Here $S^{*}M$ denotes the unit co-tangent bundle of $M$ and $d\mu$ is the Liouville measure on $S^{*}M$.
 This theorem exhibits a delicate connection between the classical and quantum settings.
@@
 We therefore say that a quantum system satisfying Properties (I) and (II) is quantum ergodic, while one satisfying Properties (I)(III) is quantum mixing.
-Our focus will be on the compact connected hyperbolic surfaces, which are inherently chaotic due to their (constant) negative curvature.
+Our focus will be on compact connected hyperbolic surfaces, which are inherently chaotic due to their (constant) negative curvature.
 In this context, the following theorem of Ratner [RAT87], with quantitative bounds due to Matheus [MAT13], will be central.
@@
 Theorem 1.2 (Position space exponential mixing).
-Let $X$ be a compact connected hyperbolic surface with geodesic flow $\phi_{t}$ and with $\lambda_{1}(X)$ being the smallest eigenvalue of the Laplace-Beltrami operator $\Delta$.
+Let $X$ be a compact connected hyperbolic surface with geodesic flow $\varphi_{t}$ and with $\lambda_{1}(X)$ being the second smallest eigenvalue of the Laplace–Beltrami operator $\Delta$.
 Let $f,g\in L^{2}(X)$.
-Then for all $t\geq 0$,
-$\bigg|\big\langle f\circ\varphi_{t},g\big\rangle_{L^{2}(T^{1}X)}-\int_{T^{1}X}f\,dx\int_{T^{1}X}g\,dx\bigg|\leq 11\,e^{\beta(\lambda_{1}(X))}\,(1+t)\,e^{-\beta(\lambda_{1}(X))t}\,\|f\|_{2}\,\|g\|_{2},$
+Then for all $t\geq 0$
+$\bigg|\big\langle f\circ\varphi_{t},g\big\rangle_{L^{2}(S^{*}X)}-\int_{S^{*}X}f\,dx\int_{S^{*}X}g\,dx\bigg|\leq 11\,e^{\beta(\lambda_{1}(X))}\,(1+t)\,e^{-\beta(\lambda_{1}(X))t}\,\|f\|_{2}\,\|g\|_{2},$
 where
 $\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4},\end{cases}$
-and where $f(x)=f(x,\theta)$ and $g(x)=g(x,\theta)$ are viewed as functions on $T^{1}X$, the unit tangent bundle.
-The more general phase space version (see [Rat87, Mat13]) implies that the geodesic flow on a compact connected hyperbolic surface is weak mixing with exponential decay which further implies ergodicity.
-By Theorem 1.1, a compact connected hyperbolic surface satisfies properties (i)(iii), and hence is both quantum ergodic and quantum weak mixing.
-Remark 1.3.
-From Theorem 1.1, we see that neither the $2$-torus nor the $2$-sphere is quantum ergodic or quantum weakly mixing.
-However, if we restrict the observables to be taken from the collection of the physical-space observables, i.e.
-the multiplicative operators corresponding to functions, then a $2$-torus with irrational aspect ratio is quantum ergodic (though still not quantum weak mixing).
-The $2$-sphere remains neither quantum ergodic nor quantum weak mixing.
-These examples can be analyzed explicitly due to the availability of closed-form expressions for their eigenfunctions and eigenvalues.
-An illustarting example showing this for a discrete circle with $N$ elements can be acquired using observable $a(x)=e^{\frac{2\pi i\,x}{N}}$; this was pointed out by Mostafa Sabri.
-This idea should translate to the $2$-dimensional surfaces as well.
-We see that the choice of observables matters.
-These phenomena on Riemannian surfaces naturally motivate the search for analogous results in other contexts, such as discrete settings.
-A pseudo-analogue of the above corollary to Theorem 1.1 was obtained for certain finite regular graphs by Anantharaman and Le Masson [AL15].
-Their result, however, is not a strict analogue, since the high-energy limit is not meaningful for finite graphs, which possess only finitely many eigenvalues.
-Instead, they considered a sequence $(G_{n})$ of graphs with increasing size $n$ and a fixed energy window $I$, as opposed to a fixed structure with increasing energy as in Shnirelman’s theorem.
-A special case of their result states that, under suitable assumptions, if $(a_{n})$ is a sequence of diagonal observables acting on functions on the vertices and if $(\Delta_{n})$ is a sequence of graph Laplacians each with eigenvalues $\{\lambda_{j}^{(n)}\}$ and eigenfunctions $\{\psi_{j}^{(n)}\}$, then
-$\frac{1}{\#\{j:\lambda_{j}^{(n)}\in I\}}\sum_{\lambda_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{\operatorname{Tr}(a_{n})}{n}\right|^{2}\xrightarrow{n\to\infty}0.$
-As in Shnirelman’s theorem, the averages of the expectation values converge to the uniform average.
-We refer to such results, in which the structure grows while energy remains bounded, as large-scale results, in contrast to the original results in which the energy frows.
+and where $f(x)=f(x,\theta)$ and $g(x)=g(x,\theta)$ are viewed as functions on $S^{*}X$, the unit co-tangent bundle.
+The more general phase space version (see [RAT87, MAT13]) implies that the geodesic flow on a compact connected hyperbolic surface is weak mixing, which further implies ergodicity.
+By Theorem 1.1, a compact connected hyperbolic surface satisfies Properties (I)(III), and hence is both quantum ergodic and quantum mixing.
+From Theorem 1.1, it follows that neither the $2$-torus nor the $2$-sphere is quantum ergodic or quantum mixing, since their geodesic flows are neither ergodic nor weakly mixing.
+However, if one restricts the class of observables to physical-space observables, that is, multiplication operators corresponding to functions, then all flat $2$-tori satisfy Property (I); see [MR12] for details.
+This illustrates that the choice of observables plays a crucial role.
+On the other hand, it is an interesting question whether Properties (II) and (III) hold for flat $2$-tori.
+For further discussion, see Appendix A.
+These phenomena on Riemannian surfaces naturally motivate the search for analogous results in other settings, for example, the discrete setting.
+Anantharaman and Le Masson [AL15] established that, for certain finite graphs, the analogue of Property (I) holds.
+Their result, however, is not a strict analogue, since finite graphs have only finitely many eigenvalues.
+Instead, they considered a sequence $(G_{n})$ of graphs with increasing size $n$ and a fixed energy window $I$; as the graph grows, the number of eigenvalues in the energy window increases.
+This differs from the quantum ergodicity theorem, in which one considers a fixed structure at increasing energy levels, so that all eigenvalues below a given energy and the corresponding eigenfunctions are considered.
+A special case of the theorem of Anantharaman and Le Masson [AL15] states that if $(a_{n})$ is a uniformly bounded sequence of diagonal matrices acting on functions on the vertices of the graph, and if $(\Delta_{n})$ is a sequence of graph Laplacians, each with eigenvalues $\{\lambda_{j}^{(n)}\}$ and eigenfunctions $\{\psi_{j}^{(n)}\}$, then under certain conditions (analogous to Properties (BSC) and (EXP) introduced in Subsection 1.2) on the structure of the graphs, we have
+$\lim_{n\to\infty}\frac{1}{\#\{j:\lambda_{j}^{(n)}\in I\}}\sum_{j:\lambda_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{\operatorname{Tr}(a_{n})}{n}\right|^{2}=0.$
+As in the quantum ergodicity theorem, the averages of the expectation values converge to the average of the observable with respect to the uniform distribution.
+We refer to such results in which the structure grows while the energy remains bounded as large-scale results, in contrast to the original results in which the energy grows.
 We refer to the latter results as large-energy results.
-The large-scale quantum ergodicity of graphs and related theorems is an active field with a vigorous interplay with the continuous settings.
-For more realted results in the discrete graph setting, see [DP12, BL13, EKYY13, TVW13, AL15, Gei15, BLL16, Ana17, AS17, BHKY17, BKY17, AS19, BS19, BHY19, McK22, MS23].
-An especially impressive result establsihed in [BHKY17, BHY19, BKY17] shows a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] for random regular graphs; the randomization effectively suppresses pathological cases helping to establish stronger results compared to the deterministic setting.
-Since the large-scale graph quantum ergodicity theorem is not a true analogue of Shnirelman’s theorem, it is natural to seek a corresponding surface result highlighting the distinctions from the original large-energy setting.
-Such an analogue for hyperbolic surfaces was established by Le Masson and Sahlsten [LS17] (see Subsection 1.2 for details).
-Their result resembles closely that of Anantharaman and Le Masson, with minor modifications reflecting the transition from discrete graphs to continuous surfaces.
-Despite the distinct nature of the two limits, there exists a heuristic correspondence between the large-energy and large-scale settings.
-In the large-energy regime, as the energy grows, the eigenfunctions exhibit shorter wavelengths, resolving increasingly fine structures.
-In the large-scale regime, as the structure grows, finer details become accessible within a fixed energy window.
-Hence, these two asymptotic regimes may be regarded as analogous in spirit.
-To our knowledge, the only existing large-scale off-diagonal results establishing analogues of properties (i), (ii), and (iii) of Theorem 1.1 (see Subsection 1.2) are those of Cipolloni, Erdős, and Schröder [CES21] for Wigner matrices.
-Their setting resembles that of Anantharaman and Le Masson, but with graphs replaced by random matrices and observables by general matrices.
-Their result is remarkably strong as they require no averaging: they prove that
-$\left|\left\langle\psi_{j}^{(n)},A_{n}\psi_{k}^{(n)}\right\rangle-\delta_{jk}\tfrac{\operatorname{Tr}(A_{n})}{n}\right|$
-decays to zero at the optimal and uniform rate of $n^{-1/2}$ for all $j,k$, thereby establishing a large-scale analogue of the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94], and in fact extending it to the off-diagonal elements.
-This result is relevant as the random matrix theory is closely related to the quantum ergodicity on graphs and surfaces, especially if Berry’s conjecture (see [Ber77]) holds.
+The study of the large-scale quantum ergodicity of graphs and related problems is an active field with a vigorous interplay with the continuous setting.
+For more related results in the discrete graph setting, see [DP12, BL13, EKY+13, TVW13, AL15, GEI15, BLL16, ANA17, AS17, BHK+17a, BKY17b, AS19, BS19, BHY19, MCK22, MS23].
+An impressive result established in [BHK+17a, BHY19, BKY17b] shows an average-free large-scale analogue of the large-scale quantum ergodicity theorem of Anantharaman and Le Masson [AL15] for random regular graphs.
+This resembles the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] in the random setting.
+However, the observable in [BHK+17a, BHY19, BKY17b] must be chosen independently of the graph.
+Nevertheless, the randomization effectively suppresses pathological cases, helping to establish stronger results when compared to the deterministic setting.
+Since the large-scale quantum ergodicity theorem for graphs by Anantharaman and Le Masson [AL15] differs from the large-energy quantum ergodicity theorem of Shnirelman [ŠNI74], Zelditch [ZEL87], and Colin de Verdière [COL85], it is natural to seek a corresponding surface result for the large-scale graph theorem that would highlight distinctions when compared to the original large-energy quantum ergodicity theorem.
+Such an analogue for hyperbolic surfaces was established by Le Masson and Sahlsten [LS17] (see Subsection 1.2 for more details).
+Their result closely resembles that of Anantharaman and Le Masson in that they prove a similar analogue of Property (I) of the original quantum ergodicity theorem; minor differences reflect the transition from discrete graphs to continuous surfaces.
+It is tempting to try to equate the large-energy setting with the large-scale setting.
+Heuristically, one might say that in the large-energy regime we use higher and higher energies to resolve finer details, whereas in the large-scale setting we zoom into the structure to capture finer features.
+However, the large-scale setting appears to be much more sensitive to the underlying structure of the manifold or graph.
+The large-energy quantum ergodicity theorem says that all Riemannian manifolds for which the geodesic flow is ergodic are quantum ergodic and that those for which the geodesic flow is weak mixing are quantum mixing; this is open for the variable-curvature large-scale setting.
+Moreover, in the case of graphs, the semiclassical measure is not the uniform measure, so it is not expected to be uniform either on manifolds of variable negative curvature, in contrast to the large-energy regime, which does not see this subtlety.
+To our knowledge, the only previously existing large-scale off-diagonal results establishing analogues of Properties (I)(III) of Theorem 1.1 (see Subsection 1.2 for the properties) are those of Cipolloni, Erdős, and Schröder [CES21] for Wigner matrices.
+Their setting resembles that of Anantharaman and Le Masson [AL15], where graphs are replaced by random matrices.
+Their result is remarkably strong, as they require no averaging: they prove that
+$\lim_{n\to\infty}\left|\left\langle\psi_{j}^{(n)},A_{n}\psi_{k}^{(n)}\right\rangle-\delta_{jk}\tfrac{\operatorname{Tr}(A_{n})}{n}\right|=0;$
+the decay to 0 has the optimal and uniform rate of $n^{-1/2}$ for all $j,k$, where $n\times n$ matrices are considered with $n$ growing.
+Thereby Cipolloni, Erdős, and Schröder [CES21] establish a result that resembles the quantum unique ergodicity conjecture of Rudnick and Sarnak [RS94] in the large-scale setting; moreover, they extend it to the off-diagonal case as well.
+We remark that, as in [BHK+17a, BHY19, BKY17b], the observable is chosen independently of the matrix.
+This result is relevant, as random matrix theory is closely related to the study of quantum ergodicity and related problems on graphs and surfaces, especially if Berry’s conjecture holds (see [BER77]).
 For more related results for random matrices, see [ESY09a, ESY09b, BY17].
-A new upcoming result by Bordenave, Letroit, and Sabri [BLS] establishes a quantum ergodicy and quantum mixing results on growing Schreier graphs where even the quantum ergodicy was previously unknown.
-To our knowledge, this is first mixing result in the graph setting.
-This is especially interesting as graphs are intuitively closer to the original setting on manifolds when compared to random matrices.
+Studying the off-diagonal setting is timely: a new result by Bordenave, Letrouit, and Sabri [BLS26] establishes quantum ergodicity and quantum mixing results on growing Schreier graphs, where even quantum ergodicity was previously unknown.
+To our knowledge, this is the first mixing result in the graph setting.
+This is especially interesting, as the graphs are intuitively closer to the original setting on manifolds when compared to random matrices.
 1.2. Main theorems
-We prove a large-scale analogue of the corollary of Theorem 1.1 for physical-space observables, that is, for functions viewed as multiplicative operators, thereby complementing the result of Le Masson and Sahlsten [LS17].
-We also prove a probabilistic analogue of this, complementing the probabilistic quantum ergodicity result of Le Masson and Sahlsten [LS24].
-The small difference is that we use the quantum Hamiltonian $\sqrt{-\Delta-1/4}$ instead of $\sqrt{-\Delta}$ but causes only a spectral shift.
-We already refer to Section 3 for more technical overview of the proof ideas as well as the novel ideas of the proofs used.
-The proof follows the ideas of Le Masson and Sahlsten in that we establish a quantitative version which yields limiting forms of the theorem.
-We again divide the argument into a spectral and a geometric component.
-However, we employ a different propagator, namely the hyperbolic wave propagator
-$P_{t}=h_{t}\!\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right),\qquad h_{t}(x)=x^{-1}\sin(tx).$
-The intuition behind $P_{t}$ is that $u_{tt}=-\Delta\,u$, the wave equation in the hyperbolic plane, has solution $u=2^{-1/2}\,P_{t}\,u_{0}$ with initial data
-$u(z,0)=0,\qquad u_{t}(z,0)=2\pi f,$
-for some function $f$. For details, see [LP82]. A convenient form of the integral kernel of $P_{t}$ is established in Section 5.
-Unlike the ball-averaging operator used by Le Masson and Sahlsten, which has no immediate connection to the underlying classical dynamics, the hyperbolic wave propagator is obtained from the quantization of the geodesic flow, thus linking the quantum and classical settings in a natural manner.
-This choice of propagator renders the spectral side particularly clean, though it makes the geometric side more involved compared to the approach of Le Masson and Sahlsten.
-To access the off-diagonal elements $\langle\psi_{j},a\psi_{k}\rangle$, where the energy levels of $\psi_{j}$ and $\psi_{k}$ may differ substantially, we also introduce a modified propagator: $P_{t,\tau}=\cos(t\tau)\,P_{t}$.
-Using $P_{t}$ and $P_{t,\tau}$ simultaneously allows us to capture the behaviour of the off-diagonal elements.
-This, however, complicates both the spectral and geometric analyses compared to the quantum ergodicity case, which concerns only the diagonal elements.
-In particular, the Heisenberg evolution of the observable $a$, $P_{t}^{*}aP_{t}$, appearing in the proof of quantum ergodicity by Le Masson and Sahlsten, is replaced here by $P_{t,\tau}^{*}aP_{t}$.
-The physical interpretation of this operator is not immediately transparent, but it can be verified that
-$P_{t,\tau}^{*}aP_{t}=P_{t,\tau/2}^{*}aP_{t,\tau/2}+P_{-t,\tau/2}^{*}aP_{-t,\tau/2},$
-meaning it equals the sum of two Heisenberg evolutions of $a$ with $P_{t,\tau/2}$, one forward and one backward in time.
-Finally, we replace rather abstract Nevo’s theorem with the exponential mixing theorem (see Theorem 1.2), which elucidates the connection between classical and quantum dynamics.
-The idea of using the exponential mixing originates in the upcoming article of the author with Lequen, Mikkelsen, Sahlsten, and Ueberschär [HLM+].
-The use of exponential mixing is both simple and transparent: the exponential decay of correlations of $a$ and $a\circ\varphi_{t}$, even for small $t$, is essential in obtaining our result.
-Small propagation times suffice because we work with observables in a macroscopic sense, unlike the large-energy quantum ergodicity and quantum weak mixing setting, which employs microlocal analysis and requires long propagation times for decorrelation to occur at the microlocal scale.
-The more technical approach taken in the proofs is presented in Section 3.
-We now state the main theorems.
+We prove a large-scale analogue of the corollary of Theorem 1.1 for physical-space observables, that is, for functions viewed as multiplication operators, thereby complementing the result of Le Masson and Sahlsten [LS17].
+We also prove a probabilistic analogue of this using the Weil–Petersson random model, complementing the probabilistic quantum ergodicity result of Le Masson and Sahlsten [LS24].
+A small difference compared to the large-energy setting is that we use the quantum Hamiltonian $\sqrt{-\Delta-\tfrac{1}{4}}$ instead of $\sqrt{-\Delta}$; however, this causes only a spectral shift.
+We proceed to introduce the main theorems now.
 Theorem 1.3 (Large-scale quantum ergodicity and quantum mixing).
-Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
+Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace–Beltrami operators $(-\Delta_{X_{n}})$. The Laplacian $\Delta_{X_{n}}$ has eigenvalues
 $0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
-together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
-$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
+together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values of the eigenvalues as follows:
+$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\tfrac{1}{4}}.$
 Assume that the sequence has the following properties:
@@
     Uniform discreteness: the sequence admits a uniform lower bound on the injectivity radius.
-Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator.
+Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplication operator.
 Then the following properties hold:
   • (i)
-    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
+    $\lim_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a_{n}\,dx\right|^{2}=0,$
   • (ii)
@@
 Properties (i)(iii) are the large-scale analogues of Properties (I)(III) in Theorem 1.1.
-Accordingly, we refer to systems satisfying Properties (i) and (ii) as large-scale quantum ergodic, and those satisfying (i)(iii) as large-scale quantum mixing.
-As in Theorem 1.1, the choice of observables plays a role, recall the example on the unit circle where quantum mixing fails discussed in the earlier remark, Remark 1.3.
-We note that this idea can be modified to prove that the $2$-torii of increasing area are not quantum mixing for physical-space observables, though there are sequences of $2$-torii that are, interestingly, quantum ergodic using the explicit formula of the eigenfunctions.
-We refer to the earlier illustrating example discussed This observation also illustrates that large-scale quantum ergodicity and large-scale quantum weak mixing are properties specific to certain sequences of hyperbolic surfaces rather than to general sequences of Riemannian manifolds.
-Le Masson and Sahlsten [LS17] have already proven that property (i) holds in this setting.
-Hence to establish Theorem 1.4, it remains to prove properties (ii) and (iii).
-To prove the previous theorem, we need the following quantitative version of it.
+Accordingly, we refer to systems satisfying Properties (i) and (ii) as large-scale quantum ergodic, and those satisfying Properties (i)(iii) as large-scale quantum mixing.
+We remark that Property (i) was proven by Le Masson and Sahlsten in [LS17].
+Property (i) is also the surface analogue of the large-scale quantum ergodicity theorem of Anantharaman and Le Masson [AL15] discussed in the previous subsection.
+In Appendix A, we construct an example of a sequence of growing flat $2$-tori for which Properties (ii) and (iii) fail.
+This demonstrates that large-scale quantum mixing is not a generic property for sequences of Riemannian surfaces.
+It remains an interesting question whether Property (i) holds for sequences of growing tori; establishing this would complement the large-energy result, analogous to Property (I) of Theorem 1.1, proved by Marklof and Rudnick [MR12].
+For the standard eigenfunctions
+$\psi_{(m,n)}(x,y)=e^{-2\pi i\left(\frac{mx}{L_{x}}+\frac{ny}{L_{y}}\right)},\qquad(m,n)\in\mathbb{Z}^{2},$
+the claim is straightforward. However, establishing the result for all eigenbases seems to be subtle.
+An example of a sequence of surfaces that is both quantum ergodic and quantum mixing in the above sense is a sequence of compact arithmetic surfaces.
+This follows from the result of Katz, Schnaps, and Vishne [KSV07], who showed that the injectivity radius grows (ensuring Properties (BSC) and (UND)), and from the result of Selberg [SEL65], who established a uniform lower bound of $\tfrac{3}{16}$ on the spectral gap (ensuring Property (EXP)).
+We remark that the result of Kim, Sarnak, and Ramakrishnan [KRS03] is the best deterministic bound currently available on the spectral gap.
+Moreover, the class of admissible sequences is quite broad, which is seen when using the probabilistic approach via the Weil–Petersson random model; for background on the Weil–Petersson random model, see [MIR13, WRI20].
+We state the probabilistic analogue of Theorem 1.3, thus showing that a sequence of growing Weil–Petersson random surfaces is quantum ergodic and quantum mixing with high probability.
+Theorem 1.4 (Probabilistic version).
+Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil–Petersson probability distribution on the moduli space $\mathcal{M}_{g}$, and where $g$ is the genus.
+Let $\Delta_{g}$ be the Laplace–Beltrami operator of $X_{g}$ with eigenvalues
+$0=\lambda_{0}^{(g)}<\lambda_{1}^{(g)}\leq\dots\leq\lambda_{j}^{(g)}\leq\dots\xrightarrow{j\to\infty}\infty$
+together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values of the eigenvalues as follows:
+$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\tfrac{1}{4}}.$
+Let $I\subset(0,\infty)$ be a compact interval, and let $(a_{g})$ be a sequence of uniformly bounded functions on $X_{g}$, each acting as a multiplication operator.
+Then Properties (i)(iii) of Theorem 1.3 hold with high probability as the genus $g$ grows to infinity.
+We remark that the use of randomization is highly active in the field of quantum chaos.
+Randomization effectively suppresses pathological cases, helping to establish strong results; the analogous deterministic settings are far from acquiring the corresponding results.
+We especially want to highlight the following important results: the Benjamini–Schramm convergence and a lower bound on the injectivity radius in the Weil–Petersson random model were established in [MIR13, MON21].
+For recent results on the spectral gap, see [AM24, LW24, MAG24, HMT25a, HMN25b, WX25].
+Another random model, the random covering model, has been used with great success to obtain results on the spectral gap; see [MN20, MNP22, HM23].
+As mentioned, Le Masson and Sahlsten have already proven that Property (i) holds in the setting of Theorem 1.3 (see [LS17]).
+They also proved that Property (i) holds in the setting of Theorem 1.4 (see [LS24]).
+Hence, to establish Theorems 1.3 and 1.4, it remains to prove Properties (ii) and (iii) in both settings.
+The approaches used to prove Theorems 1.3 and 1.4 are very similar: they rely on the quantitative version of the theorems, namely Theorem 1.5.
+Using the quantitative theorem together with Properties (BSC), (EXP), (UND), and the large-scale Weyl law (see Theorem 8.1) yields a relatively straightforward proof of Theorem 1.3.
+When proving the probabilistic theorem, Theorem 1.4, we do not have Properties (BSC), (EXP), and (UND) available, nor do we have the large-scale Weyl law.
+These tools can be replaced with those recorded in Theorem 8.2.
+We remark that Theorem 8.2 is compiled from [LS24, Thm.
+1.7, 1.9]; the references given for this theorem differ from those of Le Masson and Sahlsten [LS24], reflecting that the understanding of the spectral gap is evolving rapidly.
+Next, we state the quantitative main theorem and comment on the approach taken in its proof.
 Theorem 1.5 (Quantitative version).
-Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
+Let $X$ be a compact connected hyperbolic surface with Laplace–Beltrami operator $\Delta_{X}$ that has eigenvalues
 $0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
-together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
-$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
-Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
+together with a corresponding orthonormal eigenbasis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values of the eigenvalues as follows:
+$\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}.$
+Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplication operator. Let
 $0<\delta<\tfrac{2}{9}\min(I),\qquad T\delta=\tfrac{\pi}{2},$
 and fix $\tau\in\mathbb{R}$. Then
-$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
-where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
+$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\,\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
+where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, defined by
 $\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
-The coefficient appearing is
-$C_{8}=50000\pi.$
+The coefficient appearing is $C_{8}=5120000\pi^{2}$.
 Remark 1.6.
-We observe from the statement of Theorem 1.5 that in the proof of Theorem 1.4 in Section 8, we could relax many of the conditions.
-E.g.
-we could let the spectral gap and the injectivity radius to shrink slowly as we proceed in the sequence.
-Thereby Theorem 1.4 could be extended to encompass a larger family of sequences.
-An example of a sequence of surfaces that is both quantum ergodic and quantum mixing in the above sense is a sequence of compact arithmetic surfaces.
-This follows from Katz et al.
-[KSV07], who showed that the injectivity radius grows (ensuring properties (BSC) and (UND)), and from Kim et al.
-[KSR03], who established a uniform lower bound on the spectral gap (ensuring property (EXP)).
-The result of Kim et al.
-is the best deterministic bound currently available.
-The first, weaker, bound is Selberg’s $3/16$ bound, see [Sel65].
-Moreover, the class of admissible sequences is quite broad which is seen when using probabilistic approach using the Weil-Petersson random model; for background, see [Mir13, Wri20].
-By Le Masson and Sahlsten [LS24], for a sequence $(X_{g})$ of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is the Weil-Petersson probability measure on the moduli space $\mathcal{M}_{g}$ of hyperbolic surfaces of genus $g$, the condition (i) of Theorem 1.4 holds with high probability as $g$ grows to infinity.
-We state now a probabilistic analogue of Theorem 1.4, thus showing that a growing Weil-Petersson random surface sequence is quantum ergodic and quantum mixing with high probability and further complementing the results of Le Masson and Sahlsten [LS24].
-Theorem 1.7 (Probabilistic version).
-Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus.
-Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
-$0=\lambda_{0}^{(g)}<\lambda_{1}^{(g)}\leq\dots\leq\lambda_{j}^{(g)}\leq\dots\xrightarrow{j\to\infty}\infty$
-together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
-$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
-Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator.
-Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
-As mentioned, Property (i) was proved by Le Masson and Sahlsten [LS24].
-Hence it remains to prove Properties (ii) and (iii) where the appraoch of the proof follows closely that of Theorem 1.4.
-However, as the assumptions are lighter in the probabilistic setting we do not have Properties (BSC), (EXP), and (UND) available, nor do we have the large-scale Weyl law, Theorem 8.1.
-We need analogues of these in the proof of Theorem 1.7.
-Theorem 8.2 consisting of recent results in the random surface theory offers us the required tools.
-The use of Weil-Petersson random surfaces in quantum ergodicity is currently very active.
-Randomization effectively suppresses pathological cases helping to establish stong results that the analogous deterministic settings are far away from acquiring.
-For results in Benjamini-Schramm convergence and injectivity radius in Weyl-Peterson random model, see [Mir13, Mon21].
-For recent results on the spectral gap, see [AM24, LW24, Mag24, HMT25, HMN25, WX25].
-There are other random models and we especially refer to late groundbreaking results using the random covering model, see [MN20, MNP22, HM23].
-See an interesting application of Property (EXP) to the Bose-Einstein condensation of random hyperbolic surfaces [LS22].
+We observe from the statement of Theorem 1.5 that we could relax many of the conditions appearing in the statement of Theorem 1.3.
+For example, we could let the spectral gap and the injectivity radius slowly shrink as we proceed along the sequence.
+Thereby, Theorem 1.3 could be extended to encompass a larger family of sequences of surfaces.
+The proof of the quantitative theorem also follows the idea of Le Masson and Sahlsten [LS17] in that we divide the argument into a spectral and a geometric component.
+However, we employ a different propagator, namely the hyperbolic wave propagator
+$P_{t}=h_{t}\!\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right),\qquad h_{t}(x)=x^{-1}\sin(tx).$
+The intuition behind $P_{t}$ is that $u_{tt}=-\Delta\,u$, the wave equation in the hyperbolic plane, has solution $u=P_{t}\,f$ with initial data
+$u(z,0)=0,\qquad u_{t}(z,0)=f,$
+for some function $f$. For details, see [LP82]. A convenient form of the integral kernel of $P_{t}$ is established in Section 5.
+Unlike the ball-averaging operator used by Le Masson and Sahlsten [LS17], which does not have as immediate a connection to the underlying classical dynamics, the hyperbolic wave propagator is obtained from the quantization of the geodesic flow, thus linking the quantum and classical settings in a natural manner.
+This choice of propagator renders the spectral side particularly clean, though it makes the geometric side more involved compared to the approach of Le Masson and Sahlsten [LS17].
+To access the off-diagonal averages $\langle\psi_{j},a\psi_{k}\rangle$, where the energy levels of $\psi_{j}$ and $\psi_{k}$ may differ substantially, we introduce a modified propagator
+$P_{t,\tau}=\cos(t\tau)\,P_{t}.$
+Using $P_{t}$ and $P_{t,\tau}$ simultaneously allows us to capture the behavior of these off-diagonal averages.
+However, this complicates both the spectral and geometric analyses compared to the diagonal case studied by Le Masson and Sahlsten [LS17].
+In particular, the Heisenberg evolution of the observable $a$, $P_{t}^{*}aP_{t}$, which appears in the proof of quantum ergodicity by Le Masson and Sahlsten, is replaced here by $P_{t,\tau}^{*}aP_{t}$.
+The physical interpretation of this operator is not immediately transparent, but it can be formally verified that
+$P_{t,\tau}^{*}aP_{t}=\cos(t\tau)\,P_{t}^{*}aP_{t}=2\cos\!\left(t\frac{\tau}{2}\right)P_{t}^{*}a\cos\!\left(t\frac{\tau}{2}\right)P_{t}-P_{t}^{*}aP_{t}=2P_{t,\frac{\tau}{2}}^{*}aP_{t,\frac{\tau}{2}}-P_{t}^{*}aP_{t},$
+where we used the identity $\cos(2x)=2\cos^{2}(x)-1$.
+This shows that $P_{t,\tau}^{*}aP_{t}$ can be written as the subtraction of two Heisenberg evolutions of $a$, one with $P_{t,\frac{\tau}{2}}$ and one with $P_{t}$.
+Finally, we replace the rather abstract Nevo’s theorem with the exponential mixing theorem (see Theorem 1.2), which elucidates the connection between classical and quantum dynamics.
+To use the exponential mixing theorem, we open the squares of the Hilbert–Schmidt norm appearing in the proof, as opposed to Le Masson and Sahlsten [LS17], who opt to use Minkowski’s integral inequality.
+In this way, we lose less decay and see an interesting exponential decorrelation of Heisenberg evolutions $P_{t}^{*}aP_{t}$ and $P_{t^{\prime}}^{*}aP_{t^{\prime}}$ when $|t-t^{\prime}|$ increases.
+The idea of using exponential mixing originates in the upcoming article of the author with Lequen, Mikkelsen, Sahlsten, and Ueberschär [30].
+The use of exponential mixing is both simple and transparent: the exponential decay of correlations of $a$ and $a\circ\varphi_{t}$, even for small $t$, is essential in obtaining our result.
+Small propagation times suffice because we work with observables in a macroscopic sense, unlike the large-energy quantum ergodicity and quantum mixing setting, which employs microlocal analysis and requires long propagation times for decorrelation to occur at the microlocal scale.
+We already refer to Section 3 for a more technical overview of the proofs.
 1.3. Further discussions
@@
 This assumption plays a crucial role in our proofs.
-However, it is plausible that with a different treatment of the low-lying modes—namely, those eigenvalues and corresponding eigenfunctions with eigenvalues in [0,1/4)[0,1/4)—one could extend the results to an energy window covering the full range $0,\infty)$.
+However, it is plausible that with a different treatment of the low-lying modes, namely, those eigenvalues and corresponding eigenfunctions with eigenvalues in $[0,1/4]$, one could extend the results to an energy window covering the full range $0,\infty)$.
+Another avenue of extension concerns the choice of observables.
+In Theorem 1.1, the observable class was quite broad, zeroth-order pseudo-differential operators, so it seems that with some modifications, a larger class of observables could be accommodated in the large-scale setting as well.
+In the large-scale quantum ergodicity setting for graphs [AL15] and surfaces [ABL23], a broader class of observables is already considered when treating the diagonal averages (Property (i)): namely, observables whose integral kernels are concentrated near the diagonal, rather than being strictly supported on it as in the present work.
 In contrast to Theorem 1.1, Theorem 1.3 is not a characterization-type result linking the classical and quantum dynamics.
@@
 As noted in Remark 1.6, the sequences of surfaces may in fact possess slowly decaying injectivity radii and spectral gaps.
-Given the connection between weak mixing and the spectral gap, it is plausible to conjecture that a sequence of hyperbolic manifolds converging to a hyperbolic space in the Benjamini–Schramm sense is large-scale quantum weak mixing if and only if the spectral gap tends to zero sufficiently slowly.
-Establishing such characterizations for both large-scale quantum ergodicity and large-scale quantum mixing would be remarkable.
-While full characterization results may still be out of reach, the present results appear extendable—albeit with varying levels of difficulty.
-Several directions for possible extensions present themselves.
+Given the connection between weak mixing and the spectral gap, it is plausible that a sequence of hyperbolic manifolds converging to a hyperbolic space in the Benjamini–Schramm sense is large-scale quantum mixing if and only if the spectral gap tends to zero sufficiently slowly.
+Establishing such characterizations for both large-scale quantum ergodicity and large-scale quantum mixing would be remarkable for further illustrating the connections between the classical and quantum settings.
+While full characterization results may still be out of reach, several directions for possible extensions present themselves, albeit with varying levels of difficulty.
 For instance, can analogous results be established in other structures such as graphs?
-We conjecture that the graph result of Brooks, Lindenstrauss, and Le Masson [BLL16] could be extended to include off-diagonal elements by adapting the techniques developed here, thereby enabling a proof of large-scale weak mixing in the graph setting of Anantharaman and Le Masson.
-Doing quantum mixing on the graphs seems highly timely as Bordenave, Letroit, and Sabri have an upcoming result on mixing on growing Schreier graphs [BLS].
-Another avenue of extensions concerns the choice of observables: in Theorem 1.1, the observable class was quite broad—zeroth-order pseudodifferential operators—so it seems likely that with minor modifications, a larger class of observables could be accommodated in the large-scale setting as well.
-This has already been achieved in the large-scale quantum ergodicity contexts for graphs [AL15] and surfaces [ABL23], where the kernels of the observables are concentrated near the diagonal but not restricted to it, unlike in the present work.
-Yet another natural question is whether the theorem can be adapted to the case of variable curvature, or to accommodate more general energy operators than the Laplacian.
-For graphs, analogous results are known: Anantharaman and Sabri [AS17, AS19] established large-scale quantum ergodicity for non-regular graphs with general Schrödinger operators, under suitable restrictions.
+We conjecture that the graph result of Brooks, Lindenstrauss, and Le Masson [BLL16] could be extended to include off-diagonal averages in the setting of the large-scale quantum ergodicity theorem for graphs by Anantharaman and Le Masson [AL15] by using the ideas presented here.
+Performing quantum mixing on graphs seems highly timely, as Bordenave, Letrouit, and Sabri have a recent result on quantum mixing on growing Schreier graphs [BLS26].
+Yet another natural question is whether the results can be adapted to the case of variable curvature or to accommodate more general energy operators than the Laplacian.
+For graphs, analogous results are known: Anantharaman and Sabri [AS17, AS19] established large-scale quantum ergodicity for non-regular graphs with more general Schrödinger operators than the Laplacian, under suitable restrictions.
 Non-regular graphs can be viewed as discrete analogues of variable-curvature surfaces.
-It would also be of great interest to construct counterexamples—sequences of surfaces demonstrating that the conditions imposed on the sequences are indeed necessary.
-In the graph setting, the results of Anantharaman and Le Masson were complemented in this manner by MacKenzie [McK22]; however, constructing such examples in the surface setting remains substantially more difficult.
+It would also be of great interest to construct counterexamples that would highlight the necessity of the conditions imposed on the sequences, namely Properties (BSC), (EXP), and (UND).
+In the graph setting, the results of Anantharaman and Le Masson were complemented in this manner by McKenzie [MCK22]; however, constructing such examples in the surface setting seems substantially more difficult.
 We pose perhaps an obvious question: is it possible to remove the averaging present in our theorems?
 In the large-energy regime, this corresponds to the quantum unique ergodicity conjecture of Rudnick and Sarnak, see [RS94].
-The conjecture has been resolved in certain special cases (notably by Lindenstrauss for arithmetic surfaces [Lin06]), but remains open in general.
-Analogous large-scale results are known for graphs; for instance, large-scale quantum unique ergodicity for random regular graphs was proved by Bauerschmidt et al.
-[BHY19].
-We conjecture that the conclusions of Theorem 1.4 remain valid even without averaging.
-We refer to such statements as large-scale quantum unique ergodic and large-scale quantum unique mixing.
-These would parallel the averaging-free results of Cipolloni et al.
-[CES21].
-A possibility is to try to establish this first using random models, such as Weil-Petersson model, or by using arithmetic surfaces and thus complement Lindenstrauss [Lin06].
-In particular, for arithmetic surfaces, averaging-free off-diagonal bounds are of significant interest: strong estimates on $|\langle\psi_{j},\psi_{k}\psi_{l}\rangle|$ could contribute to a proof of the Lindelöf hypothesis, to which quantum chaos is elegantly connected via the Watson $L$-function [Wat02].
-In any case, a deeper understanding of averaging-free results may provide a pathway toward results in many-body quantum chaos—such as the Eigenstate Thermalization Hypothesis proposed by Deutsch in the physics literature [Deu18].
-Deutsch predicts that $|\langle\psi_{j},a\psi_{k}\rangle|$ goes to zero as the distance of the eigenvalues of $\psi_{j}$ and $\psi_{k}$ goes to infinity.
-Looking at the large-scale results discussed here, Theorems 1.4 and 1.7, we see that parameter $\tau$ controls the distance of the eigenvalues.
-However, the eigenvalues are taken from the fixed energy windows which means that for large enough $\tau$, the averages of the off-diagonal elements will be zero as the large distance is not possible in the energy window.
-However, if we let the energy window grow, then $\tau$ can also grow without giving trivial conclusions.
-It is natural to assume that in this setting of growing energy window and growing $\tau$, the average will close to zero, but verifying this requires careful study of the sequence of growinf surfaces, $\tau$, and the energy window.
+The conjecture has been resolved in certain special cases (notably by Lindenstrauss for arithmetic surfaces and Hecke eigenfunctions [LIN06]), but remains open in general.
+We expect that the conclusions of Theorem 1.3 and Theorem 1.4 remain valid even without averaging, possibly requiring only minor adjustments to the statements.
+Especially, the averaging-free version of the probabilistic Theorem 1.4 would parallel the averaging-free results for random graphs [BHK+17a, BHY19, BKY17b] and random matrices [CES21].
+However, the averaging-free version of the deterministic Theorem 1.3 would more closely parallel the quantum unique ergodicity conjecture; it is likely harder to establish than the probabilistic version.
+For arithmetic surfaces, averaging-free off-diagonal bounds are of significant interest: strong estimates on $|\langle\psi_{j},\psi_{k}\psi_{l}\rangle|$ could contribute to a proof of the Lindelöf hypothesis, to which quantum chaos is elegantly connected; see Watson’s work on $L$-functions [WAT02].
+In any case, a deeper understanding of averaging-free results may provide a pathway toward results in many-body quantum chaos, such as the Eigenstate Thermalization Hypothesis proposed by Deutsch in the physics literature [DEU18].
+Deutsch predicts that $|\langle\psi_{j},a\psi_{k}\rangle|$ goes to zero as the distance between the eigenvalues of $\psi_{j}$ and $\psi_{k}$ increases.
+Looking at the large-scale results discussed here, Theorems 1.3 and 1.4, we see that the parameter $\tau$ controls the distance of the eigenvalues.
+However, the eigenvalues are taken from fixed energy windows, which means that for large enough $\tau$, the averages of the off-diagonal terms will be zero, as the large distance is not possible in the energy window.
+If we let the energy window grow, then $\tau$ can also grow without giving trivial conclusions.
+It is natural to assume that in this setting of growing energy window and growing $\tau$, the average will be close to zero, but verifying this requires careful study of the sequence of growing surfaces, the parameter $\tau$, and the energy window.
 2. Preliminaries
-In this section, we recall some basic concepts and results that will be used throughout this article, without proofs.
+In this section, we recall some basic concepts and results that will be used throughout this article without proofs.
 For background on hyperbolic geometry concerning the hyperbolic plane and hyperbolic surfaces, we refer to Katok [KAT92], and for details on the Laplace–Beltrami operator, to Iwaniec [IWA02].
-2.1. The hyperbolic plane and hyperbolic surfaces
+2.1. Hyperbolic plane
 The hyperbolic plane $\mathbb{H}$ is identified with the upper half-plane
@@
 $d\mu(x+iy)=y^{-2}\,dx\,dy,$
-and we write $|A|$ for the hyperbolic volume of a set $A\subseteq\mathbb{H}$.
-The group $PSL_{2}(\mathbb{R})$, consisting of $2\times 2$ real matrices with determinant one, modulo $\pm I$, acts on $\mathbb{H}$ by Möbius transformations:
+and we denote by $|A|$ the hyperbolic volume of a set $A\subset\mathbb{H}$.
+The group $PSL_{2}(\mathbb{R})$, consisting of $2\times 2$ real matrices with determinant one, modulo $\pm\mathrm{id}$, acts on $\mathbb{H}$ by Möbius transformations:
 $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot z=\frac{az+b}{cz+d}.$
@@
 $T^{1}\mathbb{H}=\{(z,v)\in\mathbb{H}\times T_{z}\mathbb{H}:\|v\|_{z}=1\}.$
-The group $PSL_{2}(\mathbb{R})$ acts on $T^{1}\mathbb{H}$ via
+The group $PSL_{2}(\mathbb{R})$ acts on $T^{1}\mathbb{H}$ via the action
 $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot(z,v)=\left(\frac{az+b}{cz+d},\frac{v}{(cz+d)^{2}}\right).$
-Evaluating this action at the point $(i,i)$ shows that $PSL_{2}(\mathbb{R})$ is homeomorphic to $T^{1}\mathbb{H}$; hence, these two spaces can be identified.
+Evaluating this action at the point $(i,i)$ shows that $PSL_{2}(\mathbb{R})$ is homeomorphic to $T^{1}\mathbb{H}$. Thus these two spaces can be identified.
 The geodesic flow is the one-parameter family of maps
@@
 $A=\left\{a_{t}=\begin{pmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{pmatrix}:t\in\mathbb{R}\right\}.$
-We denote by $z,z^{\prime}]$ the geodesic segment connecting two distinct points $z,z^{\prime}\in\mathbb{H}$, by $z,z^{\prime}](w)$ the tangent direction at a point $w\in[z,z^{\prime}]$, and by $|[z,z^{\prime}]|$ its length.
 For any distinct points $z,z^{\prime}\in\mathbb{H}$, there exists a unique geodesic of length $r$ from $z$ to $z^{\prime}$.
@@
 $ds^{2}=dr^{2}+\sinh^{2}(r)\,d\theta^{2},\qquad d\mu(r,\theta)=\sinh(r)\,dr\,d\theta.$
-Then a point $x\in\mathbb{H}$ can be written as
+Point $x\in\mathbb{H}$ can be written as
 $x=\pi_{1}(\phi_{r}(z,\theta)),$
-where the polar coordinates are around $z$; $r$ and $\theta$ depend on $x$, and $\pi_{1}$ is the projection onto the first element of pair in $T^{1}\mathbb{H}$.
-The Laplacian on $\mathbb{H}$ is the differential operator
+where the polar coordinates are around $z$; $r$ and $\theta$ depend on $x$, and $\pi_{1}$ is the projection onto the first element of the pair in $T^{1}\mathbb{H}$.
+We denote by $z,z^{\prime}]$ the geodesic segment going from $z\in\mathbb{H}$ to another point $z^{\prime}\in\mathbb{H}$.
+By $z,z^{\prime}](w)$ we denote the direction of the tangent vector of the geodesic at point $w\in[z,z^{\prime}]$, and by $|[z,z^{\prime}]|$ we denote the length of the geodesic segment.
+The Laplace–Beltrami operator on $\mathbb{H}$ is the differential operator
 $\Delta=-y^{2}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right),$
-which commutes with all isometries of $\mathbb{H}$.
+which commutes with all isometries of $\mathbb{H}$. We refer to it as the Laplacian operator.
 2.2. Hyperbolic surfaces
@@
 $X=\Gamma\setminus\mathbb{H},$
-where $\Gamma\subset PSL_{2}(\mathbb{R})$ is a discrete subgroup consisting solely of hyperbolic elements, that is, matrices whose trace satisfies $|\mathrm{tr}(\gamma)|>2$.
-In this way, $X$ can be represented by a fundamental domain in $\mathbb{H}$.
+where $\Gamma\subset PSL_{2}(\mathbb{R})$ is a discrete subgroup consisting solely of hyperbolic elements, that is, matrices whose trace satisfies $|\mathrm{tr}(\gamma)|>2$; $\Gamma$ is a Fuchsian group.
+In this way, $X$ can be represented by a compact subset of $\mathbb{H}$ called a fundamental domain.
 A function on $X$ can equivalently be regarded as a $\Gamma$-periodic function on $\mathbb{H}$, and vice versa.
-Moreover, the integral of a function $f$ over $X$ equals the integral of the corresponding $\Gamma$-invariant function over a fundamental domain.
+Moreover, the integral of a function $f$ over $X$ equals the integral of the corresponding $\Gamma$-invariant function over a fundamental domain of $X$.
 For a hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, the unit tangent bundle can be identified with $\Gamma\setminus PSL_{2}(\mathbb{R})$, and under this identification the geodesic flow takes the form
@@
 and the (global) injectivity radius of $X$ is
-$\operatorname{InjRad}_{X}:=\sup_{z\in X}\operatorname{InjRad}_{X}(z).$
-Since the Laplacian on $\mathbb{H}$ commutes with isometries, it descends naturally to a Laplacian on the quotient $X=\Gamma\setminus\mathbb{H}$.
-To distinguish the two, we denote by $\Delta_{\mathbb{H}}$ the Laplacian acting on all smooth functions on $\mathbb{H}$, and by $\Delta_{X}$ the Laplacian acting on $\Gamma$-invariant functions.
+$\operatorname{InjRad}_{X}:=\inf_{z\in X}\operatorname{InjRad}_{X}(z).$
+Since the Laplacian on $\mathbb{H}$ commutes with isometries, it descends naturally on the quotient $X=\Gamma\setminus\mathbb{H}$.
+To distinguish the two, we denote by $\Delta_{\mathbb{H}}$ the Laplacian acting on $\mathbb{H}$, and by $\Delta_{X}$ the Laplacian acting on $\Gamma$-invariant functions.
 On a compact connected hyperbolic surface, the Laplacian $\Delta_{X}$ has a discrete spectrum
 $0=\lambda_{0}<\lambda_{1}\leq\lambda_{2}\leq\dotsb\xrightarrow{j\to\infty}\infty.$
-3. Strategy of proofs
-Three main theorems, Theorems 1.4, 1.5, and 1.7, are proven in Section 8.
-In this section we give a rough, overarching idea of the approach taken in the proofs.
-Most of the paper is dedicated for trying to establish Theorem 1.5.
-Proving this theorem is technical so the proof is split into multiple sections.
-These sections, Sections 4, 5, 6, and 7, are presented before the Section 8 so that the auxiliary results have been covered when the main theorems are under focus.
-In Theorem 1.5 we want to bound the double sum appearing in Expression (3.3) for a fixed surface $X=\Gamma\setminus\mathbb{H}$ under suitable conditions. Hence we want to bound the following:
-$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}$ (3.1)
-We can split the inner product as follows:
-$|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,\psi_{k}\,dt\right\rangle\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|},$ (3.2)
-when
+The spectral gap of $\Delta_{X}$ is the difference between its two smallest eigenvalues.
+3. Proof ideas
+Three main theorems, Theorems 1.3, 1.4, and 1.5, are proven in Section 8 using the auxiliary results established in Sections 4, 5, 6, and 7.
+In this section we give an overarching idea of all the proofs to provide a coherent narrative; we refer to the latter sections for the exact definitions of the appearing quantities.
+We remark that in the actual proof, the coefficients are written carefully to make all dependencies explicit; here, we adopt a less explicit presentation.
+The goal is to prove Properties (ii) and (iii) of Theorems 1.3 and 1.4.
+We prove the properties simultaneously: We want to show that for any $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for any $\tau\in\mathbb{R}$ we have
+$\limsup_{n\to\infty}\frac{1}{N(X_{n},I)}\,\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle|^{2}<\epsilon.$ (3.1)
+The setting is deterministic in Theorem 1.3 and probabilistic in Theorem 1.4.
+To achieve this goal, we aim to bound the double sum for a fixed surface $X$, which is the content of the quantitative main theorem, Theorem 1.5.
+We proceed with the idea of the proof of Theorem 1.5. We set out to bound the following expression:
+$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}.$ (3.2)
+We can rewrite the inner product as follows:
+$|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})h_{t}(\rho_{k})dt\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})h_{t}(\rho_{k})dt\right|}\,|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,\psi_{k}\,dt\right\rangle\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|},$ (3.3)
+where
 $P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right),\qquad h_{t}(x)=\frac{\sin(tx)}{x},$
@@
 $P_{t,\tau}=h_{t,\tau}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right),\qquad h_{t,\tau}=\cos(t\tau)\,h_{t}.$
-This split could be referred to as geometric-spectral split, since the denumerator contains the spectral data of the problem and the numerator the geometric data of the problem.
-$P_{t}$ is the hyperbolic wave propagator (see [LP82]).
-It is the quantization of the geodesic flow, hence a very natural object.
-Using this propagator makes the spectralside easy but complicates the analysis of the geometric side when comparing the setting to that of Le Masson and Sahlsten (see [LS17]) who used a ball averaging propagator which is not as natural as the wave propagator.
-$P_{t,\tau}$, a modified version of $P_{t}$, is used simultaneously with $P_{t}$ so that we can access the off-diagonal elements.
-This adds another layer of complexity to the proof when comparing to the setting in which only the diagonal elements are considered.
-The denumerator of Expression (3.2) contains the spectral data of the problem.
-We acquire bounds for the denominator in Section 4.
-Using the bounds acquired there gives an upper bound for Expression (3.1):
-$64\pi^{2}\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2},$
-where the Hilbert-Schmidt norm upper bound was acquired by enlargening the sums to include all the basis elements after the spectral bound was pulled out of the sum.
-Now the Hilbert-Schmidt norm contains the geometric data of the problem.
-Bouding the Hilbert-Schmidt norm is very technical.
-In Section 5 we establish convenient forms for the integral kernels of $P_{t}$ and $P_{t,\tau}$.
-Using the acquired formulas, we bound the Hilbert-Schmidt norm in Section 6.
-This is based on carefully bounding the Hilbert-Schmidt norm expression and finally splitting it into two cases.
-In one we consider an integral over points with small injectivity radius and in the other an integral over points with large injectivity radius.
-To bound these two cases, we especially need the exponential mixing theorem, Theorem 1.2, and bounds for certain geometric weight $F_{t,t^{\prime},\rho}$ which appears in both the small and large injectivity radius cases, containing the core geometric information of the problem.
-Bounding $F_{t,t^{\prime},\rho}$ is done in Section 7.
-We remark that the exponential mixing theorem replaces the Nevo’s theorem appearing in the same role in [LS17] as the exponential mixing here.
-With the Hilbert-Schmidt norm bounded, we acquire the final bounds for Expression (3.1) to conclude Theorem 1.5:
-$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq E\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
-where we refer to Theorem 1.5 for the quantities appearing in the upper bound.
-The limit version of the main theorems, Theorem 1.4, is proven using Theorem 1.5 together with the large-scale Weyl law of Le Masson and Sahlsten, Theorem 8.1.
-In the proof we want to prove that for any $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for any $\tau\in\mathbb{R}$ we have that
-$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle|^{2}\right)<\epsilon.$ (3.3)
-Proving this proves only conditions (ii) and (iii); Le Masson and Sahlsten’s proof of quantum ergodicity covers case (i), see [LS17].
-We remark here that the techniques used here can also be used to give a new proof for the quantum ergodicity theorem.
-Using the bounds acquired for the double sum in Theorem 1.5 gives the following upper bound for the previous expression:
-$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,E\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}(X_{n}))^{3}}\,\left(\|a_{n}\|_{2}^{2}\,+\,8\pi^{2}\|a_{n}\|_{\infty}^{2}\,|D_{n}\setminus D_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right)\right),$
-where we refer to Theorem 1.4 for the quantities appearing in the upper bound.
-Using property (BSC) helps us to get a universal lower bound for $\beta(\lambda_{1}(X_{n}))$ and using property (UND) helps us to get a universal lower bound for $\operatorname{InjRad}_{X_{n}}$.
-Weyl law together with property (BSC) shows that
-$\frac{|D_{n}\setminus D_{n}(4T)|}{|N(X_{n},I)|}\xrightarrow{n\to\infty}0,$
-as the Weyl law relates the growth of the number of eigenvalues in the energy window $I$ to that of the volume of the surface, $|X_{n}|$, and the ratio of the points with small injectivity radius and the volume of the surface goes to zero due to property (BSC).
-The core idea is to use properties (BSC), (EXP), and (UND) together with Weyl law to remove the limit supremum from the upper bound.
-After this, we have an upper bound in which $\delta$ appears.
-We can choose a suitable $\delta$ to get $\epsilon$ as an upper bound to conclude the proof.
-To finally prove the probabilistic main theorem, Theorem 1.7, we take a very similar approach as with the proof of Theorem 1.4 as these theorems are very similar.
-We only prove Properties (ii) and (iii) of Theorem 1.4 as Le Masson and Sahlsten have already proved Property (i), see [LS24].
-We remark that in probabilistic setting we do not impose Properties (BSC), (EXP), and (UND) and hence we do not have the large-scale Weyl law at out disposal.
-To remedy this, we get similar properties from the theory of random surfaces.
-These are recorded in Theorem 8.2.
-We can use these properties in the place of Properties (BSC), (EXP), and (UND) in a very similar way to get the conclusion to hold with high probability as we proceed in the sequence.
-To conclude the section, we remark that the main novelty of this paper method-wise is the use of the exponential mixing theorem, Theorem 1.2, instead of using Nevos’s theorem to get the needed extra decay to establish Theorem 1.5.
-Another novelty is to separate a weight arising from the choice of propagator, denoted here by $F_{t,t^{\prime},\rho}$, which gives much required nimbleness in the proof; it can also be approximated separately without the observable function $a$ interfering.
+$T$ depends on $\delta$ via $T\delta=\tfrac{\pi}{2}$.
+We refer to this as the geometric-spectral split, since the denominator contains the spectral data of the problem and the numerator contains the geometric data.
+Two different operators, $P_{t}$ and $P_{t,\tau}$, are used because we are dealing with off-diagonal averages where the separation of the eigenvalues is significant.
+The denominator of Expression (3.3) contains the spectral data. We acquire the following bounds for the denominator in Section 4:
+$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|^{-1}\leq 8\pi\rho_{j}\rho_{k}\leq 8\pi\max(I)^{2}.$
+The bound is obtained by carefully integrating and using the interrelations of $\delta,T,\rho_{j},$ and $\rho_{k}$.
+Since we are working in a compact energy window, $\rho_{j},\rho_{k}\leq\max(I)$.
+It is especially important that $\delta\in(0,\tfrac{2}{9}\min I)$.
+Being able to choose such a positive $\delta$ is crucial; this is why we have excluded the low-lying notes, i.e., those eigenfunctions of the Laplacian $\Delta_{X}$ with eigenvalues in the window $0,\tfrac{1}{4}]$.
+Using the geometric-spectral split and the bounds for the spectral data gives an upper bound for Expression (3.2):
+$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq 64\pi^{2}\max(I)^{4}\,\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}dt\,\psi_{k}\right\rangle\right|^{2},$
+where this has the following Hilbert-Schmidt norm upper bound by enlarging the double sum to include all basis elements:
+$64\pi^{2}\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}dt\right\|_{HS}^{2}.$
+The Hilbert-Schmidt norm contains the geometric data of the problem; it contains no eigenvalues or eigenfunctions and is now purely geometric.
+In Section 5 we establish convenient forms for the integral kernels of $P_{t}$ and $P_{t,\tau}$ which are used to bound the Hilbert-Schmidt norm in Section 6.
+$P_{t}$ has integral kernel $K_{t}^{\Gamma}$:
+$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\qquad K_{t}(x,y)=\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}.$
+$P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)K_{t}^{\Gamma}$.
+We notice that $P_{t}$ and $P_{t,\tau}$ are self-adjoint from these integral kernel formulas.
+Informally, the previous formulas for the integral kernels are obtained by using the inverse Selberg transform $\mathcal{S}^{-1}$ (e.g., see [LS17]):
+$K_{t}^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}\mathcal{S}^{-1}(h_{t})(x,\gamma y).$
+However, $h_{t}$ does not behave well enough for this to work immediately, and hence a more careful approach is needed; we approximate $P_{t}$ and $P_{t,\tau}$ to use analogous techniques.
+By using the acquired formulas for the kernels, we bound the Hilbert-Schmidt norm in Section 6.
+Opening the Hilbert-Schmidt norm using the obtained kernel formulas and carefully rewriting it gives the following equivalent expression:
+$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dzdt^{\prime}dt,$
+where $D$ is a fundamental domain of $X$. The previous expression has the following upper bound obtained by taking absolute values, moving them inside, and bounding $|\cos(\cdot)|$ by $1$:
+$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (3.4)
+Eliminating $\cos$ factors, which appear due to the simultaneous presence of $P_{t}$ and $P_{t,\tau}$, might lead to some loss of decay.
+A more careful approach could yield better estimates.
+$F_{t,t^{\prime},\rho}$ contains the core of the geometric data and is defined as follows:
+$F_{t,t^{\prime},\rho}=\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}.$
+Here $z,z^{\prime}\in\mathbb{H}$ are any points separated by $\rho$.
+$F_{t,t^{\prime},\rho}$ is studied in Section 7.
+An especially important fact is that $F_{t,t^{\prime},\rho}=0$ when $\rho>t+t^{\prime}$.
+In Expression (3.4), we split the integration over the fundamental domain $D$ into points with injectivity radius larger than $4T$, $D(4T)$, and into points with injectivity radius smaller than $4T$, $D\setminus D(4T)$.
+This approach leads to the following upper bound:
+$\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|$ + $+\ 2\,\|a\|_{\infty}^{2}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$ (3.5)
+For points with large injectivity radius, all group elements except $id\in\Gamma$ contribute zero, thus eliminating the sum.
+The latter term deals with points with small injectivity radius.
+The terms are treated separately.
+The first term of Expression (3.5) can be written using polar coordinates around $z$ and has the following upper bound:
+$\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}\left|\langle a,a\circ\phi_{\rho}\rangle\right|d\rho\leq 30\|a\|_{2}^{2}\int_{0}^{2T}\sinh(\rho)(1+\rho)e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2},$
+where the latter inequality is obtained using the exponential mixing theorem, Theorem 1.2.
+The exponential decay obtained via the exponential mixing theorem is perhaps the most crucial step in the entire proof.
+$\beta$ is a constant depending on the spectral gap.
+The second term of Expression (3.5) is bounded more roughly using Hölder’s inequality.
+The interplay of $F_{t,t^{\prime},d(z,z^{\prime})}$ and $F_{t,t^{\prime},d(\gamma z,z^{\prime})}$ ensures that only finitely many group elements contribute.
+The upper bound can be written as:
+$8\pi^{2}\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)(1+\rho)e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}.$
+Using the above bounds, we obtain an upper bound for Expression (3.4):
+$\frac{1}{T^{2}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
+This is straightforwardly bounded when we have the following bound for $F_{t,t^{\prime},\rho}$:
+$F_{t,t^{\prime},\rho}\lesssim\sinh(\max(\rho,|t-t^{\prime}|))^{-\tfrac{1}{2}}.$ (3.6)
+Obtaining this bound for $F_{t,t^{\prime},\rho}$ is the most technical part of the paper and is carried out in Section 7.
+The idea is to compute a weighted volume of the intersection of two hyperbolic balls using hyperbolic geometry and trigonometry.
+In practice, the integration is handled in multiple separate cases.
+The final upper bound is a constant times
+$\frac{1}{T\beta^{3}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
+where $|X\setminus X(4T)|$ denotes the set of points with small injectivity radius. This shows that the double sum Expression (3.2) is bounded using the Hilbert-Schmidt norm bound:
+$\max(I)^{4}\,\frac{1}{T\beta^{3}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right).$
+A constant factor is omitted. This establishes the proof idea of Theorem 1.5.
+We proceed to Theorems 1.3 and 1.4.
+We want to find an upper bound $\epsilon$ for a suitable $\delta$ for the left-hand side of Expression (3.1).
+Using the bounds from Theorem 1.5, we get the following upper bound (omitting a constant factor):
+$\max(I)^{4}\frac{1}{T}\,\limsup_{n\to\infty}\frac{1}{N(X_{n},I)}\frac{1}{\beta(\lambda_{1}(X_{n}))^{3}}\left(\|a_{n}\|_{2}^{2}+\|a_{n}\|_{\infty}^{2}|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right).$ (3.7)
+In the deterministic setting of Theorem 1.3, we use that $(a_{n})$ is uniformly bounded by $a_{max}$, $I$ is compact, (EXP) gives a universal lower bound for $(\beta(\lambda_{1}(X_{n})))$ denoted by $\beta_{min}$, and (UND) provides a uniform lower bound for $(\operatorname{InjRad}_{X_{n}})$ denoted by $\operatorname{InjRad}_{min}$.
+The upper bound is
+$\frac{\max(I)^{4}a_{max}^{2}}{\beta_{min}^{3}}\frac{1}{T}\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}+\frac{|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{N(X_{n},I)\,\operatorname{InjRad}_{min}}\right).$
+The large-scale Weyl law gives $N(X_{n},I)\sim|X_{n}|$. Using Property (BSC), the volume ratio of points with small injectivity radius shrinks to zero. Hence, the upper bound becomes
+$\frac{\max(I)^{4}a_{max}^{2}}{\beta_{min}^{3}}\frac{1}{T}.$
+Since $T\delta=\tfrac{\pi}{2}$, we may choose $\delta$ sufficiently small so that the bound is less than $\epsilon$. This completes the proof idea for Theorem 1.3.
+In the probabilistic setting of Theorem 1.4, a similar approach is taken; here $n=g$ is the genus of the surface.
+We do not have Properties (BSC), (EXP), (UND), or the large-scale Weyl law a priori.
+However, Theorem 8.2 provides analogous tools.
+For a fixed surface $X_{g}$ of genus $g$,
+$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
+$\leq\ \frac{|X_{g}|}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(1-\sqrt{1-4(\tfrac{1}{4}-\eta)}\right)^{3}}\,\left(1+\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
+with probability 1o(1)1-o(1), where $\eta\in(0,\tfrac{1}{4})$ and
+$J=\left\{x^{2}+\tfrac{1}{4}\ :\ x\in I\right\}.$
+Hence Expression (3.7) equals, with high probability in the high-genus limit (up to a constant factor),
+$a_{max}^{2}\max(I)^{4}\frac{1}{T}.$
+Since $T\delta=\tfrac{\pi}{2}$, we may choose $\delta$ sufficiently small so that the bound is less than $\epsilon$. This completes the proof idea for Theorem 1.4.
 4. Spectral data
-In this section we prove a bound for spectral quantity
+In this section, we prove a bound for the spectral quantity
 $\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}$
-under certains assumptions on $a,b,t,\tau,$ and $T$, given that
+under certain assumptions on $a,b,t,\tau,$ and $T$, given that
 $h_{t}(x)=x^{-1}\,\sin(tx),\qquad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x).$
-The spectral quantity appears in the proof of Theorem 1.5 and there it contains the spectral data of the problem.
-We state the proposition.
+This spectral quantity appears in the proof of Theorem 1.5, where it encodes the spectral data of the main problem. We state the proposition.
 Proposition 4.1 (Spectral data).
@@
 $a-b-\tau\in(-\delta,\delta),$
-where $0<\delta<\tfrac{2}{9}\,m$ and where $T\,\delta=\pi/2$. Then
+where $\delta\in(0,\tfrac{2}{9}m)$. Then for $T\delta=\tfrac{\pi}{2}$ we have that
 $\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}<8\pi ab.$
-The proof consists of relatively straightforward integration and careful application of the relationships of the variables and coefficients.
-The simplicity follows from a convenient choice of $h_{t}$ and $h_{t,\tau}$.
+The proof consists of straightforward integration and careful application of the relationships among the variables. The simplicity follows from the convenient choices of $h_{t}$ and $h_{t,\tau}$.
 Proof of Proposition 4.1.
@@
 $G_{\tau,T}=\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|$
-from below to prove the bound. $G_{t,\tau}$ is illustrated in Figure 1.
+from below. $G_{\tau,T}$ is illustrated in Figure 1.
 Figure 1.
@@
 Two pronounced “ridges” are visible along the lines $b=a+\tau$ and $b=a-\tau$.
-The ridges get pronounced as $T$ increases.
-For a fixed $\delta$, $G_{\tau,t}(a,b)$ has a positive lower bound when $a-b-\tau\in(-\delta,\delta)$.
-Notice also how the ridges decay as $a$ and $b$ grow.
-By using the definitions of $h_{t}$ and $h_{t,\tau}$, we write $G_{\tau,T}(a,b)$ as
+The ridges become more pronounced as $T$ increases and decay as $a$ and $b$ increase.
+Using the definitions of $h_{t}$ and $h_{t,\tau}$, we expand $G_{\tau,T}(a,b)$ to
 $\frac{1}{Tab}\left|\int_{0}^{T}\cos(t\tau)\sin(ta)\sin(tb)dt\right|.$ (4.1)
-By using trigonometric sum-to-product identities, we have that $\cos(t\tau)\sin(ta)$ equals
-$\frac{1}{2}\left(\sin\left(t(\tau+a)\right)-\sin\left(t(\tau-a)\right)\right),$
-and further that $\cos(t\tau)\sin(ta)\sin(tb)$ equals
-$\frac{1}{4}\left(\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\right).$
-Using this equality, we have that Expression (4.1) equals
-$\frac{1}{4Tab}\left|\int_{0}^{T}\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\,dt\right|.$
-The sine terms of the integral can all be integrated separately and thus the integral yields the following equal expression:
-$\frac{1}{4Tab}\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}-\frac{\sin(T(a+b-\tau))}{a+b-\tau}-\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|$
-Applying the reverse triangle inequality yields a following lower bound for the previous expression:
+Applying trigonometric sum-to-product identities, we have
+$\cos(t\tau)\sin(ta)=\frac{\sin\left(t(\tau+a)\right)-\sin\left(t(\tau-a)\right)}{2},$
+and hence
+$\cos(t\tau)\sin(ta)\sin(tb)=\frac{\cos(t(a-b-\tau))+\cos(t(a-b+\tau))-\cos(t(a+b-\tau))-\cos(t(a+b+\tau))}{4}.$
+Therefore, Expression (4.1) equals
+$\frac{1}{4Tab}\left|\int_{0}^{T}\cos(t(a-b-\tau))+\cos(t(a-b+\tau))-\cos(t(a+b-\tau))-\cos(t(a+b+\tau))\,dt\right|.$
+Integrating each cosine term separately gives
+$\frac{1}{4Tab}\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}-\frac{\sin(T(a+b-\tau))}{a+b-\tau}-\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|.$
+Applying the reverse triangle inequality yields a lower bound:
 $\frac{1}{4Tab}\bigg(\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}\right|-\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|-\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|\bigg).$ (4.2)
-We proceed to bound the absolute value terms separately. Consider the first absolute value term of the previous expression. Consider function
-$x\mapsto\frac{\sin(T\,x)}{x}.$
-It is positive when $x\in\left(-\pi/T,\pi/T\right)$, and it also decreases as $|x|$ grows in this interval.
-Because $T\delta=\pi/2$ and $a-b-\tau\in(-\delta,\delta)$, we can apply the observations regarding the previously studied function to get that
+We now bound the absolute value terms separately. Consider the first absolute value. The function
+$x\mapsto\frac{\sin(T\,x)}{x}$
+is positive when $x\in\left(-\frac{\pi}{T},\frac{\pi}{T}\right)$, and decreases as $|x|$ grows in this interval. Since $T\delta=\frac{\pi}{2}$ and $a-b-\tau\in(-\delta,\delta)$, we have
 $\frac{\sin\!\big(T(a-b-\tau)\big)}{a-b-\tau}\geq\frac{\sin(T\delta)}{\delta}=\frac{1}{\delta}.$
-We also notice that
-$\frac{\sin(Tx)}{x}>0$
-when $x\in\left(-\pi/T,\pi/T\right)$, which means that for every $x\in\mathbb{R}$ we have that
+Moreover, for all $x\in\mathbb{R}$:
 $\frac{\sin(Tx)}{x}\geq-\frac{1}{\pi/T}=-\frac{1}{2\delta}.$
-The previous two observations mean that
-$\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}$
-is necessarily positive and hence
-$\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}\right|=\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}>\frac{1}{2\delta}.$
-Next, consider the second and third absolute value terms of Expression (4.2). We observe that
+Thus,
+$\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}>\frac{1}{2\delta}.$
+Next, consider the second and third absolute value terms in Expression (4.2). Since
 $\left|\frac{\sin(Tx)}{x}\right|\leq\frac{1}{|x|},$
-implying that
+we have
 $\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|\leq\frac{1}{|a+b-\tau|}.$
-We also observe that
+Moreover,
 $|a+b-\tau|\geq 2b-|a-b-\tau|>2m-\delta>8\delta,$
-where we used the reverse triangle inequality together with the fact that $|a-b-\tau|$ is smaller than $\delta$ and that $2m$ is larger than $9\delta$. Using these observations gives that
-$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|\leq\frac{1}{|a+b-\tau|}<\frac{1}{8\delta}.$
-A similar approach yields that
-$\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|<\frac{1}{8\delta}.$
-These two observations mean that
+$|a+b+\tau|\geq 2a-|a-b-\tau|>2m-\delta>8\delta.$
+Hence,
+$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|<\frac{1}{8\delta},\qquad\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|<\frac{1}{8\delta},$
+so that
 $-\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|-\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|>-\frac{1}{4\delta}.$
-With all the absolute value terms bounded, Expression (4.2) has the following lower bound:
-$\frac{1}{4Tab}\,\frac{1}{4\delta}=\frac{1}{16\frac{\pi}{2}ab}=\frac{1}{8\pi ab}.$
-In the previous we used that $T\delta=\pi/2$.
-Combining these bounds gives that
-$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|\ \geq\ \frac{1}{4Tab}\cdot\frac{1}{4\delta}\ =\ \frac{1}{16\,\frac{\pi}{2}\,ab}\ =\ \frac{1}{8\pi\,ab}.$
-Therefore,
-$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}=G_{\tau,T}^{-1}<8\pi ab,$
-as claimed. ∎
+Combining these bounds, Expression (4.2) has the lower bound
+$\frac{1}{4Tab}\,\frac{1}{4\delta}=\frac{1}{16\frac{\pi}{2}ab}=\frac{1}{8\pi ab},$
+where we used $T\delta=\frac{\pi}{2}$. Hence,
+$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|>\frac{1}{8\pi ab},$
+and therefore
+$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}=G_{\tau,T}^{-1}<8\pi ab.$
 5. Propagators
@@
 $\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2},$
-the geometric data of the problem that appears in the proof of Theorem 1.5, it is necessary to analyze the kernels of the hyperbolic wave propagator $P_{t}$ and a slightly mofified version of it, $P_{t,\tau}$, for these kernels are then used to open the Hilbert-Schmidt norms.
-Although the hyperbolic wave propagator is a natural choice, it exhibits greater thechnical complexity than, for instance, the ball averaging operator (cf.
-[LS17]).
-As mentioned earlier, this choice makes the specral side easier.
+the geometric data of the problem that appears in the proof of Theorem 1.5, it is necessary to analyze the kernels of the hyperbolic wave propagator $P_{t}$ and a slightly modified version of it, $P_{t,\tau}$, since these kernels are then used to bound the Hilbert–Schmidt norms.
+Although the hyperbolic wave propagator is a natural choice, it exhibits greater technical complexity than, for instance, the ball averaging operator (cf.
+[LS17]); the choice is deliberate as it makes the spectral side easier.
 We state a proposition giving a convenient form for the integral kernels of $P_{t}$ and $P_{t,\tau}$.
@@
 $P_{t}=h_{t}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),\quad P_{t,\tau}=h_{t,\tau}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),$
-be defined on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$ with $\Delta_{X}$ being the Laplace-Beltrami operator and where
+be defined on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, with $\Delta_{X}$ being the Laplace–Beltrami operator, and where
 $h_{t}(x)=\frac{\sin(tx)}{x},\quad h_{t,\tau}(x)=\cos(t\tau)\,h_{t}(x).$
-Then $P_{t}$ admits integral kernel
+Then $P_{t}$ admits the integral kernel $K_{t}^{\Gamma}$ defined as
 $K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\quad K_{t}(x,y)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
-and $P_{t,\tau}$ admits integral kernel $\cos(t\tau)\,K_{t}^{\Gamma}$.
-Given that
-$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right)$
-and that $\tilde{P}_{t}$ is the operator with the integral kernel $K_{t}^{\Gamma}$ which was presented in the statement of the proposition, the proposition can be proved by showing that inner product $\langle P_{t}f,g\rangle$ equals inner product $\langle\tilde{P}_{t}f,g\rangle$ where $f,g\in C^{\infty}(X)$ as $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ with $X$ being compact.
-The core idea is representing $P_{t}$ as a limit of propagators $P_{t}^{(\varepsilon)}$ that have a convenient explicit formulae for their integral kernels.
-The proof relies on multiple auxiliary results presented after the proof of Proposition 5.1 for clarity’s sake
+and $P_{t,\tau}$ admits the integral kernel $\cos(t\tau)\,K_{t}^{\Gamma}$.
+Given that $\tilde{P}_{t}$ is the integral operator with integral kernel $K_{t}^{\Gamma}$, the proposition can be proved by showing that $\langle P_{t}f,g\rangle=\langle\tilde{P}_{t}f,g\rangle$ for $f,g\in C^{\infty}(X)$, since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$.
+The core idea of showing the equality of the inner products is representing $P_{t}$ as a limit of propagators $P_{t}^{(\varepsilon)}$ that have convenient explicit formulae for their integral kernels.
+The proof relies on multiple auxiliary results presented after the proof of Proposition 5.1.
 Proof of Proposition 5.1.
-The assertion for $P_{t,\tau}$ follows directly from the assertion for $P_{t}$ since $P_{t,\tau}=\cos(t\tau)P_{t}$. Hence it suffices to consider $P_{t}$ and $K_{t}$.
-Due to way $P_{t}$ is defined, it is bounded and has property
+The assertion for $P_{t,\tau}$ follows directly from the assertion for $P_{t}$ since $P_{t,\tau}=\cos(t\tau)P_{t}$. Hence, it suffices to consider $P_{t}$ and $K_{t}^{\Gamma}$.
+Due to the way $P_{t}$ is defined, it is bounded and has the property
 $P_{t}\psi_{j}=h_{t}(\rho_{j})\psi_{j},$
-where $\psi_{j}$ is an eigenfunction of the Laplacian on $X$ with eigenvalue $\lambda_{j}$ such that $\rho_{j}=\sqrt{\lambda_{j}-1/4}$. Define regularized operators for $\varepsilon>0$:
-$P_{t}^{(\varepsilon)}=h_{t}^{(\varepsilon)}\Big(\sqrt{-\Delta-\tfrac{1}{4}}\Big),\quad h_{t}^{(\varepsilon)}(x)=\frac{\sin(tx)}{x}\,e^{-\frac{\varepsilon^{2}x^{2}}{2}}.$
-Due to way $P_{t}^{(\epsilon)}$ is defined, it is bounded and has property
+where $\psi_{j}$ is an eigenfunction of $\Delta_{X}$ with eigenvalue $\lambda_{j}$ such that $\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$. Define regularized operators for $\varepsilon>0$:
+$P_{t}^{(\varepsilon)}=h_{t}^{(\varepsilon)}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),\quad h_{t}^{(\varepsilon)}(x)=\frac{\sin(tx)}{x}\,e^{-\frac{\varepsilon^{2}x^{2}}{2}}.$
+Due to the way $P_{t}^{(\varepsilon)}$ is defined, it is bounded and has the following property:
 $P_{t}^{(\varepsilon)}\psi_{j}=h_{t}^{(\varepsilon)}(\rho_{j})\,\psi_{j}.$
-Since $h_{t}^{(\varepsilon)}$ converges to $h_{t}$ pointwise as $\varepsilon$ shrinks, we have that $P_{t}^{(\varepsilon)}f$ converges to $P_{t}f$ pointwise as $\varepsilon$ shrinks for every $f\in C^{\infty}(X)$.
-Proposition 5.2 ensures the existence of integral operator $Q_{t}^{(\varepsilon)}$ on $X$ with integral kernel
+Since $h_{t}^{(\varepsilon)}$ converges to $h_{t}$ pointwise as $\varepsilon\to 0$, we have that $P_{t}^{(\varepsilon)}f$ converges to $P_{t}f$ pointwise for every $f\in C^{\infty}(X)$ due to spectral theory.
+Proposition 5.2 ensures the existence of an integral operator $Q_{t}^{(\varepsilon)}$ on $X$ with integral kernel $K_{t}^{(\varepsilon)}$ defined as
 $K_{t}^{(\varepsilon)}(x,y)=\sum_{\gamma\in\Gamma}k_{t}^{(\varepsilon)}(x,\gamma y),$
 $k_{t}^{(\varepsilon)}(x,y)=-\frac{1}{\pi i}\frac{1}{2\sqrt{2}\pi}\int_{-\infty}^{\infty}\Bigg(\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\Bigg)\rho h^{(\varepsilon)}(\rho)\,d\rho,$
-such that
+with the property:
 $Q_{t}^{(\varepsilon)}\psi_{j}=h_{t}^{(\varepsilon)}(\rho_{j})\psi_{j}.$
-Lemma 5.3 allows us to rewrite $k_{t}^{(\varepsilon)}$ in a more tractable form so that
-$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
-where
+Lemma 5.3 allows us to rewrite $k_{t}^{(\varepsilon)}$ in a more tractable form:
+$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv.$
+Here
 $\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}$
-denotes the standard Gaussian mollifier. Since an operator on $X$ is uniquely determined by its action on the orthonormal basis, it follows that $Q_{t}^{(\varepsilon)}=P_{t}^{(\varepsilon)}$.
-Let $\tilde{P}_{t}$ denote an integral operator on $X$ with kernel $K_{t}^{\Gamma}$.
-To show that $\tilde{P}_{t}$ is well-defined and bounded, we use the Schur test (see [HS12]), so it suffices to verify that for all $x\in X$
-$\int_{X}K_{t}(x,y)\,dy$
-is finite as $K_{t}^{\Gamma}(x,y)$ is symmetric in $x$ and $y$. Open the previous expression using the definition of $K_{t}$:
-$\frac{1}{2\sqrt{2}\pi}\,\int_{X}\sum_{\gamma\in\Gamma}\frac{\bm{1}_{t>d(x,\gamma y)}}{\sqrt{\cosh(t)-\cosh(d(x,\gamma y))}}dy.$
-Uniting the integral and the sum into an integral over the universal cover gives the following equal expression:
-$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}dy.$
-Writing this using polar coordinates around $x$ yields
+is the standard Gaussian mollifier. Since a bounded operator is uniquely determined by its action on an orthonormal basis, it follows that $Q_{t}^{(\varepsilon)}=P_{t}^{(\varepsilon)}$.
+Let $\tilde{P}_{t}$ denote the integral operator on $X$ with kernel $K_{t}^{\Gamma}$.
+To show that $\tilde{P}_{t}$ is well-defined and bounded, we use the Schur test (see [HS78]).
+It suffices to verify that for all $x\in X$,
+$\int_{X}K_{t}^{\Gamma}(x,y)\,dy<\infty,$
+since $K_{t}^{\Gamma}(x,y)$ is symmetric in $x$ and $y$. We open the integral using the definition of $K_{t}^{\Gamma}$:
+$\frac{1}{2\sqrt{2}\pi}\,\int_{X}\sum_{\gamma\in\Gamma}\frac{\mathbf{1}_{t>d(x,\gamma y)}}{\sqrt{\cosh(t)-\cosh(d(x,\gamma y))}}dy.$
+Uniting the integral and the sum into an integral over the universal cover gives:
+$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}dy.$
+Writing this using polar coordinates around $x$ yields:
 $\frac{2\pi}{2\sqrt{2}\pi}\int_{0}^{t}\frac{\sinh(\rho)}{\sqrt{\cosh(t)-\cosh(\rho)}}d\rho=\frac{4\pi}{2\sqrt{2}\pi}\sqrt{\cosh(t)-1},$
-where coefficient $2\pi$ appeared first from the polar coordinates since there was no dependency on the direction, and the equality followed from computing the integral.
-The result is finite and $\tilde{P}$ is well-defined and bounded.
-To show that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, as claimed in the statement of the proposition, it suffices to verify that for every $f,g\in C^{\infty}(X)$
-$\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle=0\iff\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
-since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ where $X$ is compact. To show this, start by considering the absolute value term of the previous expression. It can be written as
-$\left|\int_{X}g(x)^{*}\big(P_{t}f(x)-\tilde{P}_{t}f(x)\big)dx\right|,$
-where we opened the inner product and united the two integrals.
-By pulling the absolute values inside the integral and then bounding $|g|$ with the $\sup$ norm, we get the following upper bound for the previous expression:
+where the coefficient $2\pi$ arises from the polar coordinates, and the equality follows from computing the integral. The result is finite, and $\tilde{P}_{t}$ is well-defined and bounded.
+To show that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, as claimed in the proposition, it suffices to verify that for every $f,g\in C^{\infty}(X)$,
+$\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
+since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ where $X$ is compact. To show this, consider the absolute value:
+$\left|\int_{X}g(x)^{*}\big(P_{t}f(x)-\tilde{P}_{t}f(x)\big)dx\right|.$
+By moving the absolute value inside the integral and bounding $|g|$ by its supremum norm, we obtain the upper bound:
 $\|g\|_{\infty}\int_{X}\left|P_{t}f(x)-\tilde{P}_{t}f(x)\right|dx.$
-Focus on the absolute value integrand appearing in the previous expression.
-We can write it in the following form by writing $P_{t}f$ using the limit and then by writing the operator action using the kernels:
+We focus on the absolute value inside the previous integral. Writing $P_{t}f$ using the limit and expressing the operator action via kernels, we obtain
 $\left|\lim_{\varepsilon\to 0}\int_{X}K_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}^{\Gamma}(x,y)f(y)\,dy\right|.$
-Writing $K_{t}^{\Gamma}$ and $K_{t}^{(\varepsilon)}$ using their sum representations over $\Gamma$ and then uniting the integral and the sum into an integral over the universal cover yields the following equal expression, where we also view $f$ as $\Gamma$ periodic function:
+Expressing the kernels using the sum over $\Gamma$ and combining the sum with the integral over $X$ into an integral over the hyperbolic plane yields the following equivalent expression:
 $\left|\lim_{\varepsilon\to 0}\int_{\mathbb{H}}k_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}(x,y)f(y)\,dy\right|.$
-By pulling the absolute values inside the integral and bouding $|f|$ with the $\sup$ norm, we get the following upper bound:
-$\|f\|_{\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|k_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}(x,y)\right|dy.$
-Focus on the integral in the previous expression. By opening $k_{t}$ and $K_{t}$, the integral can be written as
-$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\left|\left(\int_{\mathbb{R}}\frac{\left(\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)\right)\,\bm{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}dv\right)-\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\right|dy.$
-By defining
-$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
-we may write the preceding expression as
+Here, $f$ is viewed as a $\Gamma$-periodic function on $\mathbb{H}$. By moving the absolute value inside and bounding $|f|$ by its supremum norm, we obtain the upper bound:
+$\|f\|_{\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|k_{t}^{(\varepsilon)}(x,y)-K_{t}(x,y)\right|dy.$
+Focus on the limit of the integral. By opening $k_{t}$ and $K_{t}$, the limit can be written as
+$\lim_{\varepsilon\to 0}\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\left|\left(\int_{\mathbb{R}}\frac{(\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v))\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}dv\right)-\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\right|dy.$
+Define
+$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}.$
+Using this, the previous integral expression becomes
 $\frac{1}{2\sqrt{2}\pi}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(t)+\chi_{d(x,y)}(-t)\right|dy,$
-where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$.
-The additional term $\chi_{d(x,y)}(-t)$ equals 0 so adding it had no effect on the equality.
-Using the triangle inequality, the previous expression is bounded by
-$\frac{1}{2\sqrt{\pi}}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|+\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(-t)\right|\,dy.$
-Finally, Lemma 5.4 guarantees that $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$.
-The line of reasoning leading to this cocnlusion means that
-$\langle P_{t}f,g\rangle=\langle\tilde{P}_{t}f,g\rangle,$
+where $\ast_{\mathbb{R}}$ denotes convolution on $\mathbb{R}$. The term $\chi_{d(x,y)}(-t)$ equals 0, so its addition has no effect. Using the triangle inequality, the expression is bounded by
+$\frac{1}{2\sqrt{2}\pi}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|+\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(-t)\right|dy.$
+Finally, Lemma 5.4 guarantees that $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. Hence, the limit equals 0, and
+$\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
 thus concluding the proof. ∎
-The remaining section contains the lemmas and propositions used in the proof of Proposition 5.1.
-First, we record a slightly modified formulation of Proposition 5 in [Mar04].
-It gives a family of operators with explicit formulae for their integral kernels on the compact connected hyperbolic surface.
-Especially, these operators share the eigenfunctions with the Laplacian of the surface and of which eigenvalues have a representation using the eigenvalues of the Laplacian.
+The remaining part of this section contains lemmas and propositions used in the proof of Proposition 5.1.
+First, we record a slightly modified formulation of [MAR12, Prop.
+5].
+It gives a family of operators with explicit formulae for their integral kernels on a compact connected hyperbolic surface.
+These operators share the eigenfunctions with the Laplace–Beltrami operator of the surface, and their eigenvalues have a representation in terms of the Laplace–Beltrami eigenvalues.
 Proposition 5.2.
-Let $h:\{x+iy:|y|<C\}\to\mathbb{C}$ satisfy:
-  1. i.
+Let $h:\{x+iy\in\mathbb{C}\ :\ |y|<C\}\to\mathbb{C}$ satisfy the following properties, where $C$ is a positive constant:
   • $h$ is analytic,
-  2. ii.
   • $h$ is even,
-  3. iii.
-    $h(x+iy)$ decays superpolynomially as $|x|\to\infty$,
+  • $h(x+iy)$ decays superpolynomially as $|x|\to\infty$.
 […]
@@
 $K(x,y)=-\frac{1}{\pi i}\frac{1}{2\sqrt{2}\pi}\int_{-\infty}^{\infty}\left(\int_{d(x,y)}^{\infty}\frac{e^{-i\rho v}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv\right)\rho h(\rho)\,d\rho,$
-satisfies property
+satisfies
 $P\psi_{j}=h(\rho_{j})\psi_{j},$
-where $\psi_{j}$ is a Laplacian eigenfunction on $X$ with eigenvalue $\lambda_{j}$ and $\rho_{j}=\sqrt{\lambda_{j}-1/4}$.
-Next, we derive a convenient expression for $k_{t}^{(\varepsilon)}$, an integral kernel that is crucial in establishing a convenient formula for the integral operator $P_{t}$.
+where $\psi_{j}$ is a Laplacian eigenfunction on $X$ with eigenvalue $\lambda_{j}$ and $\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}$.
+In the next lemma, we derive a convenient expression for $k_{t}^{(\varepsilon)}$, which is an integral kernel crucial in establishing a tractable formula for the kernel of the integral operator $P_{t}$.
+More specifically, $k_{t}^{(\varepsilon)}$ is used to approximate the integral kernel of $P_{t}$.
+We state the lemma.
 Lemma 5.3.
@@
 is the standard Gaussian mollifier.
-The proof is based on noting that there is a Fourier transform inside the expression used to define $k_{t}^{(\varepsilon)}$ in the statement.
-Then by using the inverse Fourier transform of $\phi_{\varepsilon}$ and the Fourier transform of $s_{t}(x)=x^{-1}\,\sin(t\,x)$ (in the distributional sense) together with the Fourier convolution identity, we acquire the wanted conclusion.
+The proof is based on noting that there is a Fourier transform inside the expression used to define $k_{t}^{(\varepsilon)}$.
+By using the inverse Fourier transform of $\phi_{\varepsilon}$ and the Fourier transform of $s_{t}(x)=x^{-1}\,\sin(tx)$ (in the distributional sense) together with the Fourier convolution identity, we obtain the desired conclusion.
 Proof of Lemma 5.3.
-Applying Fubini’s theorem to interchange the order of integration, we obtain that $k_{t}^{(\varepsilon)}$ equals
+By applying Fubini’s theorem to interchange the order of integration, we obtain that $k_{t}^{(\varepsilon)}$ equals
 $\frac{1}{\pi i}\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{1}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\left(\int_{-\infty}^{\infty}e^{-i\rho v}\rho h_{t}^{(\varepsilon)}(\rho)\,d\rho\right)dv.$ (5.1)
@@
 $\mathcal{F}(f)(x)=\int_{\mathbb{R}}f(y)e^{-ixy}\,dy$
-and hence the Inverse Fourier transform is
-$F^{-1}f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}f(y)\,e^{ixy}dy.$
+and the inverse Fourier transform is
+$\mathcal{F}^{-1}f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}f(y)\,e^{ixy}dy.$
 We observe that
@@
 $\mathcal{F}^{-1}(\phi_{\varepsilon})(x)=\frac{1}{2\pi}\,e^{-\frac{\varepsilon^{2}x^{2}}{2}}.$
-By using the Fourier convolution identity and the Fourier transform of sine we observe that the inner integral of Expression (5.1) equals
+The inner integral in Expression (5.1) equals
 $\mathcal{F}\left(\rho\mapsto 2\pi\,\sin(t\rho)\,\mathcal{F}^{-1}(\phi_{\varepsilon})\right)(v)=\frac{2\pi}{2\pi}\left(\mathcal{F}(s_{t})\ast\phi_{\varepsilon}\right)(v)=\pi i\left(\phi_{\varepsilon}(v+t)-\phi_{\varepsilon}(v-t)\right),$
-where $s_{t}$ denotes function $x\mapsto\sin(tx)$. Substituting this into Expression (5.1) yields the following equal expression:
-$\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(v-t)-\phi_{\varepsilon}(v+t)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(t+v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
-where the equality followed using the symmetry of the Gaussian mollifier.
-Finally, the following lemma establishes convergence of $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$.
-Interestingly, the convolution here is over the real line but the convergence is studied in the $\mathbb{H}$.
+where the first equality follows from the convolution identity, and the second equality uses
+$\mathcal{F}(s_{t})(x)=\pi i\bigl(\delta(x+t)-\delta(x-t)\bigr),$
+with $s_{t}(x)=\sin(tx)$ and $\delta$ denoting the Dirac distribution.
+Substituting this into Expression (5.1) gives:
+$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(v-t)-\phi_{\varepsilon}(v+t)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(t+v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
+where the equality follows from the symmetry of the Gaussian mollifier. ∎
+Finally, the following lemma establishes the convergence of $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$.
+Interestingly, the convolution is over the real line, but the convergence is studied in the hyperbolic plane.
 Lemma 5.4.
@@
 $\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}$
-and let
+and
 $\phi_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi\varepsilon^{2}}}e^{-\frac{x^{2}}{2\varepsilon^{2}}}.$
-Then map $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$ for every $t\in\mathbb{R}$ as $\varepsilon\to 0$, where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$.
-In the proof we show that
+Then the map $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$ for every $t\in\mathbb{R}$ as $\varepsilon\to 0$, where $\ast_{\mathbb{R}}$ denotes convolution on $\mathbb{R}$.
+In the proof, we show that
 $\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy\xrightarrow{\varepsilon\to 0}0.$
-The idea is to present this integral as a convolution $\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)$, where $A$ is a continuous function with $A(t)=0$.
-Then the conclusion can be drawn using the properties of the convolution.
+The idea is to present this integral as a convolution $\phi_{\varepsilon}\ast_{\mathbb{R}}A_{t}$, where $A_{t}$ is a continuous function with $A_{t}(t)=0$.
+Then the conclusion follows from the properties of convolution.
 Proof of Lemma 5.4.
-We want to show that the following expression goes to 0 as $\varepsilon$ goes to zero:
-$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy=\int_{\mathbb{H}}\left|\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\big(\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\big)\,dv\right|dy.$
-By the triangle inequality and Fubini’s theorem, this is bounded by
-$\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\int_{\mathbb{H}}\left|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\right|\,dy\,dv=\phi_{\varepsilon}\ast_{\mathbb{R}}A(t),$
+We want to show that the following expression goes to 0 as $\varepsilon\to 0$:
+$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy=\int_{\mathbb{H}}\left|\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\big(\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\big)\,dv\right|dy.$ (5.2)
+Here, we used that the integral over the Gaussian mollifier equals $1$:
+$\chi_{d(x,y)}(t)=\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\,\chi_{d(x,y)}(t)\,dv.$
+By the triangle inequality and Fubini’s theorem, Expression (5.2) is bounded by
+$\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\int_{\mathbb{H}}\left|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\right|\,dy\,dv=\phi_{\varepsilon}\ast_{\mathbb{R}}A_{t}(t),$
 where
-$A(v)=\int_{\mathbb{H}}|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|\,dy.$
-By using the polar coordinates around $x$ together with the triangle inequality we have that
-$A(v)\leq\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\,dy+\int_{\mathbb{H}}\frac{\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dy<\infty.$
-Hence we observe that $v\mapsto|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|$ has a dominating function with which the dominated convergence theorem can be used to observe that $A(v)\to 0$ as $v\to t$.
-$A$ is continuous, and as it is continuous, we have that
-$\lim_{\varepsilon\to 0}\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)=0,$
+$A_{t}(v)=\int_{\mathbb{H}}|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|\,dy.$
+We see that $A_{t}(t)=0$, and we want to show that $\lim_{v\to t}A_{t}(v)=0$. By the triangle inequality:
+$A_{t}(v)\leq\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\,dy+\int_{\mathbb{H}}\frac{\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dy<\infty.$
+The finiteness can be verified using polar coordinates around $x$.
+Hence, $y\mapsto|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|$ has a dominating function when $v$ is in a compact interval around $t$.
+By the dominated convergence theorem, $A_{t}(v)\to 0$ as $v\to t$.
+Since $A_{t}(t)=0$ and $A_{t}$ is continuous at $t$, the properties of convolution yield
+$\lim_{\varepsilon\to 0}\phi_{\varepsilon}\ast_{\mathbb{R}}A_{t}(t)=0,$
 establishing the claim. ∎
@@
 $\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2},$
-which is the geometric data appearing in the proof of Theorem1.5 in Section 8. $P_{t}$ and $P_{t,\tau}$ are the integral operators we extensively studied in the previous section, Section 5.
+which is the geometric data appearing in the proof of Theorem 1.5 in Section 8.
+$P_{t}$ and $P_{t,\tau}$ are the integral operators we extensively studied in the previous section, Section 5.
+We now state the proposition.
 Proposition 6.1 (Geometric data).
-Let $P_{t}$ and $P_{t,\tau}$ be integral operators on a compact connected hyperbolic surface $X$ with Laplace-Beltrami operator $\Delta_{X}$ such that $P_{t}$ has integral kernel
+Let $P_{t}$ and $P_{t,\tau}$ be integral operators on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$ with Laplace–Beltrami operator $\Delta_{X}$ such that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, defined as
 $K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\quad K_{t}(x,y)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
-and $P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)\,K_{t}^{\Gamma}$. Then
-$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D_{1}\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)$
-where $\beta=\beta(\lambda_{1}(X))$, where $\lambda_{1}(X)$ is the spectral gap of $\Delta_{X}$, and where
+and $P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)\,K_{t}^{\Gamma}$. Let $a$ be a function on $X$ with mean zero. Then
+$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)$
+where $\beta=\beta(\lambda_{1}(X))$, $|X\setminus X(4T)|$ is the volume of the points of $X$ having injectivity radius less than $4T$, $\lambda_{1}(X)$ is the spectral gap of $\Delta_{X}$, and where
 $\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
-The constant appearing is
-$C_{7}=\frac{7200}{\pi}.$
-This is the most technical part of the whole paper.
-The core idea is first to write $K_{t}^{\Gamma}$ and $K_{t,\tau}^{\Gamma}$ using $K_{t}$ while simultaneously simplifying the expressions by unifying the sums over the Fuchsian group $\Gamma$ with the integrals over the fundamental domains of the surface $X$ to get integrals over the whole hyperbolic plane.
-This careful splitting allows us to separate a certain weight function $F_{t,t^{\prime},\rho}$ into a separate integral in which it is easier to bound.
-Section 7 is dedicated for the analysis of $F_{t,t^{\prime},\rho}$.
-Finally, we split the expression into two sides, one with the points with large injectivity radius and the other with the points with small injectivity radius.
-Both cases require careful analysis to acquire reasonable bounds, especially the side with the large injectivity radius for which we need the exponential mixing theorem, Theorem 1.2.
-The proof relies on multiple auxiliary lemmas presented after the proof. This approach makes the proof clearer.
+The constant appearing is $C_{7}=80000$.
+By careful bounding and rewriting the Hilbert-Schmidt norm expression, we get a tidy expression in which the observable $a$ is separated from the geometric data arising from the choice of the propagator; this geometric data is contained in a certain weight function $F_{t,t^{\prime},\rho}$.
+This weight is studied in Section 7.
+The bound is acquired by separating an integration over the fundamental domain into a part dealing with the points with large injectivity radius and a part dealing with the points with small injectivity radius.
+In the large injectivity radius case, the exponential mixing theorem, Theorem 1.2, gives a crucial decay needed to acquire the bounds.
+The final bounds are obtained by using bounds for $F_{t,t^{\prime},\rho}$.
 Proof of Proposition 6.1.
@@
 $\int_{D}\int_{D}\left|\frac{1}{T}\int_{0}^{T}\int_{D}K_{t,\tau}^{\Gamma}a(z)K_{t}^{\Gamma}(z,y)dt\right|^{2}dydx.$
-Using that $K^{\Gamma}_{t,\tau}=\cos(t\tau)K_{t}^{\Gamma}$ and by expanding the squares, we can write the previous expression as follows:
+We also used that $P_{t,\tau}$ is self-adjoint. Using that $K^{\Gamma}_{t,\tau}=\cos(t\tau)K_{t}^{\Gamma}$ and by expanding the squares, we can write the previous expression as follows:
 $\int_{D\times D}\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D\times D}K_{t}^{\Gamma}(x,z)a(z)K_{t}^{\Gamma}(z,y)\,K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dz^{\prime}dzdt^{\prime}dtdydx.$
-Using Fubini’s theorem to reorder the integrals yields the following equals expression:
+Using Fubini’s theorem to reorder the integrals yields the following equal expression:
 $\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D\times D\times D\times D}K_{t}^{\Gamma}(x,z)a(z)K_{t}^{\Gamma}(z,y)K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dxdydz^{\prime}dzdt^{\prime}dt.$
@@
 $\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}\int_{D}\int_{D}K_{t}^{\Gamma}(x,z)a(z)K^{\Gamma}_{t}(z,y)\,K_{t^{\prime}}^{\Gamma}(x,z^{\prime})a(z^{\prime})^{*}K_{t^{\prime}}^{\Gamma}(z^{\prime},y)dydxdz^{\prime}dz\right|dt^{\prime}dt.$
-This can be written in a more compact form when gathering the factors containing $x$ into one integral and the factors containing $y$ into another integral and noticing that these two integrals are the same thus resulting into an integral squared:
+The cosines were bounded from above by $1$ in the previous step.
+The previous expression can be written in a more compact form when gathering the factors containing $x$ into one integral and the factors containing $y$ into another integral and noticing that these two integrals are the same, thus resulting in an integral squared:
 $\frac{1}{T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{D}K_{t}^{\Gamma}(x,z)K_{t^{\prime}}^{\Gamma}(x,z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt.$
 Lemma 6.2 allows us to write the previous expression using the hyperbolic plane kernels $K_{t}$ and $K_{t^{\prime}}$ instead of the surface kernels $K_{t}^{\Gamma}$ and $K_{t^{\prime}}^{\Gamma}$:
-$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K_{t}(x,z)K_{t^{\prime}}(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt$
-Coefficient $(8\,\pi^{2})^{-1}$ came from the definition of $K_{t}^{\Gamma}$:
+$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K_{t}(x,z)K_{t^{\prime}}(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt,$
+where the coefficient $(8\,\pi^{2})^{-1}$ comes from the definition of $K_{t}^{\Gamma}$:
 $K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y).$
-Lemma 6.3 allows us to rewrite the integral inside the absolute value of the previous integral expression in a very convenient form, eliminating but one sum. Thus the previous expression equals:
-$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{\mathbb{H}}a(z)a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K_{t}(x,z)K_{t^{\prime}}(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K_{t}(y,\gamma z)K_{t^{\prime}}(y,z^{\prime})dy\right)dz^{\prime}dz\right|dt^{\prime}dt$
-which further equals
-$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{\mathbb{H}}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt,$ (6.1)
-when $F_{t,t^{\prime},\rho}$ is defined to equal
+Lemma 6.3 allows us to rewrite the integral inside the absolute value in a convenient form, eliminating all but one sum. Thus the previous expression equals:
+$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{\mathbb{H}}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt,$ (6.1)
+where $F_{t,t^{\prime},\rho}$ is defined as
 $\int_{\mathbb{H}}K_{t}(x,w)K_{t^{\prime}}(x,w^{\prime})dx,$
-where $w,w^{\prime}\in\mathbb{H}$ are any points seprated by distance $\rho$.
-In Section 7, where we study $F_{t,t^{\prime},\rho}$ carefully, in Lemma 7.1 we will see that the chosen points $w,w^{\prime}$ do not matter, only their distance.
-The definition of $F_{t,t^{\prime},d(z,z^{\prime})}$ shows that $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, which follows from the definition of $K_{t}$ and $K_{t^{\prime}}$.
-Using this observation gives us the following equal expression for Expression (6.1):
-$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (6.2)
-Focus on the absolute values in the previous expression. They are bounded from above by
-$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
-$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$ (6.3)
-where the integration over $D$ is split into two: the points having injectivity radius bigger than $4T$, $D(4T)$, and the points having injecitivity radius smaller than $4T$, $D\setminus D(4T)$.
-The large injectivity radius case and the small injectivity radius case are bounded separately.
-The small injectivity radius term is bounded in Lemma 6.4. Using its bound gives the following upper bound for the previous expression:
-$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
-$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|$
-$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
-In the previous, the second term is bounded by the third term as the second term has no sum and the third one does. Hence the previous expression is bounded by
-$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
-+ $+\ 2\,\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$
-where the second term is the small injectivity radius case, which is bounded in Lemma 6.5. Using this lemma gives the following upper bound for the previous expression:
-$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ 8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$ (6.4)
-The latter term is bounded by
-$8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}11e^{\beta}\sinh(\rho)(1+\rho)e^{-\beta\rho}e^{2\beta T}F_{t,t^{\prime},\rho}^{2}d\rho,$
-since
-$11\,e^{\beta}(1+\rho)e^{-\beta\rho}e^{\beta\,2T}\geq 1.$
-Hence Expression (6.4) is bounded by:
-$11e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
-We have now acquired good bounds for the absolute values appearing in Expression (6.1), and this expression is thus bounded by
-$\frac{11}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
-As $F_{t,t^{\prime},\rho}$ equals $F_{t^{\prime},t,\rho}$ and since we integrate from 0 to $T$ both $t$ and $t^{\prime}$, we can use symmetry to integrate the inner integral of $t^{\prime}$ from 0 to only $t$ twice.
-Hence we write the previous expression as
-$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$ (6.5)
-The innermost integral can be bounded using Proposition 6.6, which uses the bound acquired for $F_{t,t^{\prime},\rho}$ in Section 7. We get the following upper bound for the previous expression:
-$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}dt^{\prime}dt,$
-where
-$C_{6}=480\pi.$
-The remaining two integrals have
-$\frac{2T\,C_{6}}{\beta^{3}}$
-as an upper bound, which is seen after evaluating these integrals. Then the previous integral expression has the following upper bound:
-$\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
-where
-$\frac{22\cdot 2\cdot C_{6}\cdot e}{8\pi^{2}}\leq\frac{7200}{\pi}=C_{7},$
-where we used that $e^{\beta}$ is less than $e$. This is also an upper bound for the Hilbert-Schmidt norm we set out to bound. It is of the wanted form which concludes the proof.
-The remaining of this section contains the lemmas needed to prove Proposition 6.1 together with the contents of Section 7.
+with $w,w^{\prime}\in\mathbb{H}$ separated by distance $\rho$; Lemma 7.1 ensures that $F_{t,t^{\prime},\rho}$ is well-defined.
+The definition of $F_{t,t^{\prime},d(z,z^{\prime})}$ shows that $F_{t,t^{\prime},d(z,z^{\prime})}=0$ if $d(z,z^{\prime})>2T$, which follows from the definitions of $K_{t}$ and $K_{t^{\prime}}$.
+Using this, Expression (6.1) can be rewritten by restricting the integration domain of $z^{\prime}$ as follows:
+$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (6.2)
+We split the integral over $D$ into points with injectivity radius greater than $4T$, $D(4T)$, and points with injectivity radius less than $4T$, $D\setminus D(4T)$:
+$\left|\int_{D}\cdots\right|\leq\left|\int_{D(4T)}\cdots\right|+\left|\int_{D\setminus D(4T)}\cdots\right|.$
+By moving the absolute value inside the latter term and bounding $|a|$ by its supremum norm, the previous expression yields the following upper bound:
+$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
+[…]
+where the large and small injectivity radius terms are bounded separately.
+The small injectivity radius term is bounded in Lemma 6.4. Using this estimate, we obtain the following upper bound:
+$30\|a\|_{2}^{2}\,\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho+2\,\|a\|_{\infty}^{2}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$
+The second term of the previous expression is bounded in Lemma 6.5, yielding the following upper bound:
+$30\|a\|_{2}^{2}\,\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ 8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
+This is further bounded by
+$8\pi^{2}\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho,$
+using that $8\pi^{2}>30$ and $e^{2T}\,e^{-\beta\rho}\,(1+\rho)\geq 1$.
+Hence, Expression (6.1) is bounded by
+$\frac{1}{T^{2}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
+By symmetry of $F_{t,t^{\prime},\rho}$ in $t$ and $t^{\prime}$, we can restrict the inner integral to $0\leq t^{\prime}\leq t$ and multiply by $2$:
+$\frac{2}{T^{2}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
+Applying Proposition 6.6 to the innermost integral, we obtain the upper bound
+$\frac{2}{T^{2}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{4}(t-t^{\prime})}\,dt^{\prime}\,dt,$
+where $C_{6}=10000$. Integrating in $t^{\prime}$ and $t$ gives an upper bound of $4TC_{6}/\beta^{3}$, so the entire expression is bounded by
+$\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
+with $2\cdot 4\cdot C_{6}=80000=C_{7}$.
+The remaining part of this section contains the lemmas needed to prove Proposition 6.1, together with the contents of Section 7.
 The following lemma allows us to write the integration containing the kernels living on the hyperbolic surface using the kernels living on the hyperbolic plane.
@@
 be integral kernels on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, where $K$ and $L$ are integral kernels on $\mathbb{H}$ that depend only on distance $d(x,y)$. Then
-$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma_{2}z^{\prime})dx,$
+$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx,$
 where $D$ is a fundamental domain of $X$.
-The proof is on uniting the sum over the Fuchsian group and the integration over the fundamental domain into an integration over the hyperbolic plane.
+The proof is based on uniting the sum over the group $\Gamma$ and the integration over the fundamental domain $D$ into an integration over the hyperbolic plane.
 Proof of Lemma 6.2.
@@
 $\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(x,\gamma_{1}z)L(x,\gamma_{2}z^{\prime})\,dx$
-by opening the kernels using their sum form. Doing a change of variables $\gamma_{2}\mapsto\gamma_{1}\gamma_{2}$ for the sum allows us to write the previous expression as
+by opening the kernels using their sum forms. Doing a change of variables $\gamma_{2}\mapsto\gamma_{1}\gamma_{2}$ for the inner sum allows us to write the previous expression as
 $\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(x,\gamma_{1}z)L(x,\gamma_{1}\gamma_{2}z^{\prime})\,dx.$
@@
 $\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}K(\gamma_{1}^{-1}x,z)L(\gamma_{1}^{-1}x,\gamma_{2}z^{\prime})dx.$
-Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover gives us the following equal expression:
+Uniting the integral and the sum over $\gamma_{1}$ into the integral over the universal cover gives us the following equal expression:
 $\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})\,dx,$
-where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. This concludes the proof.
-The following lemma allows us to simplify the notation by removing one sum over the Fuchsian group.
+where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. This concludes the proof. ∎
+The following lemma allows us to simplify an integral expression in which sums over the group $\Gamma$ appear.
 Lemma 6.3.
-Let $K$ and $L$ be integral kernels on compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$. Let $a$ be a function on $X$. Then:
-$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}$
-$=\qquad$ $\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime},$
+Let
+$K^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}K(x,\gamma y),\qquad L^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}L(x,\gamma y)$
+be integral kernels on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, where $K$ and $L$ are integral kernels on $\mathbb{H}$ that depend only on distance $d(x,y)$.
+Let $a$ be a function on $X$.
+Then
+$\int_{D}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}$
+$=\ \int_{\mathbb{H}}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime},$
 where $D$ is a fundamental domain of $X$.
-The proof is based on the same principle as in the proof of Lemma 6.2: we unite the sum over the Fuchsian group and the integral over the fundamental domain into an integration over the hyperbolic plane.
+The proof is based on the same principle as the proof of Lemma 6.2: we unite the sum over the group and the integral over the fundamental domain into an integration over the hyperbolic plane.
 Proof of Lemma 6.3.
 Focus on the following expression:
-$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}.$
+$\int_{D}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}.$
 We expand the squares and reorder the integrals and the sums using Fubini’s theorem to get the following expression:
-$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}z^{\prime})dy\right)dz^{\prime}.$
-Doing a change of variables $\gamma_{2}\mapsto\gamma_{2}^{-1}\gamma_{1}$ in the sum allows us to write the previous expression as
-$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}^{-1}\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
+$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}z^{\prime})dy\right)dz^{\prime}.$
+Doing a change of variables $\gamma_{2}\mapsto\gamma_{2}^{-1}\gamma_{1}$ in the inner sum allows us to write the previous expression as
+$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}^{-1}\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
 which further equals
-$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma_{2}y,\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
+$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma_{2}y,\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
 when using that $L(x,\gamma y)=L(\gamma^{-1}x,y)$.
-Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover is possible since $a^{*}$ can be seen as $\Gamma$-periodic function on $\mathbb{H}$.
+Uniting the integral and the sum over $\gamma_{1}$ into the integral over the universal cover is possible since $a^{*}$ can be seen as a $\Gamma$-periodic function on $\mathbb{H}$.
 This gives the following equal expression:
-$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy\right)dz^{\prime},$ (6.6)
+$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy\right)dz^{\prime},$ (6.4)
 where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. For the latter integral we have that
@@
 With this observation, we can rewrite Expression (6.4) as
-$\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime}.$
-The following lemma deals with the integration over the part of the fundamental domain with large injectivity radius, giving us the upper bound consisting of two terms: one is an integral with integrand with exponential decay and the other is an integration over the points with the small injectivity radius.
-The first term is desired and can be united with the bound acquired from the small injectivity radius case, but the second term needs to be further bounded.
-This is taken care of when we deal with all the parts dealing with the parts with small injectivity radius, which is done in Lemma 6.5.
+$\int_{\mathbb{H}}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime}.$
+The following lemma addresses integration over the part of the fundamental domain with large injectivity radius, yielding an upper bound consisting of two terms.
+The first term is an integral whose integrand exhibits exponential decay, while the second term is an integral over points with small injectivity radius.
+The latter term is bounded in Lemma 6.5.
 Lemma 6.4.
 Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$.
-$D(4T)$ denotes the points with injectivity radius bigger than $4T$.
+Let $D(4T)$ denote the points with injectivity radius larger than $4T$.
 Let $a$ be a mean zero function on $X$.
 Then
-$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right|$
-$=\ \|a\|_{2}^{2}\,\int_{0}^{2T}\,11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}\,(1+\rho)\,e^{-\beta\rho}\,d\rho$
-$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|,$ (6.7)
-where $\beta=\beta(\lambda_{1}(X))$, $\lambda_{1}(X)$ is the spectral gap of Laplace-Beltrami operator $\Delta_{X}$, and where
-$\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
-The proof is based on modifying the expression so that the exponential mixing theorem, Theorem 1.2, can be applied.
+$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
+= $=\ 30\|a\|_{2}^{2}\,\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}\,d\rho+\|a\|_{\infty}^{2}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz,$ (6.5)
+where $\beta=\beta(\lambda_{1}(X))$, $\lambda_{1}(X)$ is the spectral gap of the Laplace–Beltrami operator $\Delta_{X}$, and
+$\beta(x)=\begin{cases}1-\sqrt{1-4x},&\text{ if }x\leq\frac{1}{4},\\ 1,&\text{ if }x>\frac{1}{4}.\end{cases}$
+The proof is based on noticing that only the identity element can have a non-zero contribution.
+After this, we separate the integration over the whole fundamental domain and the integration over points with small injectivity radius using the triangle inequality.
+The crucial part is bounding the first term using the exponential mixing theorem, Theorem 1.2.
 Proof of Lemma 6.4.
-As $z\in D(4T)$, we have that $d(z,\gamma z)$ is larger than $4T$ for any non-identity group element $\gamma$.
-We also have that $z^{\prime}\in B(z,2T)$.
-This means that $d(\gamma z,z^{\prime})$ is larger than $2T$ for any non-identity group element.
-As $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, we have that only the identity element can have a non-zero contribution.
-Expression (6.4) equals
-$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.8)
-By adding and subtracting
-$\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz$
-in the absolute values of Expression (6.8) and then appyling the triangle inequality gives the following expression as an upper bound for Expression (6.8):
-$\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.9)
-Focus on the first term. We write $z^{\prime}$ using polar coordinates around $z$, so $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, and the first term equals
-$\left|\int_{D}a(z)\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)\,a^{*}(\varphi_{\rho}(z,\theta))\,F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz\right|,$
-where $a$ can be seen as a function on the unit tangent bundle by defining that $a(z,\theta)=a(z)$.
-The integrals of this expression can be reordered using Fubini’s theorem yielding the following expression:
-$\left|\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\int_{D}\int_{\mathbb{S}^{1}}a(z)\,a^{*}(\varphi_{\rho}(z,\theta))\,d\theta\,dz\,d\rho\right|.$
-By pulling the absolute values inside the first integral, we have the following upper bound, where the inner two integrals are written as an inner product:
+As $z\in D(4T)$, we have $d(z,\gamma z)>4T$ for any non-identity element $\gamma\in\Gamma$.
+Also, $z^{\prime}\in B(z,2T)$, so $d(\gamma z,z^{\prime})>2T$ for any non-identity element.
+Since $F_{t,t^{\prime},d(z,z^{\prime})}=0$ if $d(z,z^{\prime})>2T$, only the identity element contributes.
+Hence the left-hand side of Expression (6.5) equals
+$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.6)
+Adding and subtracting
+$\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz$
+and applying the triangle inequality gives the following upper bound for Expression (6.6):
+$\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.7)
+We focus on the second term of the previous expression.
+By moving the absolute value inside, bounding $|a|$ by its supremum norm, and introducing a sum, we obtain an upper bound corresponding to the second term on the right-hand side of Expression (6.5).
+Focus on the first term of Expression (6.7). Write $z^{\prime}$ in polar coordinates around $z$, so $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, giving
+$\left|\int_{D}a(z)\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)\,a(\varphi_{\rho}(z,\theta))^{*}\,F_{t,t^{\prime},\rho}^{2}\,d\rho\,d\theta\,dz\right|.$
+Reordering the integrals using Fubini’s theorem gives
+$\left|\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\int_{D}\int_{\mathbb{S}^{1}}a(z)\,a(\varphi_{\rho}(z,\theta))^{*}\,d\theta\,dz\,d\rho\right|.$
+Pulling the absolute value inside the first integral, we have
 $\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,\left|\left\langle a,a\circ\varphi_{\rho}\right\rangle_{L^{2}(D\times\mathbb{S}^{1})}\right|d\rho.$
-We can apply the Exponential mixing theorem, Theorem 1.2, to the absolute value of the inner product which then yields the following upper bound for the previous expression:
-$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}(1+\rho)e^{-\beta\rho}\,d\rho.$
-Hence Expression (6.9) and hence the original expression we set out to bound are bound by
-$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$
-The following lemma deals with the integration over the part of the fundamental domain with small injectivity radius giving an upper bound for it.
-This can then be united with the bound acquired for the large injectivity radius case in Lemma 6.4.
+Applying the Exponential Mixing Theorem, Theorem 1.2, to the inner product yields the upper bound
+$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}(1+\rho)e^{-\beta\rho}\,d\rho\leq 30\|a\|_{2}^{2}\int_{0}^{2T}F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}\,d\rho,$
+where the latter inequality followed by using that $\beta\leq 1$ and $11e\leq 30$. This concludes the proof.
+The following lemma bounds the integration over the part of the fundamental domain with small injectivity radius.
 Lemma 6.5.
-Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. $D(4T)$ denotes the points with injectivity radius larger than $4T$. Then
-$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
-is bounded from above by
-$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho$
-The proof is based on careful use of Hölder’s inequality together with a careful management of the sum elements with non-zero contribtution.
+Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$, and let $D(4T)$ denote the points with injectivity radius larger than $4T$. Then
+$\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\leq 4\pi^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
+The proof is based on the careful use of Hölder’s inequality. Of particular importance is understanding the number of group elements with non-zero contribution.
 Proof of Lemma 6.5.
 Consider
-$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
-By pulling the absolute values inside this is bounded by
-$\int_{D\setminus D(4T)}|a(z)|\int_{B(z,2T)}\left|a^{*}(z^{\prime})\right|F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz,$
-which is further bounded by the following expression when using the supremum norm bounds for $|a|$ and $|a^{*}|$:
-$\|a\|_{\infty}^{2}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$
-In the previous expression, the sum elements with a possible non-zero contribution are those $\gamma\in\Gamma$ such that $d(\gamma z,z^{\prime})$ is smaller than $2T$ for some $z^{\prime}\in B(z,2T)$.
-This collection depends on $z$ and we denote it by $\Gamma(z)$.
-We have that
-$\#\Gamma(z)\leq\frac{|B(z,4T)|}{\operatorname{InjRad}_{X}}\leq 2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}},$
-which follows from that the minimum distance of $\gamma z$ and $\gamma^{\prime}z$, where $\gamma\neq\gamma^{\prime}$ is $\operatorname{InjRad}_{X}$, and that $\gamma z$ must lie in $B(z,4T)$ for it to be possible that $d(\gamma z,z^{\prime})$ is smaller than $2T$, since $z^{\prime}\in B(z,2T)$.
-Hence the previous integral expression has the following upper bound when using Hölder’s inequality and $\Gamma(z)$ instead of $\Gamma$:
-$\|a\|_{\infty}^{2}\,\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}\right)^{\frac{1}{2}}\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}\left(\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right)^{2}\right)^{\frac{1}{2}}.$ (6.10)
-Focus on the double integral in the second factor. Using Hölder’s inequality to sum and using the bound observed for $\Gamma(z)$ gives the following upper bound for it:
-$2\pi\,\int_{D\setminus D(4T)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz.$
-By using that $d(\gamma z,z^{\prime})$ equals $d(z,\gamma^{-1}z^{\prime})$ and interchanging the order of the sum and the integral using the linearity of integral, the previous expression equals
-$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,\gamma^{-1}z^{\prime})}^{2}dz^{\prime}dz,$
-which further equals under a change of variables $z^{\prime}\mapsto\gamma z^{\prime}$ the following:
-$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
-Due to the presence of $F_{t,t^{\prime},d(z,z^{\prime})}$, for non-zero contribution, we must have that $d(z,z^{\prime})$ is less than $2T$. This means that the previous expression equals
-$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)\cap B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz,$
-which is then bounded by
-$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
-As the $\gamma$-dependency is now eliminated, the sum contributes the number of elements it sums over as a coefficient, which leads to the following equal expression:
-$4\pi^{2}\,\frac{e^{8T}}{\operatorname{InjRad}_{X}^{2}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
-Hence we get the following bound for Expression (6.10), as now we have a bound for the second factor:
-$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
-By writing $z^{\prime}$ using polar coordinates around $z$, whence $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, we get the following equal expression:
-$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz,$
-which then equals
-$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho,$
-as there was no $\theta$ dependency nor $z$ dependency.
-The following lemma bounds a complicated integral in which $F_{t,t^{\prime},\rho}$ appears. The bounds for $F_{t,t^{\prime},\rho}$ are established in Section 7.
+$\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$ (6.8)
+Only the terms $\gamma\in\Gamma$ with $d(\gamma z,z^{\prime})<2T$ for some $z^{\prime}\in B(z,2T)$ contribute. Denote this collection by $\Gamma(z)$. Then
+$\#\Gamma(z)\leq\frac{|B(z,4T)|}{\operatorname{InjRad}_{X}}\leq 2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}.$
+This is because $\gamma z$ must lie in $B(z,4T)$ for it to be possible that $d(\gamma z,z^{\prime})<2T$, as $z^{\prime}\in B(z,2T)$.
+Restricting the sum over $\Gamma(z)$ in Expression (6.8), and then applying Hölder’s inequality twice, first to the integral and then to the sum, results in the following upper bound:
+$\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}\#\Gamma(z)\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}.$ (6.9)
+Focus on the right factor. It admits the following upper bound after using Fubini’s theorem to reorder the sums and integrals, and then applying the bound on $\Gamma(z)$:
+$\left(\frac{2\pi\,e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}.$
+Changing variables via $z^{\prime}\mapsto\gamma z^{\prime}$ gives the following equal expression:
+$\left(\frac{2\pi\,e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)}F_{t,t^{\prime},d(\gamma z,\gamma z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}.$
+Since $d(\gamma z,\gamma z^{\prime})=d(z,z^{\prime})$ and $F_{t,t^{\prime},\rho}=0$ for $\rho>2T$, the domain of integration in the inner integral can be restricted to $B(\gamma^{-1}z,2T)\cap B(z,2T)\subset B(z,2T)$.
+Hence, we obtain the following upper bound for the previous integral expression:
+$\left(\frac{2\pi\,e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,\gamma z^{\prime})}^{2}\,dz^{\prime}\,dz\right)^{\frac{1}{2}}$
+$\leq\ \left(\frac{4\pi^{2}\,e^{8T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,\gamma z^{\prime})}^{2}\,dz^{\prime}\,dz\right)^{\frac{1}{2}}.$
+The inequality follows because the contents of the sum no longer depends on $\gamma$, so we can use the bound on $\Gamma(z)$.
+Hence, we obtain the following upper bound for Expression (6.9):
+$2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}\,dz^{\prime}\,dz=4\pi^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,d\rho,$
+where the equality follows from writing $z^{\prime}$ in polar coordinates around $z$ and then integrating over $z$. This concludes the proof.
+The following lemma bounds an integral in which $F_{t,t^{\prime},\rho}$ appears.
 Lemma 6.6.
-Let $\beta$ be a positive constant that is smaller than $1$. Then:
-$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}d\rho\leq\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
-where
-$F_{t,t^{\prime},\rho}\leq 32\pi\,\cdot\,\begin{cases}&\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad\text{ if }\rho<t-t^{\prime}\\ &\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},\qquad\text{ if }t-t^{\prime}<\rho<t\\ &\frac{1}{\sqrt{\sinh(\rho)}},\qquad\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
-and where
-$C_{6}=480\pi.$
-The proof uses bounds for $F_{t,t^{\prime},\rho}$.
-Acquiring good enough bounds for the integral to converge requires very careful analysis which is done in Section 7.
-There we acquire three different bounds for $F_{t,t^{\prime},\rho}$ working in separate cases.
-Hence to bound the integral, we split it into three cases and use the bounds for $F_{t,t^{\prime},\rho}$.
-One of the cases is more technical as both the Beta function and the Gamma function appear in the analysis.
+Let $\beta$ be a positive constant smaller than $1$ and let $0<t^{\prime}<t$. Then
+$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}d\rho\leq\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{4}(t-t^{\prime})},$
+when
+$F_{t,t^{\prime},\rho}\leq 50\frac{1}{\sqrt{\sinh(\max(t-t^{\prime},\rho))}}\bm{1}_{\rho\in(0,t+t^{\prime})},$
+where $C_{6}=10000$.
+The proof is based on splitting the integral into two terms according to the upper bound of $F_{t,t^{\prime},\rho}$.
 Proof.
-We have that
-$(1+\rho)\leq\frac{2}{\beta}\,e^{\frac{\beta}{2}-1}\,e^{-\frac{\beta}{2}\rho},$
-which can be verified by studying the derivatives of the functions. Using this observation gives the following upper bound for the integral we are bounding:
-$\frac{1}{\beta}\,\int_{0}^{2T}\sinh(\rho)\,e^{-\frac{\beta}{2}\rho}\,F_{t,t^{\prime},\rho}^{2}d\rho.$
-By splitting the integral and using the bounds for $F_{t,t^{\prime},\rho}$, we have the following upper bound for the integral expression we set out to bound:
-$\frac{32\pi}{\beta}\left(\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{\sqrt{\sinh(\rho)}\,e^{-\frac{\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\ +\ \int_{t-t^{\prime}}^{t}\frac{e^{-\frac{\beta}{2}\rho}}{(1-e^{t-t^{\prime}-\rho})^{\frac{1}{4}}}d\rho\ +\ \int_{t}^{t+t^{\prime}}e^{-\frac{\beta}{2}\rho}d\rho\right).$ (6.11)
-Let us focus on the first term inside the parentheses. It is bounded by
-$\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{e^{\frac{1-\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\leq\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\int_{0}^{t-t^{\prime}}\frac{1}{\sqrt{t-t^{\prime}-\rho}}d\rho,$
-where the inequality followed from using the upper boundary of the integration domain. The previous then equals after evaluating the integral:
-$2\,\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\sqrt{t-t^{\prime}}\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
-The inequality followed from the observation that a positive $x$ divided by $\sinh(x)$ is bounded from above by $1$.
-Next, let us bound the second term inside the parentheses of Expression (6.11). We begin by writing it in a different form using a change of variables $u=t^{\prime}+\rho-t$:
-$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{t^{\prime}}e^{-\frac{\beta}{2}u}\,\left(1-\frac{1}{\cosh(u)}\right)^{-\frac{1}{4}}\,du.$
-We do another change of variables $z=e^{-u}$ to get
-$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,\frac{dz}{z},$
-where especially
-$1-\frac{1}{\cosh(u)}=\frac{(1-z)^{2}}{1+z^{2}}.$
-The previous integral expression equals
-$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}-1}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,dz\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{1}z^{\frac{\beta}{2}-1}\,(1-z)^{-\frac{1}{2}}\,dz,$
-where the inequality followed from that $(1+z^{2})^{\frac{1}{4}}$ is bounded by $2$ and the integration domain is enlargened. The previous expression then equals
-$2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,B\left(\frac{\beta}{2},\frac{1}{2}\right),$ (6.12)
-where $B$ is the Beta function which can be written using Gamma functions (the reference for the Beta and Gamma functions appearing here is [Olv10, Chp.6]):
-$B\left(\frac{\beta}{2},\frac{1}{2}\right)=\frac{\Gamma\left(\frac{1}{2}\right)\,\Gamma\left(\frac{\beta}{2}\right)}{\Gamma\left(\frac{1}{2}+\frac{\beta}{2}\right)}.$
-We have that $\Gamma\left(\frac{1}{2}\right)$ equals $\sqrt{\pi}$. We also have that
-$\Gamma(x)=\frac{1}{x}\Gamma(x+1),$
-and that
-$\frac{\Gamma(x+1)}{\Gamma\left(x+\frac{1}{2}\right)}\leq(x+1)^{\frac{1}{2}},$
-for positive $x$. Hence
-$B\left(\frac{\beta}{2},\frac{1}{2}\right)\leq\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}.$
-Hence Expression (6.12) is bounded by
-$2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
-Let us focus then on the third term inside the parentheses of Expression (6.11). The integral can be evaluated thus giving the following equal expression:
-$\frac{2}{\beta}\left(e^{-\frac{\beta}{2}t}-e^{-\frac{\beta}{2}(t+t^{\prime})}\right),$
-which has the following upper bound
-$\frac{2\,e^{-\frac{\beta}{2}(t-t^{\prime})}}{\beta}.$
-All three terms inside the parentheses of Expression (6.11) are bounded and they give the final upper bound for the original integral:
-$\frac{32\,\pi}{\beta}\,e^{-\frac{\beta}{2}(t-t^{\prime})}\left(2\ +\ 2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\ +\ \frac{2}{\beta}\right)\leq\frac{480\pi}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
-where we especially noted that $\beta<1$ for the inequality.
+We seek to bound
+$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}\,F_{t,t^{\prime},\rho}^{2}\,d\rho$
+$\leq 50^{2}\left(\frac{1}{\sinh(t-t^{\prime})}\int_{0}^{t-t^{\prime}}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}\,d\rho+\int_{t-t^{\prime}}^{t+t^{\prime}}(1+\rho)\,e^{-\beta\rho}\,d\rho\right),$
+where the inequality follows from the bound on $F_{t,t^{\prime},\rho}$ and splitting the integration domain. Using
+$(1+\rho)\leq\frac{2}{\beta}\,e^{\frac{\beta}{2}-1}\,e^{\frac{\beta}{2}\rho}\leq\frac{2}{\beta}e^{\tfrac{\beta}{2}\rho},$
+which can be verified by analyzing derivatives, we obtain the following upper bound for the previous integral expression:
+$\frac{2\cdot 50^{2}}{\beta}\left(\frac{1}{\sinh(t-t^{\prime})}\int_{0}^{t-t^{\prime}}\sinh(\rho)\,e^{-\frac{\beta}{2}\rho}d\rho+\int_{t-t^{\prime}}^{t+t^{\prime}}e^{-\frac{\beta}{2}\rho}d\rho\right).$ (6.10)
+We have
+$\sinh(\rho)\,e^{-\tfrac{\beta}{2}\rho}\leq\sinh(t-t^{\prime})\,e^{-\tfrac{\beta}{2}(t-t^{\prime})}$
+in the first term of the previous integral expression.
+Applying this inequality to the first term of Expression (6.10) and then integrating both integrals yields the following upper bound for Expression (6.10):
+$\frac{2\cdot 50^{2}}{\beta}\left((t-t^{\prime})\,e^{-\tfrac{\beta}{2}(t-t^{\prime})}+\frac{2}{\beta}\left(e^{-\tfrac{\beta}{2}(t-t^{\prime})}-e^{-\tfrac{\beta}{2}(t+t^{\prime})}\right)\right).$ (6.11)
+Using that
+$xe^{-bx}\leq\frac{2}{e}\frac{1}{b}e^{-bx/2}$
+for all $x>0,b\in(0,1)$, Expression (6.11) is bounded by
+$\frac{10000}{\beta^{2}}\,e^{-\frac{\beta}{4}(t-t^{\prime})}.$
 7. Weight function
@@
 $=\ \int_{B(w,t)\cap B(w^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}}.$ (7.1)
-The proof is based on using the properties of the isometries in the hyperbolic plane.
+The proof is based on the properties of isometries in the hyperbolic plane.
 Proof of Lemma 7.1.
-There exists an isometry $f$ such that $f(z)$ equals $w$ and that $f(z^{\prime})$ equals $w^{\prime}$.
+There exists an isometry $f$ such that $f(z)=w$ and $f(z^{\prime})=w^{\prime}$.
 Using a change of variables $x\mapsto f^{-1}(x)$, we get the following equal expression for the left-hand side of Expression (7.1):
@@
 $=\ \int_{f\left(B(z,t)\cap B(z^{\prime},t^{\prime})\right)}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,w))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,w^{\prime}))}},$
-where the equality followed from the fact that $d(f^{-1}(x),z)=d(x,f(z))$ which then equals $d(x,w)$, and similarly $d(f^{-1}(x),z^{\prime})=d(x,w^{\prime})$.
+where the equality follows from the fact that $d(f^{-1}(x),z)=d(x,f(z))$, which then equals $d(x,w)$, and similarly $d(f^{-1}(x),z^{\prime})=d(x,w^{\prime})$.
 As $f$ maps $z$ to $w$ and $z^{\prime}$ to $w^{\prime}$, we have that
@@
 The main goal of this section is to give an upper bound for $F_{t,t^{\prime},\rho}$.
-Here, we additionally assume that $t>t^{\prime}$ for the case $t^{\prime}<t$ is symmetric.
-$F_{t,t^{\prime},\rho}$ is a crucial weight appearing in Section 6 when proving bounds for the geometric data appearing in the proof of Theorem 1.5 in Proposition 6.1.
-Bounding this weight is done in three lemmas, Lemmas 7.3, 7.4, and 7.5, of which settings are illustrated, respectively, in the following three figures:
+$F_{t,t^{\prime},\rho}$ is a crucial weight appearing in Section 6 when proving bounds for the geometric data of the main problem.
+The upper bound is given in the following proposition.
+Proposition 7.2.
+Let
+$F_{t,t^{\prime},\rho}=\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}},$
+where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. Assume that $t>t^{\prime}>0$.
+$F_{t,t^{\prime},\rho}\leq C_{1}\,\frac{1}{\sqrt{\sinh(\max(t-t^{\prime},\rho))}}\bm{1}_{\rho\in(0,t+t^{\prime})},$
+where $C_{1}=50$.
+Remark 7.3.
+The case of $t>t^{\prime}>0$ is symmetric to the case of $t^{\prime}>t>0$.
+Despite the simplicity of the upper bound, obtaining it is tedious.
+The proof is based on carefully using hyperbolic geometry.
+In case $\rho\in(-\infty,0)\cup(t+t^{\prime},\infty)$, $F_{t,t^{\prime},\rho}=0$ since the integration domain is an empty set.
+The remaining data for the bound is obtained in three separate lemmas, Lemmas 7.4, 7.5, and 7.6, as the settings differ significantly from case to case.
+The remainder of the section is dedicated to stating and proving these three lemmas.
+We state the first lemma concerning the case of $\rho\in(0,t-t^{\prime})$.
+Lemma 7.4.
+Let
+$F_{t,t^{\prime},\rho}=\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}},$
+where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. Assume that $t>t^{\prime}>0$ and that $\rho\in(0,t-t^{\prime})$. Then
+$F_{t,t^{\prime},\rho}\leq\frac{C_{2}}{\sqrt{\sinh(t-t^{\prime})}},$
+where $C_{2}=40$.
+The following figure, Figure 2, illustrates the integration domain.
 Figure 2.
-The case of $t>t^{\prime}+\rho$.
-The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is just the smaller ball $B(z^{\prime},t^{\prime})$ for it is enveloped by the bigger ball $B(z,t)$.
-The bound in this case is established using the polar coordinates around $z^{\prime}$.
+The case of $\rho\in(0,t-t^{\prime})$.
+The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is just the smaller ball $B(z^{\prime},t^{\prime})$ since it is contained within the larger ball $B(z,t)$.
+In the proof, polar coordinates around $z^{\prime}$ are used to present the integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})=B(z^{\prime},t^{\prime})$.
+The main idea is to carefully split the integration into an integration over small angles and an integration over large angles; the small angles are especially delicate, as here $\partial B(z,t)$ and $\partial B(z^{\prime},t^{\prime})$ can be problematically close.
+Proof of Lemma 7.4.
+We have that $B(z,t)\cap B(z^{\prime},t^{\prime})=B(z^{\prime},t^{\prime})$.
+Hence we write $x$ using polar coordinates around $z^{\prime}$, taking $\theta=0$ in the direction $z,z^{\prime}](z^{\prime})$.
+Hence $x=\pi_{1}(\varphi_{\rho}(z^{\prime},\theta))$.
+$R(r,\theta)$ is $d(x,z)$ in the polar coordinates.
+Thus:
+$F_{t,t^{\prime},\rho}=2\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\left(\int_{0}^{\pi}\frac{d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}}\right)dr,$ (7.2)
+where only half of the angles are considered due to symmetry, thus giving the coefficient $2$ in front. We notice that by the hyperbolic law of cosines
+$\cosh(R(r,\theta))=\cosh(r)\cosh(\rho)+\sinh(r)\sinh(\rho)\cos(\theta).$ (7.3)
+We also remark that
+$\cosh(t)=\cosh(t^{\prime})\cosh(t-t^{\prime})+\sinh(t^{\prime})\sinh(t-t^{\prime}).$ (7.4)
+Focus on the inner integral of Expression (7.2). First, write it using Expressions (7.3) and (7.4):
+$\int_{0}^{\pi}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}.$
+We split it into three terms by splitting the integration domain:
+$\int_{0}^{\theta_{0}}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}$
+$+\ \int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}$
+$+\ \int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}.$ (7.5)
+$\theta_{0}\in(0,\frac{\pi}{2})$ is specified later; in particular, $\theta_{0}$ may depend on $r$.
+We bound the first term of Expression (7.5). Since $t-t^{\prime}>\rho$ and $t^{\prime}>r$, we have the following upper bound for it:
+$\int_{0}^{\theta_{0}}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})-\cosh(r)\cosh(\rho)}}=\frac{\theta_{0}}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})-\cosh(r)\cosh(\rho)}}$
+$\leq\ \frac{\theta_{0}}{\sqrt{\cosh(t-t^{\prime})}\sqrt{\cosh(t^{\prime})-\cosh(r)}},$
+where the inequality follows from $t-t^{\prime}>\rho$.
+Let us bound the second term of Expression (7.5). Since $t-t^{\prime}>\rho$ and $t^{\prime}>r$, we have the following upper bound for it:
+$\int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\sinh(t-t^{\prime})\sinh(t^{\prime})-\sinh(r)\sinh(\rho)\cos(\theta)}}\leq\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{1-\cos(\theta)}},$
+where the inequality follows from $t-t^{\prime}>\rho$ and $t^{\prime}>r$. We have that
+$1-\cos(x)\geq\frac{2}{\pi^{2}}x^{2},\qquad x\in(0,\pi),$
+which means that the previous integral expression is bounded from above by
+$\frac{\pi}{\sqrt{2}}\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\,\int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\theta}\leq\frac{\pi^{2}}{2\sqrt{2}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\,\frac{1}{\theta_{0}}.$
+Let us bound the third term of Expression (7.5). Since $t-t^{\prime}>\rho$ and $t^{\prime}>r$, we have the following upper bound for it:
+$\int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}=\frac{\pi}{2}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}$
+We now have an upper bound for Expression (7.5) and thus an upper bound for Expression (7.2), the latter being:
+$2\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\cosh(t^{\prime})-\cosh(r)}\frac{\theta_{0}}{\sqrt{\cosh(t-t^{\prime})}}+\frac{\pi^{2}}{\sqrt{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\,\frac{1}{\theta_{0}}$
+$+\ \pi\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}\sqrt{\sinh(t-t^{\prime})\sinh(t^{\prime})}}$
+We choose
+$\theta_{0}=\min\left(\frac{\pi}{2},\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha}\right),$
+with $\alpha\in(0,\frac{1}{2})$. The previous integral expression has the following upper bound:
+$2\,\frac{1}{\sqrt{\cosh(t-t^{\prime})}\,\sinh(t^{\prime})^{\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{1-\alpha}}$
+$+\ \frac{\pi^{2}}{\sqrt{2}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{\frac{1}{2}+\alpha}}$
+$+\ \pi\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}\sqrt{\sinh(t-t^{\prime})\sinh(t^{\prime})}}$
+The upper bound was obtained by replacing $\theta_{0}$ with
+$\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha},$
+which is possible since the middle term would vanish if the previous ratio is larger than $\frac{\pi}{2}$.
+The previous integral expression has the following equivalent expression obtained by integrating:
+$2\,\frac{1}{\sqrt{\cosh(t-t^{\prime})}\,\sinh(t^{\prime})^{\alpha}}\,\frac{(\cosh(t^{\prime})-1)^{\alpha}}{\alpha}+\frac{\pi^{2}}{\sqrt{2}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\frac{(\cosh(t^{\prime})-1)^{\frac{1}{2}-\alpha}}{\frac{1}{2}-\alpha}$
+$+\ \pi\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}}.$
+By choosing $\alpha=\frac{1}{4}$, we have the following upper bound for the previous expression:
+$40\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}}.$
+This is also the desired upper bound for $F_{t,t^{\prime},\rho}$.
+We state the second lemma concerning the case of $\rho\in(t-t^{\prime},t)$.
+Lemma 7.5.
+Let
+$F_{t,t^{\prime},\rho}=\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}},$
+where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. Assume that $t>t^{\prime}>0$. Assume also that $\rho\in(t-t^{\prime},t)$. Then
+$F_{t,t^{\prime},\rho}\leq\frac{C_{3}}{\sqrt{\sinh(t-t^{\prime})}},$
+where $C_{3}=50$.
+The following figure, Figure 3, illustrates the integration domain.
 Figure 3.
-The case of $\rho<t<t^{\prime}+\rho$, where $a+b=\rho$.
+The case of $\rho\in(t-t^{\prime},t)$.
 The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls.
-The middle point $z^{\prime}$ of the smaller ball is contained in the bigger ball.
-The integration domain is split into two parts, $D_{1}$ and $D_{2}$ by two geodesics of length $t^{\prime}$ going from $z^{\prime}$ to points $w$ and $w^{\prime}$ which are the points where the two balls intersect.
-The bound in this case is established separately for $D_{1}$ and $D_{2}$.
-For $D_{1}$, we use the polar coordinates around $z^{\prime}$ and for $D_{2}$ the polar coordinates around $z$.
+The center $z^{\prime}$ of the smaller ball is contained in the larger ball.
+The integration domain is split into two parts, $D_{1}$ and $D_{2}$, by two geodesics of length $t^{\prime}$ going from $z^{\prime}$ to the points $w$ and $w^{\prime}$ where the two balls intersect.
+In this case, the integration domain is not a ball, since $\partial B(z,t)$ and $\partial B(z^{\prime},t^{\prime})$ intersect; call the intersections $w$ and $w^{\prime}$.
+The segments $z^{\prime},w]$ and $z^{\prime},w^{\prime}]$ split the integration domain into two disjoint parts, which are treated separately in the proof.
+For the other side ($D_{1}$ in Figure 3), the idea is to use a similar approach as in the proof of Lemma 7.4: use polar coordinates around $z^{\prime}$ and deal with the small and large angles separately.
+The remaining side ($D_{2}$ in Figure 3) is easier and can be handled using polar coordinates around $z$.
+Proof of Lemma 7.5.
+By our assumption, we know $\partial B(z,t)$ and $\partial B(z^{\prime},t^{\prime})$ intersect at exactly two points.
+Call these points $w$ and $w^{\prime}$.
+The segments $z^{\prime},w]$ and $z^{\prime},w^{\prime}]$ split the integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ into two disjoint parts, $D_{1}$ and $D_{2}$.
+Hence $F_{t,t^{\prime},\rho}$ can be written as
+$\int_{D_{1}}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}$
+$+\ \int_{D_{2}}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}.$ (7.6)
+We bound the two terms separately.
+We begin by bounding the integration over $D_{1}$.
+We write $x$ using polar coordinates around $z^{\prime}$, taking $\theta=0$ as the angle in the direction $z,z^{\prime}](z^{\prime})$.
+Hence $x=\pi_{1}(\varphi_{r}(z^{\prime},\theta))$.
+$R(r,\theta)$ is $d(x,z)$ in the polar coordinates.
+We notice that the integration domain of the angles is $(\Theta,2\pi-\Theta)$, where $\Theta$ is the angle between $z,z^{\prime}](z^{\prime})$ and $z^{\prime},w](z^{\prime})$.
+The following figure, Figure 4, illustrates this.
 Figure 4.
-The case of $t<\rho$.
-The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls.
-The middle point $z^{\prime}$ of the smaller ball is not contained in the bigger ball.
-The bound in this case is established using the polar coordinates around $m$ which is the intersection of geodesics $z,z^{\prime}]$ and $w,w^{\prime}]$, where $w$ and $w^{\prime}$ are the intersections of the two balls.
-From these three lemmas we get the following bound for $F_{t,t^{\prime},\rho}$.
-Proposition 7.2.
-$F_{t,t^{\prime},\rho}\leq C_{1}\,\cdot\,\begin{cases}\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad&\text{ if }\rho<t-t^{\prime}\\ \frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{\frac{1}{8}},\qquad&\text{ if }t-t^{\prime}<\rho<t\\ \frac{1}{\sqrt{\sinh(\rho)}},\qquad&\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
-where
-$C_{1}=\max\left(C_{2},C_{3},C_{4}\right)=\max(32,32\pi,99)=32\pi.$
-The following lemma gives the bound for $F_{t,t^{\prime},\rho}$ in the first case, see Figure 2.
-Lemma 7.3.
-Assume that $t$ is larger than $t^{\prime}+\rho$. Then
-$F_{t,t^{\prime},\rho}\leq\frac{C_{2}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
-where
-$C_{2}=32.$
-The proof of lemma is based on using two different lower bounds for $\cosh(t)-\cosh(R)$, when working in the polar coordinates around $z^{\prime}$.
-Using these two bounds simultaneously allows us to deal with the problematic small angles, when the boundaries of two balls are close to each other.
-Proof of Lemma 7.3.
-We write the integral using the polar coordinates around $z^{\prime}$, taking $\theta=0$ in the direction $z,z^{\prime}](z^{\prime})$.
-Hence $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$.
-The situation is illustrated in Figure 2.
-Hence $F_{t,t^{\prime},\rho}$ is written as
-$\int_{0}^{t^{\prime}}\int_{-\pi}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
-$=\qquad$ $2\int_{0}^{t^{\prime}}\int_{0}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}},$ (7.2)
-where half of the angles are dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ using the polar coordinates $r$ and $\theta$.
-We want to find a lower bound for $\cosh(t)-\cosh(R)$. We remark that we have the following inequality since $t>t^{\prime}+\rho>R$:
-$\cosh(t)-\cosh(R)\geq\cosh(t)-\cosh(t^{\prime}+\rho).$
-This implies the following inequality by using the mean value theorem and the convexity of the hyperbolic sine for the positive input:
-$\cosh(t)-\cosh(R)\geq\sinh(t^{\prime})\,\sinh(\rho)\,(t-t^{\prime}-\rho).$ (7.3)
-We can also get another inequality. We write $\cosh(t)$ as follows:
-$\cosh(t)=\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime}).$
-Then, by using the hyperbolic law of cosines, we have that
-$\cosh(R)=\cosh(\rho)\cosh(r)+\sinh(\rho)\sinh(r)\cos(\theta).$
-Hence $\cosh(t)-\cosh(R)$ equals
-$\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$
-Since $r<t^{\prime}$ and $\rho<t-t^{\prime}$, we have the following inequality:
-$\cosh(t)-\cosh(R(r,\theta))\geq(1-\cos(\theta))\,\sinh(t)\,\sinh(t-t^{\prime}).$ (7.4)
-We write Expression (7.2) as
-$2\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}d\theta dr,$
-which has the following upper bound
-$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(1-\cos(\theta))^{\frac{1}{4}}}d\theta dr,$
-which was acquired by using the inequalities appearing in Expressions (7.3) and (7.4). The following equal expressions are acquired by evaluating the integrals:
-$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,2^{\frac{5}{4}}\,K\left(\frac{1}{2}\right)dr$
-$=\ \frac{2^{2+\frac{5}{4}}\,K(\frac{1}{2})\,\sqrt{\cosh(t^{\prime}-1)}}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
-where $K$ is the complete elliptic integral of the first kind (see [Olv10, Chp.19]). The previous expression has a further upper bound, where we also bound the elliptic integral by $2$:
-$\frac{2^{\frac{17}{4}}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\leq 32\,\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}.$
-We proceed to a lemma bounding the second case, see Figure 3.
-Lemma 7.4.
-Assume that $\rho<t<\rho+t^{\prime}$. Then:
-$F_{t,t^{\prime},\rho}\leq C_{3}\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},$
-where
-$C_{3}=32\pi.$
-The idea of the proof is to split the integration domain into two separate domains.
-The bounding of the other domain resembles the approach of the previous lemma in that we establish two different lower bounds for $\cosh(t)-\cosh(R)$, which allow us to get a bound when used simultaneously.
-Bounding the other domain turns out to be simple.
-The following figure, Figure 5, further illustrates these separate domains:
-((a)) The integration domain $D_{1}$ consists of those points that are outside the triangle formed in the ball $B(z^{\prime},t^{\prime})$.
-In this case, the polar coordinates around $z^{\prime}$ are used.
-((b)) The integration domain $D_{2}$ consists of those points that are inside the triangle formed in the ball $B(z^{\prime},t^{\prime})$. In this case, polar coordinates around $z$ are used.
-Figure 5. The case is split into two case which are bounded separately.
-Proof of Lemma 7.4.
-The integration domain of $F_{t,t^{\prime},\rho}$ can be split into a sum of integration over two disjoint domains, $D_{1}$ and $D_{2}$, see Figure 3.
-We bound these separately.
-$D_{1}$ is better illustrated in Figure 5(a) and $D_{2}$ in Figure 5(b).
-Focus on $D_{1}$ first.
-We write it using the polar coordinates around $z^{\prime}$.
-Thus $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$, where $\theta=0$ corresponds to direction $z^{\prime},z](z^{\prime})$.
-In this way, the integral over $D_{1}$ equals
-$\int_{-\Theta}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
-= $=\ 2\int_{0}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
-where half the angles were dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ in the polar coordinates $r$ and $\theta$.
-Notice that the maximum angle $\Theta$ is larger than $\pi/2$.
-We split the integral into two terms as follows:
-$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
-+ $+\ 2\int_{\frac{\pi}{2}}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$ (7.5)
-By using the hyperbolic law of cosines we get that
-$\cosh(t)=\cosh(\rho)\cosh(t^{\prime})-\sinh(\rho)\sinh(t^{\prime})\cos(\Theta),$ (7.6)
-$\cosh(R)=\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$ (7.7)
-Hence $\cosh(t)-\cosh(R)$ equals
-$\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\cos(\theta)\sinh(r)-\cos(\Theta)\sinh(t^{\prime})).$ (7.8)
-Hence we have the following inequalities:
-$\cosh(t)-\cosh(R)\geq\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))$ (7.9)
-and
-$\cosh(t)-\cosh(R)\geq\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r)).$ (7.10)
-Using the inequality of Expression (7.9), we get that the first term of Expression (7.5) is bounded by
-$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
-By evaluating the integral over $r$, we get the following equal expression:
-$\frac{4}{\sqrt{\sinh(\rho)}}\,\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)-\cos(\Theta)}}\,d\theta.$
-This has the following upper bound, when we notice that both $\cos(\theta)$ and $\cos(\Theta)$ are positive. Additionally the ratio with the hyperbolic sine and cosine is bounded by $1$:
-$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)}}\,d\theta.$
-The following bound follows from the well known upper bound for the integral:
-$\frac{12}{\sqrt{\sinh(\rho)}}.$
-Let us focus on the second term of Expression (7.5). We write it as follows by interchanging the integrals using Fubini’s theorem:
-$2\,\int_{0}^{t^{\prime}}\int_{\frac{\pi}{2}}^{\Theta}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}\,d\theta\,dr.$
-We use the two inequalities appearing in Expressions (7.9) and (7.10) to get the following upper bound for the previous expression:
-$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\cos(\theta)-\cos(\Theta))^{\frac{1}{4}}}\,d\theta\,dr.$
-Using the mean value theorem, we get the following upper bound:
-$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\Theta-\theta)^{\frac{1}{4}}}d\theta dr.$
-Integrating the inner integral yields
-$\frac{4}{3}\,\left(\Theta-\frac{\pi}{2}\right)^{\frac{3}{4}}.$
-Hence we have the following bound for the previous integral expression:
-$\frac{4\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}dr$
-$=\ \frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}},$ (7.11)
-where the equality followed from the evaluation of the remaining integral.
-Focus momentarily on $\sin(\Theta)$. Expression (7.6) can be modified into the following form using the hyperbolic identities:
-$\cos(\Theta)=1-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}.$
-The subtracted division term must be larger than $1$ as the cosine is negative due to $\Theta$ being in $(\frac{\pi}{2},\pi)$. We have that
-$\sin(\arccos(1-x))=\sqrt{x}\sqrt{2-x}\geq\sqrt{2-x},$
-when $x>1$. Hence:
-$\sin(\Theta)\geq\sqrt{2-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}}=\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\sinh(t^{\prime})\sinh(\rho)}},$
-where the equality followed using hyperbolic identities. The previous implies the following inequality:
-$\sin(\Theta)\geq\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\cosh(t^{\prime}+\rho)}}=\sqrt{1-\frac{\cosh(t)}{\cosh(t^{\prime}+\rho)}}.$
-This then gives us the following inequality:
-$\sin(\Theta)\geq\sqrt{1-\frac{1}{\cosh(t^{\prime}+\rho-t)}}.$ (7.12)
-Applying the inequality of Expression (7.12) in Expression (7.11) gives the following upper bound:
-$\frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
-This has the following further upper bound:
-$\frac{16\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
-The integration over $D_{1}$ has the following upper bound, when both of the terms of Expression (7.5) are added together:
-$\frac{1}{\sqrt{\sinh(\rho)}}\left(12+16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\right)$
-$\leq\ 16\pi\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
-Next, we will focus on bounding the integral over $D_{2}$.
-We write it using the polar coordinates around $z$.
-Thus $x=\pi_{1}(\phi_{R}(z,\theta))$, where $\theta=0$ corresponds to direction $z,z^{\prime}](z)$.
-Then the integral equals:
-$\int_{-\Psi}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta$
-= $=\ 2\int_{0}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}dRd\theta,$
-where the equality followed from using the symmetry.
-$r=r(R,\theta)$ is $d(x,z^{\prime})$ in the polar coordinates $R$ and $\theta$.
-$R(\theta)$ denotes the minimum distance that needs to be propagated into direction $\theta$ so that we are in $D_{2}$.
-Notice also that $\Psi$ is less than $\pi/2$.
-By using the hyperbolic law of cosines we have that
-$\cosh(t^{\prime})=\cosh(\rho)\cosh(t)-\sinh(\rho)\sinh(t)\cos(\Psi),$
-$\cosh(r)=\cosh(\rho)\cosh(R)-\sinh(\rho)\sinh(R)\cos(\theta).$
-As $\cosh(r)$ grows as $R$ grows, we have the following lower bound for $\cosh(t^{\prime})-\cosh(r)$:
-$\sinh(\rho)\sinh(t)(\cos(\theta)-\cos(\Psi)).$
-Hence the integral expression has the following upper bound:
-$\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\frac{1}{\sqrt{\sinh(t)}}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}}dRd\theta.$
-This has an upper bound when we integrate from 0 instead of $R(\theta)$; the following expression is acquired after we evaluate the integral:
-$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{\sqrt{\cosh(t)-1}}{\sqrt{\sinh(t)}}d\theta.$
-This is then bounded by
-$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\Psi}\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}d\theta,$
-which has the following upper bound using the mean value theorem:
-$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\sin(\theta)}}.$
-Using the well-known bound for sine, we get the following upper bound:
-$\frac{4}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{1}{\sqrt{\Psi-\theta}\,\sqrt{\frac{2}{\pi}\theta}}d\theta.$
-This is actually the Beta integral (see [Olv10, Chp. 5]), so we get the following equal expression:
-$\frac{4\sqrt{\pi}\pi}{\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)}}.$
-The integration over both $D_{1}$ and $D_{2}$ is done. Uniting the bounds of these two cases gives us the final bound for $F_{t,t^{\prime},\rho}$:
-$\frac{1}{\sqrt{\sinh(\rho)}}\left(16\pi\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}\ +\ \frac{4\pi\sqrt{\pi}}{\sqrt{2}}\right)$
-$\leq\qquad$ $\frac{32\pi}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
-The following lemma takes care of the last case, see Figure 4.
+The integration domain $D_{1}$ consists of the points that are outside the sector formed by $w,z^{\prime},$ and $w^{\prime}$ in ball $B(z^{\prime},t^{\prime})$.
+Polar coordinates around $z^{\prime}$ are used.
+Hence the first term of Expression (7.6) equals
+$2\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\left(\int_{\Theta}^{\pi}\frac{d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}}\right)dr,$ (7.7)
+where only half of the angles are considered due to symmetry, giving the coefficient $2$ in front. Using the hyperbolic law of cosines, we record two useful equalities:
+$\cosh(t)=\cosh(\rho)\cosh(t^{\prime})+\sinh(\rho)\sinh(t^{\prime})\cos(\Theta),$ (7.8)
+$\cosh(R(r,\theta))=\cosh(\rho)\cosh(r)+\sinh(\rho)\sinh(r)\cos(\theta).$ (7.9)
+Focus on the inner integral of Expression (7.7). First, we write it using Expressions (7.8) and (7.9):
+$\int_{\Theta}^{\pi}\frac{d\theta}{\sqrt{\cosh(\rho)\cosh(t^{\prime})+\sinh(\rho)\sinh(t^{\prime})\cos(\Theta)-\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta)}}.$
+We then split it into three terms by dividing the integration domain:
+$\int_{\Theta}^{\Theta+\theta_{0}}\frac{d\theta}{\sqrt{\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))}}$
+$+\ \int_{\Theta+\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))}}$
+$+\ \int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))}}.$ (7.10)
+We remark that $\Theta\in(0,\frac{\pi}{2})$ since $z^{\prime}$ is contained in $B(z,t)$ as $\rho\in(t-t^{\prime},t)$.
+$\theta_{0}\in(0,\frac{\pi}{2}-\Theta)$ and will be specified later.
+We note that $\theta_{0}$ can depend on $r$.
+We bound these terms separately.
+We start by bounding the first term of Expression (7.10). It is bounded from above by
+$\int_{\Theta}^{\Theta+\theta_{0}}\frac{d\theta}{\sqrt{\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r))}}=\frac{\theta_{0}}{\sqrt{\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r))}},$
+since $\sinh(\rho)(\sinh(t^{\prime})\cos(\Theta)-\sinh(r)\cos(\theta))$ is positive.
+Next, we bound the second term of Expression (7.10). It is bounded from above by
+$\int_{\Theta+\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\sinh(\rho)}\,\sqrt{\sinh(t^{\prime})}\sqrt{\cos(\Theta)-\cos(\theta)}},$
+since $\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r))$ is positive and $t^{\prime}>r$. By examining the behavior of the function, we notice that
+$\cos(\Theta)-\cos(\theta)\geq 1-\cos(\theta-\Theta),\qquad\theta\in(\Theta,\tfrac{\pi}{2}).$
+Using this gives the following upper bound for the previous integral:
+$\frac{1}{\sqrt{\sinh(\rho)}\,\sqrt{\sinh(t^{\prime})}}\int_{\Theta+\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{1-\cos(\theta-\Theta)}}\leq\frac{1}{\sqrt{\sinh(\rho)}\,\sqrt{\sinh(t^{\prime})}}\frac{\frac{\pi}{2}}{\sqrt{1-\cos(\theta_{0})}}.$
+This is bounded from above by
+$\frac{\pi^{2}}{2\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\frac{1}{\theta_{0}}$
+using
+$1-\cos(x)\geq\frac{2}{\pi^{2}}x^{2},\qquad x\in(-\pi,\pi).$
+Finally, we bound the third term of Expression (7.10). It is bounded from above by
+$\int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})\,(-\cos(\theta))}}\leq\frac{3}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}},$
+since $\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))$ and $\sinh(\rho)\sinh(t^{\prime})\cos(\Theta)$ are positive.
+The inequality is obtained by integrating $\theta$ and then bounding the result from above by $3$.
+The three terms of Expression (7.10) are bounded. Hence we have an upper bound for Expression (7.10), and thus also for Expression (7.7), which is:
+$\frac{2}{\sqrt{\cosh(\rho)}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,\theta_{0}\,dr}{\cosh(t^{\prime})-\cosh(r)}+\frac{\pi^{2}}{\sqrt{2}}\,\frac{1}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\theta_{0}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
+$+\ \frac{6}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
+We choose
+$\theta_{0}=\min\left(\frac{\pi}{2}-\Theta,\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha}\right),$
+with $\alpha\in(0,\frac{1}{2})$. The previous integral is then bounded from above by
+$\frac{2}{\sqrt{\cosh(\rho)}\sinh(t^{\prime})^{\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{1-\alpha}}$
+$+\ \frac{\pi^{2}}{\sqrt{2}}\frac{1}{\sqrt{\sinh(\rho)}\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{\frac{1}{2}+\alpha}}$
+$+\ \frac{6}{\sqrt{\sinh(\rho)\,\sinh(t^{\prime})}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
+The upper bound was obtained by replacing $\theta_{0}$ with
+$\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha};$
+this is possible since the middle term would vanish if
+$\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha}\geq\frac{\pi}{2}-\Theta.$
+Evaluating the integrals gives
+$\frac{2}{\alpha}\,\frac{(\cosh(t^{\prime})-1)^{\alpha}}{\sinh(t^{\prime})^{\alpha}}\,\frac{1}{\sqrt{\cosh(\rho)}}+\frac{\pi^{2}}{\sqrt{2}(\frac{1}{2}-\alpha)}\,\frac{(\cosh(t^{\prime})-1)^{\frac{1}{2}-\alpha}}{\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\,\frac{1}{\sqrt{\sinh(\rho)}}+6\,\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\,\frac{1}{\sqrt{\sinh(\rho)}}.$
+Choosing $\alpha=\frac{1}{4}$ yields the upper bound:
+$\frac{42}{\sqrt{\sinh(\rho)}}.$
+This finishes the bound for the integration over $D_{1}$ in Expression (7.6).
+We proceed to bound the integration over $D_{2}$ in Expression (7.6).
+We write $x$ using polar coordinates around $z$, taking $\theta=0$ in the direction $z,z^{\prime}](z)$.
+Hence $x=\pi_{1}(\varphi_{R}(z,\theta))$.
+$r(R,\theta)$ is $d(x,z^{\prime})$ in the polar coordinates.
+We notice that the integration domain of the angles is $(-\Psi,\Psi)$, where $\Psi$ is the angle between $z,z^{\prime}](z)$ and $[z,w](z)$.
+Furthermore, the integration domain of the distances depends on the angle: $R\in(R(\theta),t)$, where $R(\theta)$ is the minimum distance required to reach the integration domain in the direction $\theta$.
+Figure 5 illustrates this.
+Figure 5.
+The integration domain $D_{2}$ consists of points in the area surrounded by the segments $z^{\prime},w]$ and $z^{\prime},w^{\prime}]$ and the arc between $w$ and $w^{\prime}$ on $\partial B(z,t)$.
+Hence the second term of Expression (7.6) equals
+$2\int_{0}^{\Psi}\int_{R(\theta)}^{t}\frac{\sinh(R)}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r(R,\theta))}}dR\,d\theta,$ (7.11)
+where only half of the angles are considered due to symmetry, giving the coefficient $2$ in front. Using the hyperbolic law of cosines, we record two useful equalities:
+$\cosh(t^{\prime})=\cosh(t)\cosh(\rho)-\sinh(t)\sinh(\rho)\cos(\Psi),$
+$\cosh(r(R,\theta))=\cosh(R)\cosh(\rho)-\sinh(R)\sinh(\rho)\cos(\theta).$
+We remark that $r(R,\theta)$ grows as $R$ grows. Hence we have the following inequality using the previous equalities:
+$\cosh(t^{\prime})-\cosh(r(R,\theta))\geq\sinh(t)\,\sinh(\rho)\,(\cos(\theta)-\cos(\Psi)).$ (7.12)
+Focus on the inner integral of Expression (7.11). We acquire an upper bound by using the inequality of Expression (7.12):
+$\int_{R(\theta)}^{t}\frac{\sinh(R)\,dR}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\sinh(\rho)\,\sinh(t)\,(\cos(\theta)-\cos(\Psi))}}$
+$\leq\ \frac{\sqrt{\cosh(t)-1}}{\sqrt{\sinh(t)}}\,\frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{1}{\sqrt{\sinh(\rho)}}$
+$\leq\ \frac{1}{\sqrt{\cos(\theta)-\cos(\Psi)}}\,\frac{1}{\sqrt{\sinh(\rho)}},$
+where the first inequality is obtained by evaluating the integral over the enlarged domain $(0,t)$.
+We now have a bound for the inner integral of Expression (7.11), giving the following upper bound for the whole expression:
+$\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{d\theta}{\sqrt{\cos(\theta)-\cos(\Psi)}}\leq\frac{2}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{d\theta}{\sqrt{\Psi-\theta}\,\sqrt{\sin(\theta)}}.$
+The inequality follows from the mean value theorem. The previous expression is further bounded by
+$\sqrt{2}\sqrt{\pi}\,\frac{1}{\sqrt{\sinh(\rho)}}\int_{0}^{\Psi}\frac{d\theta}{\sqrt{\theta}\,\sqrt{\Psi-\theta}},$
+using that
+$\sin(x)\geq\frac{2}{\pi}x,\qquad x\in(-\frac{\pi}{2},\frac{\pi}{2}).$
+The integral is Euler’s beta integral and equals $\pi$, see [LOZ03, Chp. 5.12.]. Hence the expression equals
+$\sqrt{2}\pi\sqrt{\pi}\,\frac{1}{\sqrt{\sinh(\rho)}}.$
+We have now bounded the integration over $D_{2}$ in Expression (7.6).
+Combining the bounds for $D_{1}$ and $D_{2}$, we get the final upper bound for $F_{t,t^{\prime},\rho}$:
+$(42+\sqrt{2}\pi\sqrt{\pi})\,\frac{1}{\sqrt{\sinh(\rho)}}\leq 50\,\frac{1}{\sqrt{\sinh(\rho)}}.$
+This concludes the proof of Lemma 7.5.
+We state the third lemma concerning the case of $\rho\in(t,t+t^{\prime})$.
 Lemma 7.6.
@@
 $F_{t,t^{\prime},\rho}\leq C_{4}\,\frac{1}{\sqrt{\sinh(\rho)}},$ (7.13)
-where
-$C_{4}=99.$
+where $C_{4}=38$.
+The following figure, Figure 6, illustrates the integration domain.
+Figure 6.
+The case of $\rho\in(t,t+t^{\prime})$.
+The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls.
+The middle point $z^{\prime}$ of the smaller ball is not contained in the bigger ball.
+The integration domain is split into two parts by $w,w^{\prime}]$.
+Notice that $a+b=\rho$, the separation of $z$ and $z^{\prime}$.
 The proof is again based on splitting the integration domain into two.
-This time the splitting is done by $w,w^{\prime}]$ which splits the integration into two sides, one closer to $z$ and one closer to $z^{\prime}$.
-The technique used to bound both sides is the same, based on establising carefully bounds using hyperbolic geometry.
+This time the splitting is done by $w,w^{\prime}]$, which separates the domain into two sides: one closer to $z$ and one closer to $z^{\prime}$.
+Here, $w$ and $w^{\prime}$ are points of $\partial B(z,t)\cap\partial B(z^{\prime},t^{\prime})$.
+The technique used to bound both sides is the same and relies on hyperbolic geometry, in particular the hyperbolic law of cosines.
 Proof of Lemma 7.6.
@@
 Let $w$ and $w^{\prime}$ be the intersections of the boundaries of the balls, and let $m$ be the intersection of $z,z^{\prime}]$ with $w,w^{\prime}]$.
-We parametrize the domain using the polar coordinates around $m$ with $\theta=0$ corresponding to direction $z,z^{\prime}](m)$.
+We parametrize the domain using polar coordinates around $m$, with $\theta=0$ corresponding to the direction $z,z^{\prime}](m)$.
 Hence $x=\pi_{1}(\phi_{r}(m,\theta))$.
-To easen the notation, we call $|[z,m]|=a$, $|[m,z^{\prime}]|=b$, and $|[m,w]|=h$.
-We will present $d(z,x)$ and $d(z^{\prime},x)$ using polar coordinates, denoted by $R=R(r,\theta)$ and $P=P(r,\theta)$, respectfully.
-For illustration, see Figure 4.
-Hence $F_{t,t^{\prime},\rho}$ is in polar coordinates
+To ease notation, we let $|[z,m]|=a$, $|[m,z^{\prime}]|=b$, and $|[m,w]|=h$.
+We denote $d(z,x)$ and $d(z^{\prime},x)$ in polar coordinates by $R=R(r,\theta)$ and $P=P(r,\theta)$, respectively.
+Then $F_{t,t^{\prime},\rho}$ can be written as
 $\int_{0}^{2\pi}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}},$
-where $r(\theta)$ is the maximum distance we can go from $m$ to direction $\theta$ to still stay in the integration domain. We can split the integration into two parts:
+where $r(\theta)$ is the maximal distance from $m$ in direction $\theta$ within the integration domain. We split the integration into two parts:
 $\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}}$
 $+\ \int_{\frac{\pi}{2}}^{\frac{3\pi}{2}}\int_{0}^{r(\theta)}\frac{\sinh r\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(P(r,\theta))}}.$ (7.14)
-The first term integrates over the side closer to $z^{\prime}$ while the second term integrates the side closer to $z$, these sides are separated by $w,w^{\prime}]$.
-Let us focus on the integrand of the first term. Using the hyperbolic law of cosines, we get the following equalities:
-$\cosh(t)=\cosh(a)\cosh(r(\theta))+\sinh(a)\sinh(r(\theta))\cos(\theta),$
-$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
-$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
-where we note that $\cos(\theta)$ is positive. We also notice that by using the hyperbolic law of cosines we get that
+The first term integrates over the side closer to $z^{\prime}$, while the second integrates over the side closer to $z$; these sides are separated by $w,w^{\prime}]$.
+For the first term of Expression (7.14), the hyperbolic law of cosines gives:
+$\cosh(t)=\cosh(a)\cosh(r(\theta))+\sinh(a)\sinh(r(\theta))\cos(\theta),$ (7.15)
+$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$ (7.16)
+$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$ (7.17)
+where $\cos(\theta)>0$. Also, we have
 $\cosh(t)=\cosh(a)\cosh(h),$
 $\cosh(t^{\prime})=\cosh(b)\cosh(h),$
-which implies that
+which implies
 $\cosh(t^{\prime})=\frac{\cosh(b)}{\cosh(a)}\cosh(t).$
-Inserting the earlier equality of $\cosh(t)$ into the previous yields the following:
-$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))+\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))\cos(\theta).$
-We want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
-$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\sinh(a)\left(\sinh(r(\theta))-\sinh(r)\right),$
-which then has a lower bound:
-$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right).$
-Next, we want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
-$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right),$
-and then has a lower bound
-$\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
-Thus $(\cosh(t)-\cosh(R))(\cosh(t^{\prime})-\cosh(P))$ is bounded from below by
-$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right).$
-This equals
-$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(b)\sinh(a)\sinh(r(\theta))+\cosh(a)\sinh(b)\sinh(r)\right),$
-and is then bounded from below by
-$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(b)\sinh(a)+\cosh(a)\sinh(b)\right).$
-As $a+b=\rho$, the previous equals
-$\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
-Now, let us focus on the integrand of the second term of Expression (7.14). Using the hyperbolic law of cosines, we get the following equalities:
-$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))-\sinh(b)\sinh(r(\theta))\cos(\theta),$
-$\cosh(R)=\cosh(a)\cosh(r)+\sinh(a)\sinh(r)\cos(\theta),$
-$\cosh(P)=\cosh(b)\cosh(r)-\sinh(b)\sinh(r)\cos(\theta),$
-where $\cos(\theta)$ is negative. We also notice that, using the hyperbolic law of cosines,
-$\cosh(t)=\cosh(a)\cosh(h),$
-$\cosh(t^{\prime})=\cosh(b)\cosh(h),$
-which implies that
-$\cosh(t)=\frac{\cosh(a)}{\cosh(b)}\cosh(t^{\prime}).$
-The previous gives the following equality after inserting an earlier equality of $\cosh(t^{\prime})$ in:
-$\cosh(t)=\cosh(a)\cosh(r(\theta))-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))\cos(\theta).$
-We want to get a lower bound for $\cosh(t^{\prime})-\cosh(P)$. This equals
-$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)-\cos(\theta)\sinh(b)\left(\sinh(r(\theta))+\sinh(r)\right).$
-This has a lower bound
-$\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right).$
-Next, we want to get a lower bound for $\cosh(t)-\cosh(R)$. This equals
-$\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right)+\cos(\theta)\left(-\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))-\sinh(a)\sinh(r)\right),$
-and then has a lower bound
-$(-1)\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
-Thus $(\cosh(t)-\cosh(R(r,\theta)))(\cosh(t^{\prime})-\cosh(P(r,\theta)))$ is bounded from below by
-$(-1)\,\cosh(b)\left(\cosh(r(\theta))-\cosh(r)\right)\,\cos(\theta)\left(\frac{\cosh(a)}{\cosh(b)}\sinh(b)\sinh(r(\theta))+\sinh(a)\sinh(r)\right).$
-This equals
-$(-1)\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\left(\cosh(a)\sinh(b)\sinh(r(\theta))+\cosh(b)\sinh(a)\sinh(r)\right),$
-and is then bounded from below by
-$(-1)\,\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\left(\cosh(a)\sinh(b)+\cosh(b)\sinh(a)\right).$
-As $a+b=\rho$, the previous equals
+Inserting Equality (7.15) gives
+$\cosh(t^{\prime})=\cosh(b)\cosh(r(\theta))+\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))\cos(\theta).$ (7.18)
+Lower bounds for the differences are
+$\cosh(t)-\cosh(R)\geq\cosh(a)\left(\cosh(r(\theta))-\cosh(r)\right),$
+$\cosh(t^{\prime})-\cosh(P)\geq\cos(\theta)\left(\frac{\cosh(b)}{\cosh(a)}\sinh(a)\sinh(r(\theta))+\sinh(b)\sinh(r)\right),$
+and their product satisfies
+$(\cosh(t)-\cosh(R))(\cosh(t^{\prime})-\cosh(P))\geq\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
+For the second term of Expression (7.14), using the same approach (the roles of $t,R$ and $t^{\prime},P$ change), we have the following lower bound:
 $(-1)\,\cos(\theta)\,(\cosh(r(\theta))-\cosh(r))\,\sinh(r)\,\sinh(\rho).$
-These observations for the integrands of the terms of Expression (7.14) imply that we get the following upper bound for $F_{t,t^{\prime},\rho}$:
-$\frac{1}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\int_{0}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\,d\theta$ (7.15)
-Focus on the inner integral. Assume first that $r(\theta)<2$. The integral is bounded by the following by using the mean value theorem:
-$\int_{0}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)}.$
-This is bounded by $2\sqrt{2}$. Then, assume that $r(\theta)>2$. We split the integration as follows:
-$\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr.$
-$a$ is specified later. We have the following upper bound for the previous:
-$\frac{1}{\sqrt{\sinh(a)}}\,\int_{0}^{a}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}}\,dr\ +\ \int_{a}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr.$
-The first term was bounded using the boundary of the integration domain and the second term was bounded using the mean value theorem.
-By evaluating the integrals, we acquire the following equal expression:
-$\frac{2\,\sqrt{\cosh(r(\theta))-\cosh(a)}}{\sqrt{\sinh(a)}}\ +\ 2\,\sqrt{r(\theta)-a}\leq\frac{2\,\sqrt{r(\theta)-a}\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta))}}\ +\ 2\,\sqrt{r(\theta)-a}.$
-The inequality followed from using the mean value theorem. We choose $a$ to equal $r(\theta)-1$. Then the previous equals
-$\frac{2\,\sqrt{\sinh(r(\theta))}}{\sqrt{\sinh(r(\theta)-1)}}\ +\ 2\leq\frac{2\,\sqrt{\sinh(2)}}{\sqrt{\sinh(1)}}\ +\ 2,$
-where for the inequality we used that $r(\theta)>2$. The previous is bounded by $9$.
-The previous observations mean that Expression (7.15) has the following upper bound:
-$\frac{9}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\,d\theta.$
-The integral is bounded from above by $11$, so we can conclude that
-$F_{t,t\,\rho}\leq\frac{99}{\sqrt{\sinh(\rho)}}.$
+Notice that 1-1 comes from the fact that $\cos(x)$ is negative when $x\in(\tfrac{\pi}{2},\tfrac{3\pi}{2})$. These bounds give the following upper bound for $F_{t,t^{\prime},\rho}$:
+$\frac{1}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\int_{0}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\,d\theta.$ (7.19)
+For the inner integral, if $r(\theta)<2$, the mean value theorem gives the following upper bound:
+$\int_{0}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)}\leq 2\sqrt{2}.$
+If $r(\theta)>2$, we split the inner integral of Expression (7.19) at $c\in(0,r(\theta))$ to write it as:
+$\int_{0}^{c}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr\ +\ \int_{c}^{r(\theta)}\frac{\sinh(r)}{\sqrt{\cosh(r(\theta))-\cosh(r)}\,\sqrt{\sinh(r)}}\,dr.$ (7.20)
+c is specified later.
+The first term of the previous expression has the following upper bound using the boundary of the integration domain:
+$\frac{1}{\sqrt{\cosh(r(\theta))-\cosh(c)}}\int_{0}^{c}\frac{e^{\tfrac{r}{2}}}{\sqrt{2}}\,dr$ $=\frac{2}{\sqrt{2}}\frac{e^{\tfrac{c}{2}}-1}{\sqrt{\cosh(r(\theta))-\cosh(c)}}\leq\sqrt{2}\frac{\sinh\!\left(\tfrac{c}{2}\right)}{\sqrt{\sinh(c)\,(r(\theta)-c)}},$
+where the inequality follows from the mean value theorem together with the bound for $\sinh$. Since
+$\frac{\sinh\!\left(\tfrac{x}{2}\right)}{\sqrt{\sinh(x)}}\leq 1,$
+we conclude that the first term of Expression (7.20) is bounded from above by
+$\frac{\sqrt{2}}{\sqrt{r(\theta)-c}}.$
+The second term of Expression (7.20) is bounded from above by the following expression using the mean value theorem:
+$\int_{c}^{r(\theta)}\frac{1}{\sqrt{r(\theta)-r}}\,dr=2\sqrt{r(\theta)-c}.$
+With both terms bounded, Expression (7.20) has the following upper bound by choosing $c=r(\theta)-1$:
+$\frac{\sqrt{2}}{\sqrt{r(\theta)-c}}+2\sqrt{r(\theta)-c}=2+\sqrt{2}.$
+Hence, Expression (7.19) is bounded by
+$\frac{2+\sqrt{2}}{\sqrt{\sinh(\rho)}}\int_{-\pi}^{\pi}\frac{1}{\sqrt{|\cos(\theta)|}}\,d\theta\leq\frac{38}{\sqrt{\sinh(\rho)}},$
+where the last integral is bounded from above by $11$. This completes the proof of Lemma 7.6.
 8. Proof of main theorems
-In this section we give proofs to the main theorems of this paper, Theorems 1.4, 1.5, and 1.7.
-We first prove the quantitative main theorem, Theorem 1.5.
-Most of the paper is dedicated to this end.
-Proving it is highly technical and thus for the clarity of the proof we have presented auxiliary results in separate sections, Sections 4, 5, 6, and 7.
-After proving Theorem 1.5, we prove the limit form main theorem, Theorem 1.4. It follows straightforwardly from Theorems 1.5 and 8.1.
-Proving Theorem 1.7 relies on using Theorems 1.5 and 8.2.
-The proof is very similar to that of Theorem 1.4 but due to the probabilistic setting imposing lighter conditions on the sequence, we need Theorem 8.2 to record some key results in random surface theory.
-We proceed to prove Theorem 1.5. We restate the theorem for the convenience of the reader.
+In this section, we give proofs of the main theorems of this paper, Theorems 1.3, 1.4, and 1.5.
+We first prove the quantitative main theorem, Theorem 1.5, as it is needed for the proofs of Theorems 1.3 and 1.4.
+For clarity of the proof, we have presented auxiliary results in separate sections: Sections 4, 5, 6, and 7.
+After proving Theorem 1.5, we prove the limit-form main theorem, Theorem 1.3.
+This result follows straightforwardly from Theorems 1.5 and 8.1 using Properties (BSC), (EXP), and (UND), where Theorem 8.1 is the large-scale Weyl law.
+Proving Theorem 1.4 relies on Theorems 1.5 and 8.2.
+The proof is very similar to that of Theorem 1.3, but as the probabilistic setting does not impose Properties (BSC), (EXP), and (UND), we need Theorem 8.2 to record some key results in random surface theory to replace these properties.
+We proceed to prove Theorem 1.5 and restate the theorem for the convenience of the reader.
 Theorem.
-Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
+Let $X$ be a compact connected hyperbolic surface with Laplace–Beltrami operator $\Delta_{X}$ that has eigenvalues
 $0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
-together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
-$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
-Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
+together with a corresponding orthonormal eigenbasis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values of the eigenvalues as follows:
+$\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}.$
+Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplication operator. Let
 $0<\delta<\tfrac{2}{9}\min(I),\qquad T\delta=\tfrac{\pi}{2},$
 and fix $\tau\in\mathbb{R}$. Then
-$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
-where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
+$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\,\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
+where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, defined by
 $\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
-The coefficient appearing is
-$C_{8}=50000\pi.$
-The proof presented here is clear as the technical details are dealt in different sections, see Sections 6, 5, 6, and 7.
-The core idea of the proof is to split the problem into a geometric side and a spectral side using a well-chosen propagator.
-The geometric side and the spectral side can then be dealt separately.
+The coefficient appearing is $C_{8}=5120000\pi^{2}$.
+The proof presented here is clear since the technical details are dealt with in different sections; see Sections 4, 5, 6, and 7.
+The core idea is to split the problem into a geometric side and a spectral side using a well-chosen propagator.
+These two sides can then be dealt with separately.
 Proof of Theorem 1.5.
@@
 $P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right),\qquad P_{t,\tau}=h_{t,\tau}\left(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\right).$
-Then due to way the operators are defined, we have that
+Then, due to the way the operators are defined, we have
 $P_{t}\psi_{j}=h_{t}(\rho_{j})\psi_{j},\qquad P_{t,\tau}\psi_{j}=h_{t,\tau}(\rho_{j})\psi_{j}.$
-Using the previous, we can write that
+Using the above, we can write
 $|\langle\psi_{j},a\psi_{k}\rangle|=\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|^{-1}\left|\left\langle\psi_{j},\,\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\,\psi_{k}\right\rangle\right|.$
-The left factor containing spectral data is bounded using Proposition 4.1,which can be used since $I$ is compact in $(0,\infty)$, $T\delta=\pi/2$, and $\delta$ is smaller than $(2/9)\cdot\min(I)$:
+The left factor containing the spectral data is bounded using Proposition 4.1, which can be applied since $I$ is compact in $(0,\infty)$, $T\delta=\tfrac{\pi}{2}$, and $\delta<\tfrac{2}{9}\min(I)$:
 $\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})h_{t}(\rho_{k})\,dt\right|^{-1}\leq 8\pi\rho_{j}\rho_{k}\leq 8\pi\max(I)^{2}.$
@@
 $\leq 64\pi^{2}\,\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2},$
-where the Hilbert-Schmidt norm bound is acquired when we consider all the basis elements. The Hilbert-Schmidt norm contains the geometric data of the problem.
-Next we want to apply Proposition 6.1 to bound the Hilbert-Schmidt norm.
-This requires understanding the integral kernels of $P_{t}$ and $P_{t,\tau}$ for which a convenient form is given in Proposition 5.1.
-We apply Proposition 6.1 to acquire the following upper bound for the previous expression:
-$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
-where we opted to use $|X\setminus X(4T)|$ to denote the volume of the points of $X$ with injectivity radius smaller that $4T$. Also,
-$C_{7}=\frac{7200}{\pi}.$
-Hence we can conclude that
-$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
-where
-$C_{7}\cdot 64\pi^{2}\leq 50000\pi=C_{8}.$
-We now turn to Theorem 1.4, the limit form main theorem. We also restate the theorem for convenience.
+where the Hilbert-Schmidt norm bound is obtained when considering all basis elements. The Hilbert-Schmidt norm contains the geometric data of the problem.
+Next, we apply Proposition 6.1 to bound the Hilbert-Schmidt norm:
+$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
+where $|X\setminus X(4T)|$ denotes the volume of the points of $X$ with injectivity radius smaller than $4T$. Also, $C_{7}=80000$. Hence we can conclude that
+$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
+where $C_{7}\cdot 64\pi^{2}=5120000\pi^{2}=C_{8}$.
+We now turn to Theorem 1.3, the limit-form main theorem. We also restate the theorem for convenience.
 Theorem.
-Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
+Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace–Beltrami operators $(-\Delta_{X_{n}})$. The Laplacian $\Delta_{X_{n}}$ has eigenvalues
 $0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
-together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
-$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
+together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values of the eigenvalues as follows:
+$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\tfrac{1}{4}}.$
 Assume that the sequence has the following properties:
@@
     Uniform discreteness: the sequence admits a uniform lower bound on the injectivity radius.
-Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator.
+Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplication operator.
 Then the following properties hold:
   • (i)
-    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
+    $\lim_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a_{n}\,dx\right|^{2}=0,$
   • (ii)
@@
 Properties (ii) and (iii) remain to be proven; Property (i) was established by Le Masson and Sahlsten (see Theorem 1.1 [LS17]).
-We remark that the proof presented here for properties (ii) and (iii) could be adapted with slight additions to also include property (i), thus giving a new proof for the large-scale quantum ergodicity theorem of Le Masson and Sahlsten.
-Passing to the limit and thus acquiring a small upper bound $\epsilon$ is based on carefully using Properties (BSC), (EXP), and (UND) together with the large-scale Weyl law which is used to asymptotically relate the volume of the surface with the number of eigenvalues in an energy window, when we proceed in the sequence of surfaces.
+We remark that the proof presented here for Properties (ii) and (iii) could be adapted with slight modifications to also include Property (i), thus giving a new proof of the large-scale quantum ergodicity theorem of Le Masson and Sahlsten.
+Passing to the limit and thus acquiring a small upper bound $\epsilon$ is based on carefully using Properties (BSC), (EXP), and (UND) together with the large-scale Weyl law, which asymptotically relates the volume of the surface with the number of eigenvalues in an energy window as we proceed through the sequence of surfaces.
 Before the proof, we state the large-scale Weyl law (see [LS17, Sec.9]).
 Theorem 8.1 (Large-scale Weyl law).
-Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces satisfying property (BSC) stated in Theorem 1.4. $X_{n}$ has Laplace-Beltrami operator $\Delta_{X_{n}}$ with eigenvalues
+Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces satisfying Property (BSC) stated in Theorem 1.3. $X_{n}$ has Laplace–Beltrami operator $\Delta_{X_{n}}$ with eigenvalues
 $0=\lambda_{0}^{(n)}<\lambda_{1}^{(n)}\leq\dots\leq\lambda_{j}^{(n)}\leq\dots\xrightarrow{j\to\infty}\infty$
-together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values to eigenvalues as follows:
-$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
+together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values of the eigenvalues as follows:
+$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\tfrac{1}{4}}.$
 Let $I\subset(0,\infty)$ be a compact interval and let
-$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
+$J=\left\{x^{2}+\tfrac{1}{4}:x\in I\right\}.$
 Then
@@
 where $C$ is a constant depending on the sequence of surfaces and the interval $I$.
-We have included using both $\lambda_{j}^{(n)}$ and $\rho_{j}^{(n)}$ since Le Masson and Sahlsten proved their theorem using $\lambda_{j}^{(n)}$ approach.
-However, the theorem works just as well for $\rho_{j}^{(n)}$ as was seen in the statement of the previous theorem.
+We include both $\lambda_{j}^{(n)}$ and $\rho_{j}^{(n)}$ since Le Masson and Sahlsten proved their theorem using the $\lambda_{j}^{(n)}$ approach.
+However, the theorem works just as well for $\rho_{j}^{(n)}$.
+We especially remark that the large-scale Weyl law follows from the Benjamini–Schramm convergence property, Property (BSC).
 We proceed to prove Theorem 1.3.
@@
 $N(X_{n},I)=\#\{j:\rho_{j}^{(n)}\in I\}.$
-Let us define a following function:
+Let us define the following function:
 $\bar{a}_{n}=a_{n}-\frac{1}{|X_{n}|}\int_{X_{n}}a_{n}\,dx.$
-$\bar{a}_{n}$ is a mean 0 normalization of $a_{n}$. Due to orthogonality of $\psi_{j}^{(n)}$ and $\psi_{k}^{(n)}$, we have that
+$\bar{a}_{n}$ is a mean-zero normalization of $a_{n}$. Due to orthogonality of $\psi_{j}^{(n)}$ and $\psi_{k}^{(n)}$, we have
 $\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle=\langle\psi_{j}^{(n)},\bar{a}_{n}\psi_{k}^{(n)}\rangle.$
@@
 $\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\sum_{j:\rho_{j}^{(n)}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}^{(n)}-\rho_{k}^{(n)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(n)},\bar{a}_{n}\psi_{k}^{(n)}\rangle\right|^{2}$
-We can apply Theorem 1.5 to the double sum given that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
-$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(n)})^{3}}\,\left(\|\bar{a}_{n}\|_{2}^{2}\,+\,8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{n}\setminus X_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right),$ (8.2)
-where $|X_{n}\setminus X_{n}(4T)|$ is the volume of the points of $X_{n}$ with injectivity radius being smaller than $4T$, $T\,\delta=\pi/2$, and where
+We can apply Theorem 1.5 to the double sum, given that we assume $\delta\in(0,\tfrac{2}{9}\min(I))$. We get the following upper bound:
+$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(n)})^{3}}\,\left(\|\bar{a}_{n}\|_{2}^{2}\,+\,\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{n}\setminus X_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right),$ (8.2)
+where $|X_{n}\setminus X_{n}(4T)|$ is the volume of the points of $X_{n}$ with injectivity radius smaller than $4T$, $T\delta=\tfrac{\pi}{2}$, and
 $\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{if }x\leq\frac{1}{4},\\ 1,\qquad&\text{if }x>\frac{1}{4}.\end{cases}$
-The constant appearing is
-$C_{8}=50000\pi.$
-Then we apply the properties of the sequence to tidy the previous expression. First we note that
+The constant appearing is $C_{8}=5120000\pi^{2}$.
+Then we apply the properties of the sequence to simplify the previous expression. First we note that
 $\|\bar{a}_{n}\|_{2}\leq\sqrt{|X_{n}|}\,\|\bar{a}_{n}\|_{\infty}.$
-And then, since $(a_{n})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
-$\|\bar{a}_{n}\|_{\infty}\leq 2\,a_{max}.$
-Property (EXP) means that there is a universal lower bound for the spectral gaps which then implies that $\beta(\lambda_{1}^{(n)})$ has a universal lower bound.
+Since $(a_{n})$ is uniformly bounded, we have
+$\|\bar{a}_{n}\|_{\infty}\leq 2\|a_{n}\|\leq 2\,a_{max},$
+where $a_{max}$ is the universal bound for $(a_{n})$.
+Property (EXP) implies a positive universal lower bound for the spectral gaps, which ensures $\beta(\lambda_{1}^{(n)})$ has a universal positive lower bound.
 Denote this by $\beta_{min}$.
-Finally, Property (UND) means that there is a universal lower bound for the injectivity radii.
+Finally, Property (UND) implies a universal positive lower bound for the injectivity radii.
 Denote this by $\operatorname{InjRad}_{min}$.
 Using these observations gives the following upper bound for Expression (8.2):
-$\limsup_{n\to\infty}\ 2\cdot C_{8}\,\frac{\max(I)^{4}\,a_{max}^{2}}{N(X_{n},I)\,T\,\beta_{min}^{3}}\left(|X_{n}|+\frac{8\pi^{2}\,|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{\operatorname{InjRad}_{min}}\right)$
-$=\ \frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\left(\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right)+\frac{8\pi^{2}\,e^{6T}}{\operatorname{InjRad}_{min}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}\setminus X_{n}(4T)|}{N(X_{n},I)}\right)\right),$
-where the equality followed from splitting the limit supremum into two terms.
-Consider the second limit supremum inside the parentheses of the previous expression.
-Due to the large-scale Weyl law, Theorem 8.1, we have that $N(X_{n},I)$ grows asymptotically like the volume of surfaces $|X_{n}|$.
-This means that asymptotically the ratio inside the limit supremum closes to 0 as we proceed in the sequence since Property (BSC) forces the ratio of the volume of the points with small injectivity radius and the volume of the surface to go to zero.
-Hence the previous expression equals
+$\limsup_{n\to\infty}\ 2\cdot C_{8}\,\frac{\max(I)^{4}\,a_{max}^{2}}{N(X_{n},I)\,T\,\beta_{min}^{3}}\left(|X_{n}|+\frac{|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{\operatorname{InjRad}_{min}}\right)$
+$=\ \frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\left(\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right)+\frac{e^{6T}}{\operatorname{InjRad}_{min}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}\setminus X_{n}(4T)|}{N(X_{n},I)}\right)\right),$
+where the equality follows from splitting the limit supremum into two terms.
+Consider the second limit supremum inside the parentheses.
+Due to the large-scale Weyl law, Theorem 8.1, $N(X_{n},I)$ grows asymptotically like the volume $|X_{n}|$.
+Hence the ratio inside the limit supremum approaches 0 since Property (BSC) forces the volume of points with small injectivity radius relative to the total volume to go to zero.
+Therefore, the previous expression equals
 $\frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right).$
-Using the large-scale Weyl law again to the remaining limit supremum, we see that the ratio approaches some constant $C_{9}$. Hence the previous expression equals
-$\frac{2C_{8}\,C_{9}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}$
-$=\ \frac{4\,C_{8}\,C_{9}\,\delta\,\max(I)^{4}}{\beta_{min}^{3}\,\pi},$
-where for the equality we used that $T\,\delta=\pi/2$. If we had chosen originally that
-$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{8}\,\max(I)^{4}}{\beta_{min}^{3}\,\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
-we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
-In the end we prove the probabilistic theorem, Theorem 1.7. We restate the theorem for convenience.
+Using the large-scale Weyl law, the ratio approaches some constant $C_{9}$. Thus we have
+$\frac{2C_{8}\,C_{9}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}=\frac{4\,C_{8}\,C_{9}\,\delta\,\max(I)^{4}}{\beta_{min}^{3}\,\pi},$
+where the equality uses $T\delta=\tfrac{\pi}{2}$. If we choose
+$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{9}\,\max(I)^{4}}{\beta_{min}^{3}\,\pi}\right)^{-1},\tfrac{2}{9}\,\min(I)\right),$
+then the previous expression is bounded by $\epsilon$. We assume this choice in the beginning of the proof to conclude the proof.
+In the end, we prove the probabilistic theorem, Theorem 1.4. We restate the theorem for convenience.
 Theorem.
-Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus.
-Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
+Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil–Petersson probability distribution on the moduli space $\mathcal{M}_{g}$, and where $g$ is the genus.
+Let $\Delta_{g}$ be the Laplace–Beltrami operator of $X_{g}$ with eigenvalues
 $0=\lambda_{0}^{(g)}<\lambda_{1}^{(g)}\leq\dots\leq\lambda_{j}^{(g)}\leq\dots\xrightarrow{j\to\infty}\infty$
-together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
-$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
-Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator.
-Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
+together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values of the eigenvalues as follows:
+$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\tfrac{1}{4}}.$
+Let $I\subset(0,\infty)$ be a compact interval, and let $(a_{g})$ be a sequence of uniformly bounded functions on $X_{g}$, each acting as a multiplication operator.
+Then Properties (i)(iii) of Theorem 1.3 hold with high probability as the genus $g$ grows to infinity.
 Properties (ii) and (iii) remain to be proven; Property (i) was already proven by Le Masson and Sahlsten [LS24].
-The proof is very close to that of Therem 1.4.
-Now we do not assume Properties (BSC), (EXP), and (UND), so these need to be replaced.
-This is done by the following theorem acquired from the theory of the Weil-Petersson random surfaces.
+The proof is very close to that of Theorem 1.3.
+Now, we do not assume Properties (BSC), (EXP), and (UND), so these need to be replaced.
+The following theorem records analogous tools acquired from the theory of Weil–Petersson random surfaces.
 Theorem 8.2.
-Let $X$ be a $\mathbb{P}_{g}$-random surface, where $\mathbb{P}_{g}$ is the Weil-Petersson probability distribution on the moduli space $\mathcal{M}_{g}$ with genus $g$. It has eigenvalues
+Let $X$ be a $\mathbb{P}_{g}$-random surface, where $\mathbb{P}_{g}$ is the Weil–Petersson probability distribution on the moduli space $\mathcal{M}_{g}$ with genus $g$. It has eigenvalues
 $0=\lambda_{0}<\lambda_{1}\leq\dots\leq\lambda_{j}\leq\dots\xrightarrow{j\to\infty}\infty$
 together with associated values
-$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}.$
-Then with probability 1o(1)1-o(1) $X$ has the following properties:
+$\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}.$
+Then $X$ has the following properties with probability 1o(1)1-o(1) for large genus:
   • (a)
@@
 where $I\subset(0,\infty)$ is a compact interval and
-$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
-The theorem is acquired from two articles: [Mon22, HMT25].
-Properties (a), (b), and (d) together with the rates is acquired from [Mon22], and Property (c) together with rate is acquired from [HMT25].
+$J=\left\{x^{2}+\tfrac{1}{4}:x\in I\right\}.$
+The theorem is taken from two articles: [MON22, HMT25a].
+Properties (a), (b), and (d), together with the rates, are taken from [MON22], and Property (c), together with the rate, is taken from [HMT25a].
 When comparing the proof of Theorem 1.4 with the proof of Theorem 1.3, Property (a) replaces Property (UND), Property (b) replaces Property (BSC), and Property (c) replaces Property (EXP).
-Additionally, Property (d) is a suitably modified version of the large-scale Weyl law of Le Masson and Sahlsten working in this probabilistic setting.
+Additionally, Property (d) is a suitably modified version of the large-scale Weyl law of Le Masson and Sahlsten, applied in this probabilistic setting.
 Proof of Theorem 1.4.
-We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)$ such that for every $\tau\in\mathbb{R}$, we have that
-$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle\right|^{2}<\epsilon,$ (8.3)
+We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)$ such that for every $\tau\in\mathbb{R}$, we have
+$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)}\in I,\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle\right|^{2}<\epsilon,$ (8.3)
 where
 $N(X_{g},I)=\#\{j:\rho_{j}^{(g)}\in I\}.$
-Let us define a following function:
+Let us define the following function:
 $\bar{a}_{g}=a_{g}-\frac{1}{|X_{g}|}\int_{X_{g}}a_{g}\,dx.$
-$\bar{a}_{g}$ is a mean 0 normalization of $a_{g}$. Due to orthogonality of $\psi_{j}^{(g)}$ and $\psi_{k}^{(g)}$, we have that
+$\bar{a}_{g}$ is a mean-zero normalization of $a_{g}$. Due to the orthogonality of $\psi_{j}^{(g)}$ and $\psi_{k}^{(g)}$, we have
 $\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle=\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle.$
 Hence the left-hand side of Expression (8.3) equals
-$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle\right|^{2}.$
-We can apply Theorem 1.5 to the double sum givem that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
-$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(g)})^{3}}\,\left(\|\bar{a}_{g}\|_{2}^{2}+8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right),$
-where $|X_{g}\setminus X_{g}(4T)|$ is the volume of the points of $X_{g}$ with injectivity radius less than $4T$, $T\,\delta=\pi/2$, and where
+$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)}\in I,\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle\right|^{2}.$
+We can apply Theorem 1.5 to the double sum, given that we assume $\delta\in(0,\tfrac{2}{9}\min(I))$. We get the following upper bound for the previous expression:
+$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(g)})^{3}}\,\left(\|\bar{a}_{g}\|_{2}^{2}+\|\bar{a}_{g}\|_{\infty}^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right),$
+where $|X_{g}\setminus X_{g}(4T)|$ is the volume of the points of $X_{g}$ with injectivity radius less than $4T$, $T\delta=\tfrac{\pi}{2}$, and where
 $\beta(x)=\begin{cases}1-\sqrt{1-4x},\qquad&\text{ if }x\leq\frac{1}{4},\\ 1,\qquad&\text{ if }x>\frac{1}{4}.\end{cases}$
-The constant appearing is
-$C_{8}=50000\pi.$
-First notice that
-$\|a_{g}\|_{2}\leq\sqrt{|X_{g}|}\,\|a_{g}\|_{\infty}.$
-Since the sequence $(a_{g})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
-$\|\bar{a}_{g}\|\leq 2\,a_{max}.$
-Hence the previous limit supremum expression is bounded from above by
-$2\cdot C_{8}\,\frac{a_{max}^{2}}{T}\,\limsup_{g\to\infty}\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right).$ (8.4)
-Using Theorem 8.2, we have that
-$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+8\pi^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
-$\leq\ \frac{1}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(\frac{3}{16}-\eta\right)^{3}}\,\left(|X_{g}|+8\pi^{2}\,|X_{g}|\,\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
-with propability $1-\mathcal{O}(g^{-c(\eta)})$, where $\eta\in(0,3/16)$, $c(\eta)$ is a positive constant depending on $\eta$, and where
-$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
+The constant appearing is $C_{8}=5120000\pi^{2}$.
+First, notice that
+$\|\bar{a}_{g}\|_{2}\leq\sqrt{|X_{g}|}\,\|\bar{a}_{g}\|_{\infty}.$
+Since the sequence $(a_{g})$ is uniformly bounded, we have
+$\|\bar{a}_{g}\|_{\infty}\leq 2\|a_{g}\|_{\infty}\leq 2a_{max},$
+where $a_{max}$ is the universal bound for $(a_{g})$. Hence the previous limit supremum expression is bounded from above by
+$2\cdot C_{8}\,\frac{a_{max}^{2}}{T}\,\limsup_{g\to\infty}\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right).$ (8.4)
+By using Theorem 8.2, we have
+$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
+$\leq\ \frac{|X_{g}|}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(1-\sqrt{1-4(\tfrac{1}{4}-\eta)}\right)^{3}}\,\left(1+\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
+with probability 1o(1)1-o(1), where $\eta\in(0,\tfrac{1}{4})$ and
+$J=\left\{x^{2}+\tfrac{1}{4}:x\in I\right\}.$
 The previous expression has a tidier upper bound:
-$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+8\pi^{2}\,e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right).$
-This means that with high propability Expression (8.4) equals
-$2\,C_{8}\,C_{10}\,\frac{a_{max}^{2}}{T}=\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\,\delta$
-where equality followed from using that $T\,\delta=\pi/2$. $C_{10}$ is a constant upper bound coming from
-$\frac{1}{\left(\frac{3}{16}-\eta\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}.$
-If we had chosen originally that
-$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
-we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
+$\frac{1}{\left(1-\sqrt{4\eta}\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right).$
+This means that when passing to the high-genus limit, we have
+$\frac{1}{\left(1-\sqrt{4\eta}\right)^{3}\,\mathcal{O}(\max(J)-\min(J))}\,\left(1+e^{6T}\,\frac{\mathcal{O}(g^{-\frac{1}{4}})}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right)\xrightarrow{g\to\infty}C_{10}$
+for some constant $C_{10}$. This means that with high probability, Expression (8.4) has the following upper bound:
+$2\,C_{8}\,C_{10}\,\frac{a_{max}^{2}}{T}=\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\,\delta,$
+where the equality follows from using $T\delta=\tfrac{\pi}{2}$.
+If we had originally chosen
+$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{10}\,a_{max}^{2}}{\pi}\right)^{-1},\tfrac{2}{9}\min(I)\right),$
+we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice at the beginning of the proof to conclude the argument.
+Appendix A On quantum mixing of physical-space observables on flat 2-tori
+Let $\mathbb{T}_{L}^{2}$ be a flat $2$-torus with side length $L$.
+We consider the sequence $\{\mathbb{T}_{j}^{2}\}_{j\in\mathbb{N}}$ of growing flat $2$-tori obtained by letting $L=j$.
+The torus $\mathbb{T}_{j}^{2}$ has eigenvalues and corresponding eigenfunctions
+$\lambda_{(m,n)}^{(j)}=\frac{4\pi^{2}}{j^{2}}(m^{2}+n^{2}),\qquad\psi_{(m,n)}^{(j)}(x,y)=e^{\frac{2\pi i}{j}(mx+ny)},$
+where $(m,n)\in\mathbb{Z}^{2}$.
+Let $\rho_{(m,n)}^{(j)}=\sqrt{\lambda^{(j)}_{(m,n)}-\tfrac{1}{4}}$.
+We consider Properties (II) and (III) of Theorem 1.3 and show that the corresponding properties fail for the sequence $\{\mathbb{T}_{j}^{2}\}$.
+In fact, it suffices to show that Property (II) fails, since this already implies the failure of Property (III).
+Proposition A.1 (Large-scale quantum mixing breaks).
+Let $I\subset(0,\infty)$ be a compact interval. There exists $\epsilon>0$ and a sequence of bounded functions $\{a_{j}:\mathbb{T}_{j}^{2}\to\mathbb{C}\}$ such that for every $\delta>0$ we have that
+$\limsup_{j\to\infty}\frac{1}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}\sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\,\sum_{\begin{subarray}{c}(p,q)\in\mathbb{Z}^{2}:\rho_{(p,q)}^{(j)}\in I,(p,q)\neq(m,n),\\ |\rho_{(m,n)}^{(j)}-\rho_{(p,q)}^{(j)}|<\delta\end{subarray}}|\langle\psi_{(m,n)}^{(j)},a\psi_{(p,q)}^{(j)}\rangle|^{2}\geq\epsilon.$
+Proof of Proposition A.1.
+Let $\epsilon>0$ be specified later, and define $a_{j}(x,y)=e^{\frac{2\pi i}{j}(x+y)}$. Let $\delta>0$ be arbitrary; in particular, it is independent of $\epsilon$ and the choice of $a_{j}$.
+We estimate the double sum appearing in the statement from below. A convenient lower bound is obtained by restricting to pairs of indices of the form $(m,n)$ and (m1,n1)(m-1,n-1):
+$\sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\bm{1}_{\begin{subarray}{c}\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta\end{subarray}}\,|\langle\psi_{(m,n)}^{(j)},a_{j}\psi^{(j)}_{(m-1,n-1)}\rangle|^{2}$
+$=\ \sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\bm{1}_{\begin{subarray}{c}\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta\end{subarray}},$
+where the equality follows from the identity
+$a_{j}\psi_{(m-1,n-1)}^{(j)}=\psi_{(m,n)}^{(j)}.$
+Consequently, the $\limsup$ in the statement admits the lower bound
+$\limsup_{j\to\infty}\frac{1}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}\sum_{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I}\bm{1}_{\begin{subarray}{c}\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta\end{subarray}}$
+$=\ \limsup_{j\to\infty}\frac{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)},\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}|<\delta\}}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}.$
+As $j\to\infty$, the spacing between $\rho_{(m,n)}^{(j)}$ and $\rho_{(m-1,n-1)}^{(j)}$ tends to zero uniformly for indices with $\rho_{(m,n)}^{(j)}\in I$.
+In particular, for sufficiently large $j$, the condition $|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}^{(j)}|<\delta$ is automatically satisfied whenever both terms lie in $I$.
+Moreover, we have the asymptotic relation
+$\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)},\rho_{(m-1,n-1)}^{(j)}\in I\}\sim\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}.$
+It follows that
+$\limsup_{j\to\infty}\frac{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)},\rho_{(m-1,n-1)}^{(j)}\in I,|\rho_{(m,n)}^{(j)}-\rho_{(m-1,n-1)}|<\delta\}}{\#\{(m,n)\in\mathbb{Z}^{2}:\rho_{(m,n)}^{(j)}\in I\}}=C$
+for some constant $C>0$. Choosing $\epsilon=\tfrac{C}{2}$ yields the desired conclusion.
+In Subsection 1.1, we noted that it is an interesting problem whether analogues of Properties (II) and (III) of Theorem 1.1 hold for a fixed flat $2$-torus.
+The example above does not readily extend to that setting: when the torus is fixed and one considers higher and higher eigenvalues, the spectrum does not exhibit the same clustering, and nearby eigenvalues do not accumulate in the same way.
 Acknowledgements.
 The author would like to thank Prof.
-Mostafa Sabri, Prof.
-Tuomas Sahlsten, and Prof.
-Farrell Brumley for discussions that led to this article.
+Farrell Brumley, Prof.
+Mostafa Sabri, and Prof.
+Tuomas Sahlsten for discussions that led to this article.
 The author also thanks Prof.
@@
 Joe Thomas for their guidance while writing this article.
-Our use of the exponential is inspired by the upcoming article [HLM+] of the author with Félix Lequen, Søren Mikkelsen, Tuomas Sahlsten and Henrik Ueberschär on quantum ergodicity for eigenfunctions for Schrödinger operators on hyperbolic surfaces.
+Finally, the author wants to thank Prof.
+Zeév Rudnick and Prof.
+Mostafa Sabri for their helpful comments on the earlier drafts.
+Our use of the exponential is inspired by the upcoming article [30] by the author with Félix Lequen, Søren Mikkelsen, Tuomas Sahlsten, and Henrik Ueberschär on quantum ergodicity of eigenfunctions for Schrödinger operators on hyperbolic surfaces.
 We thank the other authors of that paper.
-The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, Research council of Finland project: Quantum chaos on large and many body systems grant numbers 347365, 353738, and Aalto University.
+The research conducted here is supported by the Vilho, Yrjö and Kalle Väisälä Foundation, and the Research Council of Finland, project: Quantum chaos on large and many-body systems, grant numbers 347365, 353738, and Aalto University.